A Complete Bibliography of Computers and Security, 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

26 December 2021 Version 1.01

Title word cross-reference

2 [NTH04]. 24 × 7 [Han00y]. GF(2^{m}) [CL05]. H [Arm02b]. H_{2}A [CBB05]. I [Kov00b]. K [LV02]. Z_{n} [LWL09].

-Nearest [LV02].

.NET [PE06].

1 [Ano00e, Ano00i, Ano00b, Ano00c, Ano00f, Ano00d, Ano00h, Ano00g, Ano01-46, Ano01-45, Ano01-42, Ano01-38, Ano01-43, Ano02-38, Ano02-37, Ano02-39, Ano02-40]. 1st [Ano01-40].

2000 [Ano00e, Ano00i, Ano00b, Ano00c, Ano00f, Ano00d, Ano00h, Ano00g, Ja02]. 2001 [Ano01k, Ano01-46, Ano01-45, Ano01-39, Ano01-41, Ano01-42, Ano01-38, Ano01-44, Ano01-40, Ano01-43]. 2002 [Ano02-38, Ano02-37, Ano02-43, Ano02-41, Ano02-42, Ano02-39, Ano02-40]. 2003 [Ano03-33, Ano03-39, Ano03-40, Ano03-34, Ano03-36, Ano03-35, Ano03-38, Ano03-37]. 2004 [Ano04o, Ano04i, Ano04k, Ano04l, Ano04p, Ano04j, Ano04n, Ano04m]. 2005 [Ano05g, Ano05d, Ano05f, Ano05k, Ano05e, Ano05j, Ano05i, Ano05h]. 2006 [Ano06t, Ano06v, Ano06w, Ano06-27, Ano06u, Ano06z, Ano06y, Ano06x]. 2007 [Ano07-35, Ano07-37, Ano07-38, Ano07-39, Ano07-32, Ano07-34, Ano07-40, Ano07-33, Ano07-36]. 2008 [Ano08-27, Ano08-28, Ano08-29, Ano08-30, Ano08u, Ano08v, Ano08w, Ano08x, Ano08y, Ano08z]. 2009 [Ano09y, Ano09-28, Ano09-27, Ano09z]. 21 [Kov01b, Nos00a]. 21-st [Kov01b, Nos00a]. 21st [Sch01a]. 22 [McK04]. 2nd [Mou05].

3 [Col00]. 31 [Ano01-39, Ano01-41, Ano01-44, Ano02-43]. 3G [DP06].

9-11 [Beq02a].

= [FvS05].

A-Changin [Hin01i]. AAIs [LOP04]. Abstracts [Ano02a, Ecc01a, Ecc01b, Ecc01c, Mey02, Nic00a, Nic00b, Nic00c, Nic00d, Pal00a, Pal00b, Pal00c, PM01, Pal01a, Pal01b, Pal01c, Pal01d, Pal01e]. academic [Bis02, DKS07, Sch07c]. Acceptance [CFRR02]. Access [BE01, HS00, WS02, YL04, BS09, CC02, CE04, GS03, HW03, JCKL04, KLL09, KHKK06, OKE09, PE09, PK07, Pur02, RK06, SSF^{+}09, SV07, TE06, Zho02, ZL03]. accountability [BAS09]. accurate [Lea04, TXL09]. accurately [Sch06g]. Accused [Han00-85]. achieve [GRNR09]. Achieving [KK00]. Achilles [Hin00a]. acquisition [Wol03c]. ACSAC [Ecc01d]. Act [Beq03, Ja02, Pou03b]. active [LG07a, LG07b, LG07c, LGW07]. Activities [Han00-51]. ad [KVD06, KVD07, OCB09, TC05]. Adaptable [KK07]. Adapting [PE09]. Adaptive [CBB05, YZ07, ZCW04]. Additional [MZ00]. address [PE09]. addresses [Sch07a]. Addressing [AGAS01, AOG^{+}02]. Ado [Nik02c]. Adopters [Han01r]. Advanced [CF07, And05]. adversum [Hin01g]. Advert [Ano09p]. Advocates [Han01q]. adware [Sch03c]. After [Nov07]. Again [Han00d, Han00-44, Han00-89, Hin02c]. Against [Zen01a, BNSZ09, CLC04, GVL09, LS05, MFDVGT08a, MFDVGT08b, MFDVGT08c, Zen01b, ZCW04]. Age [Web00, GvS05]. Agencies [Han00a]. Agent [BE01, GFV05, VSR07, DGY^{+}05, LCC04]. Agent-Based [GFV05, BE01, DGY^{+}05]. agents [DR04]. Agree [Han00-45]. Agreement [Han00-87, CMY07, JL07a, JL07b, JL07c, Jua04, MK06, SW06, Tse07, YC09]. Agreements [Han00-45]. Aims [Ano09q, Ano09r, Ano09s]. Airwaves [Han00x]. aka [Nik02a]. al. [HC04b]. alert [SG09, ZLCF08a, ZLCF08b, ZLCF08c]. alerts [LCK^{+}06]. algorithm [CO09, CC02, LG07a, LG07b, LG07c, YH07d]. Aligning [DF06, Sch05g].

Alive [Han01y]. alleged [BLP05]. Alliance [WZM^{+}00]. Alternative [Han00-64, PMRZ01a, SLG04]. America [Beq00a]. American [Han00b, Woo01]. Analysis [GvS01, GM00, KFS^{+}03, Nos00b, RF01, SSMJ05, Asl04b, Bis02, BCS03, BvS03, BP03, CH04, DS09, HCS06, JPL04, KLL09, KS05, KS06, Kon09, Kum09, LCK^{+}06, MP08a, MP08b, MP08c, OZ07, RS04, Sch07a, SAF09, Wol03d, YH07d]. analytic [KL07]. analyze [SJ07]. analyzing [AMR07, LGW07]. Anomaly [GTDVMFV09, CP03, CC04, MP08a, MP08b, MP08c, OL03, OG07, Wan05, WGZY06, YZ07]. Anomaly-based [GTDVMFV09]. Anonymity [Han00t, MK06, WH04]. anonymous [CC03, CL06, CJC04, Chi08a, Chi08b, Chi08c, JPL05, MGMSML^{+}09]. Answer [Bar00b, Bar01b]. Anti [GRJ07, Pou02a, GT06, WC07]. Anti-keylogging [GRJ07]. anti-spam [GT06]. anti-spamming [WC07]. Anti-Terrorism [Pou02a]. any [Ano05q]. AOL [Han00d]. Apache [Han00e]. APIs [MvO09]. Application [MP08a, MP08b, MP08c, TE06, DMWS09, HCBLETRG06, Kin04, MFDVGT08a, MFDVGT08b, MFDVGT08c, Nab05]. Application-based [MP08a, MP08b, MP08c]. Applications [Fra01, HS00, DP06]. applied [Zuc07]. Applying [KK02, Kin04, MC03a]. Appraisal [SPLS01]. Approach [BvS00, BE01, EvS00b, Nos00b, DSX06, HKP02, KVD07, Lea04, LG09, Li04, Mac05, MPD06, MLD08, OZ07, SJ07, SS07, SK03, SYL09, TXL09, TS05a, Wan05, ZHH09]. Approaches [EvS00a, FBB06, Hin03f]. April [Ano00e, Ano01-46, Ano03-33]. architectural [SLG04]. Architecture [GKL^{+}00, KKHR01, TM00, DMWS09, GBBS08a, GBBS08b, GBBS08c, KHKK06, SSF^{+}09, TM09]. Arrested [Han00-81]. art [MPD06]. Articles [Ano02a, Ecc01a, Ecc01b, Mey02, Nic00a, Nic00b, Nic00c, Nic00d, Pal00a, Pal00b, Pal00c, PM01, Pal01a, Pal01b, Pal01c, Pal01d, Pal01e, Ecc01c]. artificial [KPsH^{+}08a, KPsH^{+}08b, KPsH^{+}08c]. aspect [MLD08, MSLD09]. aspect-oriented [MLD08, MSLD09]. aspects [AN03, GvS08a, GvS08b, GvS08c]. Assesses [Han00q]. Assessing [FBP07, LPF06, SE02, KK06]. Assessment [Ber07, CCC09, DKS07, Fur07a, Fur07b, Fur07c, Hin03d]. assignment [CJ03, HW03, Lin01, WL05]. assisted [Art04]. Association [PPBH00]. Assumption [Hin00b]. Assurance [BB05, Kov01a, GC09, SKM08a, SKM08b, SKM08c]. assurances [FvS05]. asymmetric [Mat03]. attachment [CL03]. Attack [Han00-40, Han00-63, CK03, Kum06, MFDVGT08a, MFDVGT08b, MFDVGT08c, Rob07, ZLCF08a, ZLCF08b, ZLCF08c, ZHH09, ZCW04]. Attackers [Sch03f, CK03]. Attacks [CZ03, Han00-62, McK04, Woo01, ÁP03, BIC08a, BIC08b, BIC08c, BLP05, BNSZ09, CLC04, CLK09, GVL09, GKS^{+}06, HMM09, HH05, HFC^{+}08, LS05, Li04, Li06, MS09, MK04, PML09, Sch02a, SYL09, ZC05, Her09b]. Attitudes [FDIR00, CF05]. attribute [OKE09, vdHvS03]. Attributes [Wil00]. Auction [Han00-49, CC03, Che04, JPL05]. Audit [All01, BS07a, BS07b, BS07c, WGZY06]. auditing [FBV06]. August [Ano02-38, Ano05g, Ano07-35]. authenticated [Jua04, WLH06]. Authentication [BCC02, BC05, CC01, CJT02, CFRR02, CF05, FDIR00, KVD07, LOP04, MW06, PMRZ00b, Ano05q, Art04, Asl04a, Asl04b, Ayo06, CBB05, CZ03, Chi08a, Chi08b, Chi08c, CF07, DSGP06, FLZ02, FCZ05, Her09a, HLTJ09, sHCP09, JPL04, LFHT07, LH03, McK04, OHB08, Pha06, RFR07a, RFR07b, RFR07c, SBS09, SLH03, SW06, Tsa08, WLT03, WDCJ09, WC03, YW04, YWL05, YC09, YRY05]. authenticator [jLC07]. AUTHMAC DH [Asl04b]. Author [Ano00a, Ano01a, Ano02b, Ano03b]. Authority [CB01]. authorization [LOP04]. authorizations [GRNR09]. Authors [Ano01m, Ano01n, Ano01r, Ano01s, Ano01t, Ano02o, Ano02r, Ano02t, Ano03n, Ano03o, Ano03-43, Ano06g, Ano06h, Ano06i, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano09i, Ano01o, Ano01p, Ano01q, Ano02p, Ano02q, Ano02s, Ano03p, Ano03q]. autoexecution [Sch03c]. Automated [BIC08a, BIC08b, BIC08c, Pal08a, Pal08b, Pal08c, BS09, LL09]. availability [RS03]. averaged [Li06]. avoid [ZC05]. Award [Han00c]. aware [KHKK06, OHB08]. awareness [CITN07, DCF07, DKS07, KS08a, KS08b, KS08c, KK06, KK08a, KK08b, KK08c, RM08a, RM08b, RM08c, Sch04i]. awareness-fitting [Sch04i].

B2B [Han00-98]. B2C [vKH00]. back [Sch06e]. backdrop [Pou01b]. bad [Kum06]. balanced [NCS06]. Balancing [Nos00a]. Baltimore [Han00f]. bandwidth [KLL09]. Banking [CDD^{+}02, MvO09]. barcodes [NTH04]. barriers [FTP08a, FTP08b, FTP08c]. Based [CC01, GFV05, PMRZ00b, PMRZ00c, SC02, TM00, AJS06, Ano05q, BE01, CLX05, CBR06, CO09, CL06, CC02, CCH03, CZG^{+}05, CL05, CC04, DGY^{+}05, DS09, ES09, FLZ02, GTDVMFV09, GSK09, GLK08a, GLK08b, GLK08c, GVL09, GYLC09, HC03, HR08a, HR08b, HR08c, HAGTA08a, HAGTA08b, HAGTA08c, Her09a, HFC^{+}08, HLL05, sHCP09, JPL04, JHgJ05, KC05, jLC07, LCLL09, LG09, LTH03, Li04, LG07a, LG07b, LG07c, LGW07, LWT^{+}09, LZXW09, LCC04, MW06, MP08a, MP08b, MP08c, MP07, OKH07, Pha06, PW06, RK06, SPP07a, SPP07b, SPP07c, SLH03, SV07, SYL09, TXL09, TNG04, TE06, Tsa08, VSR07, WLT03, WL05, WGZY06, WZM09, Woo04, WC05, YC09, YZPL09, vSvS06a]. Baselines [JS02]. basics [Han02]. basis [CL05]. batch [NCS06]. Batten [Han01d]. battle [Sch03l]. Battles [Han00y]. BCM [Nos00b]. Be [Han00-97, vKH00, PMRZ00a]. Becomes [Han00-76]. Beer [Hin00b]. behavior [CZG^{+}05, Kum09, MP07, OL03, RKR09, WGZY06]. behavior-based [MP07]. behaviors [SSMJ05]. behaviour [Lea03b]. behavioural [VvS04]. beliefs [FTP08a, FTP08b, FTP08c]. Benefits [AGAS01, AOG^{+}02, RR06]. Best [Han00k]. better [Sch02g]. Between [Gup00, Han01r, AH09, BTY^{+}07, BEP04, JS03, LK07, Sch02d, Sch04e]. Beyond [FN05, Hei03]. bids [CC03]. Big [Han00-101]. bigger [Sch02g]. bilinear [DSGP06]. Bill [Ano02-31, Ano02f]. billion [LCLL09]. binary [SKBM06]. bioinformatics [TXL09]. biological [LK07]. Biometric [MS00, PMRZ00b, Rob07, SWA^{+}04, NTH04, RFR07a, RFR07b, RFR07c]. Biometric-Based [PMRZ00b]. Biz [Han00-47, Han00-103]. Black [Kov00a]. Blame [Han00-100]. Blind [GSK09, CLX05, LCC04]. Block [Gup00, PW06]. Blocks [Ano01-49]. bloom [GVL09]. Blue [Hin02a]. Bluetooth [Bar00b, Bar01b]. Board [Ano01-30, Ano01-31, Ano01-32, Ano01-33, Ano01-34, Ano01-35, Ano01-36, Ano02-32, Ano02-33, Ano02-34, Ano02-35, Ano02-36, Ano03z, Ano03-27, Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano03-32, Ano06m, Ano06n, Ano06o, Ano07l, Ano07m, Ano07n, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano07t, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Han00-90, Ano01-37]. Book [Ecc01e, Ecc01f]. boon [Sch04f]. boosting [JST^{+}07]. both [vS05a]. bottleneck [KLL09]. boundaries [Sch06c]. Box [Sch03c, Sch07b]. Brand [Car01]. brave [Sch07e]. Breach [Han00-39, Nov07]. breaches [LSK09]. Bricks [McG02]. Bridging [BTY^{+}07]. Brief [Ano02-28, Ano02-31, Ano02-29, Ano02-30, CR02, Ecc01g, Hil02a, MH02c]. bring [Bha03]. Brother [Han00-101]. Bruce [Hei03]. Bug [Han00c, Han00-69, Han00-80]. Bug-in-Your-Site [Han00c]. Bugbear [Arm02a]. bugs [Hin02a]. Building [LWT^{+}09, McG02, OKE09, ZLCF08a, ZLCF08b, ZLCF08c, Hun02]. Builds [Han00-98]. Business [Bar00a, Fin00, Han00-37, Nab05, OZ07, RM05, vSvS05]. businesses [And05]. Busted [Han00d, Han00-53]. Busts [Han00-44]. Buy [Han00-49].

Calendar [Ano01j, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano02c, Ano03a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano06c, Ano06a, Ano06b, Ano06j, Ano06k, Ano06l, Ano07k, Ano04a, Ano05c]. Calender [Ano06d, Ano04e, Ano04f, Ano04g, Ano04h]. Call [Ano06e, Ano09a, Ano09b, FGG^{+}04]. Calls [Han00-47]. Can [FL05, Han00-49, Hin01d, PK07, RSS02]. Canadian [Han00g, Han00h]. cannot [Fur05]. can’t [Her09b, Sch03e]. Capital [HD05]. capture [AMB06]. Card [CJT02, Han00b, Han00g, Han00-81, Ano05q]. cards [FCZ05, Jua04, Pha06, SLH03, WC03, YW04]. Care [Han00-42, Kov00b, SE02, vSvS06b]. Careless [Hin03a]. carrying [Art04]. case [BTY^{+}07, Dou03, HD05, Kin04, KK08a, KK08b, KK08c, MvO09, Sch04d]. cash [CLX05, CL03]. categories [VE02, VEL08]. causal [ZLCF08a, ZLCF08b, ZLCF08c]. Cause [Han00-93, LSK09]. Caused [Han00g]. Causes [Han00j]. CD [HM04]. Cellular [Han00i]. Centre [Hil02b, McK03b]. centric [BE01, CS09]. Century [Sch01a, Kov01b, Nos00a]. CERT [Han00l, Han00-65]. certificate [Ung04].

Certificates [Han00p, PMRZ01b, GKH09]. Certification [CB01, DeM02, EvS00b, vSvS01, Sch05i]. Certifications [Han00-84]. Certified [Han00-84]. Chained [BC05]. chaining [CBB05, PW06]. chains [Hin02f]. Challenge [PMRZ00c, Luo06]. Challenge-Based [PMRZ00c]. Challenges [OKH07, Bra07a, Bra07b, Bra07c, DFSC09, FJK06, GTDVMFV09, Hun02]. Change [Kov01b, Li06, Sch04a]. changes [Sch07d]. Changin [Hin01i]. changing [Sch06b]. channel [YZPL09]. channels [Wu05]. Characteristics [Wil00, RFR07a, RFR07b, RFR07c]. Characterization [CLC04]. Charged [Han00-80]. Charges [Han00-56, Han00-80]. charter [McK03a]. Chat [Han00j]. checking [BKPA09, KN03]. Chen [ZL03]. Chen-Chung [ZL03]. chief [Ano02n, Ano03k, Ano03l, Ano03m, Sch01b, Sch02b, Sch02c, Han01a, Wol08h, Wol08i, Wol08j]. Child [Han00-53]. China [Han01b]. Chinese [CC02]. Chips [Han00-46]. chosen [ZCW04]. Chung [ZL03]. CIA [Han00j, Han00-51]. CIAO [Kov01a]. CIDS [DGY^{+}05]. CIO [Han00k]. Ciphers [Gol01, DS09, SK03]. circle [Hin02b]. circuit [BNSZ09]. circumstances [Wol03g]. Classes [Han00-62]. Classification [LCLL09]. Classified [Han00-82]. classifier [LV02]. Clean [Han00-92]. click [Bha03]. Client [KJKV09, KLL09, MSZ05]. Client-side [KJKV09, KLL09, MSZ05]. clients [WLH06]. CLIQUES [CMY07]. closing [Ber07]. Clustering [SV07, OL03, OZ07]. clusters [AP04]. CMU [Han00l]. COBIT [vS05a]. Code [Han00r, Han00-99, JS03, Kum09, MLD08, RS03, SF04, Ano01-49]. Colander [CG00]. collaboration [OZ07]. collaborations [LZXW09]. collar [Beq02b]. Collection [Kov00c]. color [CO09]. Combat [Han00-36]. Combine [EvS00b]. Commandments [Han00-96]. commands [JST^{+}07, KC05]. comment [ZL03]. Commerce [Beq00a, Han00r, Nab05, TMM01, Wri01, Zuc04, Zuc07]. commercial [Kja06]. Committee [Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-28, Ano02u, Ano03x, Ano09t, Ano09u, Ano09v, Ano01-29, Ano02v, Ano02w, Ano02x, Ano02y, Ano02z, Ano02-27, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano03w, Ano03y]. Common [Gup00]. communication [Kum09, NCS06]. Communications [BCC02, Ja02, LL02, Seb06, SKM08a, SKM08b, SKM08c]. Comparative [LFHT07, LOP04]. Comparing [PE06]. comparison [BEP04, BCV01, Hin03f, Kja06, SP06, VFMP05]. Compiled [Ano02-31, Ano02f]. complete [Hin02b]. Compliance [vS05b, BAS09, FN05, VvS04]. comprehensive [Pou02b]. compression [ASP03]. compromises [Sch05b]. Compsec [Hil02b, Hin00f, Hin02b, Ano01k, Hil02b, McK03b]. Comput [McK04]. compute [YH07d]. Computer [AGAS01, AOG^{+}02, BA04, Cus01, DM01, Han00-82, Han01y, Hin03b, Luo06, PPBH00, Wol03a, Beg04, Cae02, DSX06, GBBS08a, GBBS08b, GBBS08c, GS03, HH05, HYJ^{+}09, HD05, JEL03, Kja06, KCC09, LK07, MK04, PdVGA08a, PdVGA08b, PdVGA08c, Pou02b, Rog03, RS04, Sam09, WGZY06, WH04, hZxZtZ06]. Computers [Ano03b, Ano03c, Han00-49]. computing [Sch07b, WLH06]. concept [TS05b]. Concepts [Fin00, Ber03]. conceptual [VE04]. Concern [Han01n]. concerning [Ano03l, Sch06d]. Concerns [Beq01c, Dhi01, Han00-93, MSLD09]. concise [Kon09]. Conducting [Han00-51]. Conference [Eva01, Hil02b, McK03b]. conferences [Ano01j, Ano02c, Ano03a, BG06]. Confidence [HMM09]. configurable [GT06]. Confirms [Han00-95]. conic [LWL09]. Conquers [Hin00d]. Consensus [KK08a, KK08b, KK08c]. consequences [GRJ07]. consideration [MPPB04]. Considerations [Zun00, PW06]. consolidation [KL07]. constrained [Tse07]. Construct [Gol01]. constructs [MSLD09]. Containing [PF08a, PF08b, PF08c]. containment [BIC08a, BIC08b, BIC08c]. Content [Wri01, GSK09]. Contents [Ano06p, Ano06q, Ano06r, Ano07u, Ano07v, Ano07w, Ano07x, Ano07y, Ano07z, Ano07-27, Ano07-28, Ano07-29, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano08t, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h]. contest [MC03b]. context [Hin03c, KHKK06, MP07]. context-aware [KHKK06]. contextual [JS03, KKK05]. Continue [Han00y]. Continuous [FBV06, FvS05]. contract [DMWS09]. contributor [Sch05c]. contributory [GYLC09]. Control [BE01, WS02, vSvS06a, CC02, CE04, HW03, JCKL04, KHKK06, OKE09, PE09, Pur02, RK06, Sch04l, dSSW04, SV07, TE06, YL04, Zho02, ZL03, vdHvS03]. controllers [Asl05]. Controlling [Wri01]. Controls [BvS00, TE01, PE09]. Conundrum [Han00g]. convenience [WDCJ09]. Convention [Pou01b, Pou01a]. converged [PF09]. Convicted [Han00-48]. cooperative [PF08a, PF08b, PF08c]. coordinates [CLK09]. copy [Ano05c]. copyright [HLC07, KA09]. Corporate [Kov01a, Web00, Zen01b, vS01a]. correlation [LCK^{+}06, SG09, ZLCF08a, ZLCF08b, ZLCF08c]. cost [FBB06, Kon09, Pol05, SK03]. cost$ [Hin03h]. Could [Han00-44, Han00-97]. Council [Pou01b, Pou01a]. counter [GKS^{+}06]. countermeasures [PML09]. country [BTY^{+}07]. courses [SS07]. Court [Han00-95]. coverage [Ano02d, Ano02e, Ano02f, MH02a, MH02b]. Cracker [Han00g]. Crazed [Han00-102]. creaking [Hin03e]. Creates [Han00b, Han00-50, Han00-67, Han00-86]. Creative [Han00p]. credential [SV07]. credential-based [SV07]. credentialing [GKH09]. credentials [Sch04d]. Credit [Han00b, Han00g, Han00-81]. Credit-Card [Han00-81]. Crime [Beq01a, Col00, Pou01a, Beq02a, Beq02b, Pou01b]. Crimes [Phi01b, Web00, DM01]. Criminal [Han00-51, Rog03]. Crisis [Han00-104, Han02]. Criteria [Pal08a, Pal08b, Pal08c]. Critical [Beg08a, Beg08b, Beg08c, Rat01, FL05, Her09a]. Criticized [Han00-89]. crop [Hin02e]. cross [KJKV09]. cross-site [KJKV09]. crosses [Sch05a]. Cryptanalyses [HW03, WL05]. Cryptanalysis [CC01, GLMZ08a, GLMZ08b, GLMZ08c, Pha06, WLT03, YW04]. Cryptographic [Gol01, SVW00, HW03, TNG04]. Cryptography [TMM01, AN03, CJ03, Hei03, Lin01, Pip03, Sch02d, Sch04e, Van03]. cryptosystem [CCH03, YC09]. CSP [SBS09]. cues [KPsH^{+}08a, KPsH^{+}08b, KPsH^{+}08c]. culture [Fur07d, RMC07, vSvS04b]. Cumulative [LG04]. Current [Phi01b, CH06]. curve [CCH03, LWL09, Van03, YC09]. Cyber [Beq01a, Beq01b, Hin03c, Pou01b, Pou01a, CITN07, Hin01e, Hin03d]. Cyber-Crime [Pou01a]. Cyber-Space [Beq01b]. Cyber-terrorism [Hin03c, Hin01e]. Cyberconditions [Han00-50]. Cybercrime [Han00-36, Han00-34, Han00-47, Han00-88, Han00-87, Phi01b, Hin03d]. Cyberspace [Car01, Han00y]. Cyberstalking [Han00m]. Cyberterrorism [GF02]. Cyberthreats [Hin01c]. Cybertools [Han00-94]. Cycle [vSvS06a]. Czar [Han00-60].

D [NTH04]. damage [Hin01e, dSSW04]. Data [Beq01b, Han00-52, Han00-82, Pou02a, SK03, Zun00, AMB06, BS07a, BS07b, BS07c, CO09, FvS05, GC09, KPsH^{+}08a, KPsH^{+}08b, KPsH^{+}08c, KN03, OZ07, WGZY06]. Database [Han00-88]. databases [Kam09]. Dataflow [FKSF01]. date [CL03]. date-attachment [CL03]. DDOS [Li04, Li06, Han00n, Han00-83, LS05]. deadly [Hin02a, vSvS04a]. Dealing [BC05, JS03]. Death [Hin02a]. December [Ano00i, Ano01-45, Ano03-39, Ano04o, Ano07-37, Ano07-38, Ano07-39, Ano08-27, Ano08-28, Ano08-29]. Deceptions [Coh00, CK03]. decision [JST^{+}07, KL07, ZHH09]. Decode [Gra01]. Decoding [Bra07d]. deduction [CAT05]. Defaced [Han00e]. DefCon [Han00o]. defective [HD05]. Defence [Han00n]. defences [Rob07]. Defending [LS05, WZM09]. Defense [Han00-88, BNSZ09, CLC04, CS09]. defenses [Her09b]. Defensive [Coh00]. definition [And03, TvS05]. Déjà [Hin02c]. Delay [Han00-47, YZPL09]. delegated [GRNR09]. delegation [MW06]. Denial [CBR06, Han00-63, CLC04, ZC05]. Denial-of-Service [CBR06]. Department [Han00-88, Han00-89, Han00-51]. dependability [Hun02]. dependent [SK03]. deploy [And05]. deploying [GLK08a, GLK08b, GLK08c]. deriving [vdHvS03]. Design [BFG^{+}05, DS09, ESZ06, Sam09, BP03, CAT05, CJC04, Hof04, KVD06, Lea04, MGMSML^{+}09]. desktop [BFC09]. Detect [Han00-79, YH07a, YH07b, YH07c]. Detecting [HC03, KLL09, Beg08a, Beg08b, Beg08c, GVL09, Zen01a]. Detection [Bar01a, BCV01, GFV05, AJS06, Ano03l, ASS^{+}06, Beg04, Bis02, BvS03, CBR06, CAT05, CP03, CC04, CLK09, DGY^{+}05, ES09, GBBS08a, GBBS08b, GBBS08c, GTDVMFV09, Ihe04, JST^{+}07, JSD05, KC05, LCK^{+}06, LG07a, LG07b, LG07c, LWT^{+}09, LV02, MP08a, MP08b, MP08c, MP07, OL03, SPP07a, SPP07b, SPP07c, SYL09, TC05, VE02, Wan05, WGZY06, YH07d, YZ07]. detector [HFC^{+}08]. deterrent [PE09]. develop [And05]. Developer [Han00f]. developing [BTY^{+}07]. Development [JS02, WS02, CS09, VFMP05]. Devices [CFRR02, CF07, sHCP09, OKH07, YC09]. DFANS [LL09]. dictionary [GKS^{+}06]. Differentially [HMDC02]. difficult [Pur02]. Digest [Ano02d, Ano02e, Ano02f, MH02a, MH02b]. Digital [Han00p, Han00-64, LWL09, TMM01, YH01, ZCW04, AH09, Bra07d, HLC07, KK02, LG04, LG09, Sch06c, WC05]. Dilemmas [Sch06c, Sch05k]. dimensional [GKH09]. Direct [vSvS06a]. Direct-Control [vSvS06a]. directed [OBB05]. direction [Ste04, Woo04]. direction-based [Woo04]. Director [Han00-51]. disaster [Hin01d, Sch05g]. Discipline [vS01b]. Disclosures [Han00-69]. discovery [Luo06]. Discrete [SC02, LTH03]. discussion [FBV06]. disk [LG06]. dispatch [DR04]. Disposable [Han00b]. dissected [Nik02b]. dissemination [WZM09]. Dissimilar [Gup00]. distinguishing [JS03]. distributed [AP04, Asl04c, Asl05, BAS09, CLC04, CCC09, Mac05, MP09, OCB09, TM09, WH04]. distribution [Asl04b, PF08a, PF08b, PF08c, PRZ03, PRZ04, Sch06d, WH04, YWB^{+}04, YZPL09]. Divide [Han01z, AH09]. DNS [Her09a]. DNS-based [Her09a]. Do [Hin00a, Nos00a, Woo01, HE02, Sch05i]. Document [BE01]. Document-centric [BE01]. Documents [PMRZ01b, HC04a]. DOD [Han00q, Ano01-49]. DOE [Han00-78]. Doesn’t [Han00t, Han00-75]. doing [DMWS09, Pol05, RSS02]. Domino [Han00-55]. dominoes [Hin03e]. double [CL05]. double-exponentiation [CL05]. Down [Han00-70, Han01d]. drives [LeV05]. DRM [BSC09]. Dropped [Han00-80]. Drugs [Han00-49]. Dual [HLC07, Sam09]. dual-computer [Sam09]. Dual-wrapped [HLC07]. Due [Han00-57, vSvS06b]. Duh [Han00-43]. Dutch [Han00-44]. duty [MPPB04]. Dynamic [Bar00a, PdVGA08a, PdVGA08b, PdVGA08c, DR04, HW03, KHKK06, KK07, LZXW09, MP08a, MP08b, MP08c]. dynamics [CY05, sHCP09, jLC07, YC04]. dynamics-based [sHCP09, jLC07].

E-Biz [Han00-103]. e-cash [CL03]. E-Commerce [Han00r, TMM01, Wri01, Nab05]. e-learning [GRNR09]. E-mail [Han00-92, Han00-95, Han01c, Han00s]. e-mails [LG09]. E-Security [Han00-96]. e-services [KHKK06]. E-tailers [Han01d]. e-voting [BFG^{+}05, Gri02]. Each [PMRZ00a]. ebanking [WDCJ09]. economic [Mat03, TS05a]. economy [McK02b]. Edinburgh [Eva01]. editor [Ano02n, Ano03k, Ano03l, Ano03m, Sch01b, Sch02c, Ano01l, Car00, Han00z, Han00-27, Han00-28, Han00-29, Han00-30, Han00-31, Han00-32, Han00-33, Han01f, Han01g, Han01h, Han01i, Han01j, Han01k, Han01l, Sch02b, Wol08h, Wol08i, Wol08j]. editor-in-chief [Ano02n, Ano03k, Ano03l, Ano03m, Sch01b, Sch02c, Wol08h, Wol08i, Wol08j, Sch02b]. Editorial [Ano06m, Ano06n, Ano06o, Ano07l, Ano07m, Ano07n, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano07t, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, GK09b, GL09, GF09, GK09a, Wol08a, Wol08b, Wol08c, Wol08d, Wol08e, Wol08f, Wol08g, Wol09]. Education [Hor01, DSX06, RM08a, RM08b, RM08c]. Effect [PPBH00, Ihe04]. Effective [BvS00, BvS03, OZ07]. effectiveness [TKKK05]. effects [Pou03a]. efficacy [RKR09]. Efficient [CC03, CCH03, CJT02, CP03, HR08a, HR08b, HR08c, Jua04, SW06, Tsa08, WH04, ÁP03, DS09, JPL05, LL09, SJ07, YWB^{+}04].

Eldorado [Beq01b]. election [Sch07d]. elections [Lia04]. electro [HYJ^{+}09]. electro-magnetic [HYJ^{+}09]. Electronic [CDD^{+}02, GKL^{+}00, HSH00, Ja02, PMRZ01b, CLX05, Lia04, Luo06, SP06, SKM08a, SKM08b, SKM08c, TS05b, Wol03f, Zuc04, Zuc07]. Eliminates [Han00-46]. Elizabeth [Hil02b, McK03b]. elliptic [CCH03, Van03, YC09]. email [Her09a, Sch07a]. Embedding [NTH04, CO09]. emergence [Pou02b]. Emerging [Hor01]. Empirical [KC05, LCLL09, LSK09, MK04, Sch07a, SKBM06]. Enablement [Sho00]. encoding [ÁP03]. Encountering [Wol03b]. Encrypted [MMZ00, AMB06, LHL04]. encryption [SK03, Wol03b]. end [AMB06, FJK06, LBE08a, LBE08b, LBE08c, MF05, PF09, RKR09, RMC07, SSMJ05]. end-to-end [AMB06, LBE08a, LBE08b, LBE08c, PF09]. end-user [RMC07]. end-users [FJK06]. Enemy [Zen01a, DM01, Han00a]. Energy [Han00-89]. Enforcement [Eva01, Sch02e, Dav02b]. Enforcers [Han00-94]. Enforcing [HSKL08a, HSKL08b, HSKL08c]. Engineering [LLL02, Boo06, Lea03a, Lea04, SKBM06, Zuc04]. engines [AN04, Sch05c]. English [Che04]. Enhanced [LHL04, CZ03, McK04, WLT03, YWB^{+}04]. enhancement [ESZ06, FLZ02, SLH03]. enhancing [LCGS03]. Enigma [Thi03]. Enough [vKH00, Bha03]. ensemble [CAT05]. Enterprise [AC08a, AC08b, AC08c, McK02a, She00, Boo06, Dou03]. entity [CSL09]. Environment [Bar00a, KK00, Che04, DKS07, LG06, LCC04]. environments [BNSZ09, KK07]. Erratum [Ano02g, McK04]. error [LSK09]. Escalate [Han00s]. Escrow [VBD01]. establish [LPF06]. estimated [ZHH09]. estimation [CC04]. ethical [TE01]. Europe [Han00-87, Han01e, Han01q, Pou01b, Pou01a]. European [BEP04, Han00t, Hin03f, Pou01c]. evaluate [MC03b, Pal08a, Pal08b, Pal08c]. evaluates [Han00-61]. Evaluating [CB01, MF01, PK07]. Evaluation [BvS00, EvS00b, MFDVGT08a, MFDVGT08b, MFDVGT08c, CZG^{+}05, KC05, MGMSML^{+}09]. Events [Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano02m, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano04e, Ano04f, Ano04g, Ano04a, Ano04h, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano06a, Ano06d, Ano06f, Ano06j, Ano06k, Ano06l, Ano07k, Ano01j, Ano02c, Ano03a, Ano06c, Ano06b, SBS09]. Everyone [Han00-64]. everything [FVA05]. Evidence [All01, Ano03-41, Wol03c, Wol03d, Wol03e, Wol03f]. Evolution [Bar01a, Hin03g, LeV05]. examination [SKBM06]. Examining [BFC09]. example [GRJ07]. Exchange [KKHR01, CL06, GLMZ08a, GLMZ08b, GLMZ08c, HAGTA08a, HAGTA08b, HAGTA08c, LHL04, LFHT07, LC07, WLH06]. execution [BNSZ09]. Exhibitors [Ano01k]. Existing [GM00]. exit [GYLC09]. Expanded [Han00-91]. Expected [RR06]. experience [BFC09]. exploit [TXL09]. exploit-based [TXL09]. exploration [OKE09]. exploratory [RM08a, RM08b, RM08c]. Explorer [Sch05k]. exponentiation [CL05]. Express [Han00b]. expressions [JCKL04]. Extending [RF01, RMC07]. extension [KMWD06]. Extensions [BM00]. Extortion [Han00u]. extra [vSM04]. extraction [Kum09].

face [GKH09]. FaceHashing [TNG04]. Faces [Han00-104]. factor [Sch05h]. Factors [Nos00a, KCC09]. FACTs [Wil00]. Failures [Hin00e]. Fair [KKHR01, HAGTA08a, HAGTA08b, HAGTA08c, LCC04]. fairly [Sch06g]. Far [Han01r, RSS02]. Fast [CLK09]. Fault [BP03, Seb06, TM09]. fault-tolerant [Seb06]. Fear [Hei03]. feasible [HFC^{+}08]. Feature [CAT05, LWT^{+}09]. Features [Wil00]. February [Ano00b, Ano03-40, Ano04i, Ano05d, Ano06t, Ano07-32, Ano09w]. February-March [Ano09w]. Feeling [Han00v]. Few [Han01r]. Fighting [Han00-91, Han01e, Zen01a]. file [LG06, dSSW04, WZM09]. Files [Han00w, SKBM06]. filter [GT06]. Filtering [JCKL04, CH06, Ede03, PF08a, PF08b, PF08c, Woo04]. filters [GVL09]. Financial [Phi01b, ES09]. Finds [Han00q]. fingerprint [LS05]. fingerprinting [IKE06]. Firewall [FKSF01, Han00-39]. Firewall-1 [Han00-39]. Firewalls [FKSF01, KFS^{+}03, Woo04]. Fireworks [Hin00b]. First [Han00x]. fitting [Sch04i]. five [And05]. Flash [Dav03]. flawed [MvO09]. Flaws [Han00r]. flexible [CL03, HLL05]. flood [Li04, Li06]. flooding [GVL09]. flow [MP08a, MP08b, MP08c, SAF09]. flows [CP03]. focus [McK02a]. focused [DKS07]. forecasting [VE04]. Forensic [Wol03e, BS07a, BS07b, BS07c]. forensics [Luo06, Rog03, RS04, Wol03a, Wol03f]. Formal [Ung04, SAF09]. Formalized [BvS00]. Format [Gup00]. Formats [GM00]. Former [Han00-51]. Formulation [PPBH00]. forthcoming [Ano01j, Ano02c, Ano03a]. Forum [Eva01]. Fourth [vS06]. Framework [EvS00a, FKSF01, GM08a, GM08b, GM08c, TE01, ASS^{+}06, Ayo06, BAS09, HLL05, LSK09, MP07, PvS04, SG09, SP06, Sch02a, SAF09, SLG04, SV07, Trè03, Zuc07]. France [Han00-97]. Fraud [Beq00a, Han00-81, ES09, Hil02b, Pol05]. Freeware [Han00f]. freewheeling [CC03]. frequency [Rob06, WGZY06]. frequent [SG09]. friendly [WC03, YW04]. front [Ede03]. FTC [Han00-75]. FTKM [Seb06]. Fuel [Han00-85]. fully [OCB09]. function [Kum09, Tsa08]. Functional [KMZ00, LK07]. Functions [MZ00, DS09, PW06, YL04]. Further [JPL04]. fusion [CZG^{+}05]. future [Ano05q, Hin03b, McK02b, RS04, Sch03b, Sch06f, SK03]. fuzzy [BvS03].

G8 [Han00-34]. game [CITN07, HCBLETRG06]. games [HCBLETRG06]. gap [BTY^{+}07, KS06, Sch02d, Sch04e]. gaps [Ber07]. Gartner [Ano03l]. general [BTY^{+}07, HCBLETRG06, Lia04]. Generalization [LTH03, HC04b]. generalized [LWL09]. generate [TXL09]. Generation [PMRZ00a, PMRZ01a, CH06, CH04, TNG04, Van03]. generators [SWA^{+}04]. Generic [HSH00]. GeoCities [Han00-35]. Get [Han00-45, Han00k, Han00p]. Gets [Han00d, Han00n]. Getting [Han00-36, Han01z, Sch04h]. Giga [Dav02a]. Global [DeM02, Eva01, Kov00b, Kov00c, WZ00, GBBS08a, GBBS08b, GBBS08c, Hun02]. Go [HCBLETRG06, RSS02]. Goes [Beq01a, Han00c]. Going [Han00u]. good [McK03a]. Governance [vS01a, vS05b, vSvS06a, MC03a, PvS04, vS05a, vSvS06b]. Government [Han00-67, Kja06, Han00-90, Han00-91, Han00-92, Han01n, Sch02g, Sch04h, Sch05b]. Government- [Sch02g]. Governments [Pou03b]. grand [Han00-36]. Grant [SJ07]. graph [ZLCF08a, ZLCF08b, ZLCF08c]. Graphical [Pur01]. graphs [CK03]. Gray [Kov00c]. Great [Han01z]. greater [LeV05]. Greylists [GT06]. Grief [Han00j]. Group [Han00-47, Han00-69, KMWD06, LZXW09, NCS06, Tse07]. group-based [LZXW09]. Grow [Han01q]. Growing [Han00-42, Sch05c]. Guaranteed [PMRZ00a]. guard [McK03a]. Guessing [PPBH00]. GUI [LG09]. Guide [Ano01m, Ano01n, Ano01o, Ano01p, Ano01q, Ano01r, Ano01s, Ano01t, Ano02o, Ano02p, Ano02q, Ano02r, Ano02s, Ano02t, Ano03n, Ano03o, Ano03p, Ano03q, Ano03-43, Ano06g, Ano06h, Ano06i, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano09i]. Guidelines [MMZ00, Zen01b]. Guilty [Han00h, Han00-59].

Hacker [Han00-48, Han00-37, Han00-44, Han00-38, Han00-59]. Hackers [Han00-39, Han00-43, Han00-40, Han00-81, Sch02f]. Hacking [Han00-85]. Hacktivism [Han00-41]. Halfpennies [Hin00b]. Hand [Han01m]. Hand-held [Han01m]. Handling [MPD06]. handmaiden [Beq02b]. hands [SS07]. hands-on [SS07]. Happens [Han00-82]. Harbor [Han01r]. Hard [Han00-42, McG02]. hardening [MLD08, MSLD09]. Hardware [Han00-52, DS09, HSKL08a, HSKL08b, HSKL08c]. Hash [CBB05, Wu05, DS09, PW06, Tsa08, YL04]. Hash-chaining [CBB05]. Hatches [Han01d]. Hazards [Han00i]. head [RFR07a, RFR07b, RFR07c]. Header [MW00, ASP03, WC07]. headlines [Sch03e]. Health [Han00-42, SE02]. Healthcare [JS02]. healthy [WZM09]. heats [Sch03l]. heck [DMWS09]. Heel [Hin00a]. held [Han01m]. Help [Han00-36]. here [Sch05i]. hidden [CP03]. Hiding [YH01]. Hierarchical [CC02, EvS00a, Lin01, CJ03, ZL03]. hierarchies [MPPB04]. hierarchy [Asl04c, Asl05, HW03, YL04, Zho02, hZxZtZ06]. high [JSD05, SK03]. high-speed [JSD05]. higher [RM08a, RM08b, RM08c]. Highlighted [Han01n]. highly [LL09, Sam09]. Hire [Han00-48]. histograms [CO09]. Hit [Han00-56, Sch03f]. Hits [Han00x, Han00-83]. hoaxes [Hin02f]. hoc [KVD06, KVD07, OCB09, TC05]. hole [Sch04i]. Holes [CG00, Han00-55]. Holistic [Zuc04, Zuc07]. Home [Han00-43, Han00-79]. Homeland [Pou02d]. honeypot [ASS^{+}06]. Host [Hof05, LCC04]. hosting [Sch03f]. hot [Sch06e]. House [Han00-98, Han00-99]. However [Han00-44]. Huffman [GYLC09]. Huffman-based [GYLC09]. huge [Sch04f]. Human [KCC09, JEL03, LSK09, RFR07a, RFR07b, RFR07c, Sch05h]. Hurst [Li06]. Hwang [YRY05]. Hybrid [CBB05, VBD01, ASS^{+}06, Asl04a]. hypergraph [JL07a, JL07b, JL07c].

I-Worm.Lentin [Nik02a]. IBC [Ano06j, Ano06k, Ano06l, Ano07k]. ICT [BTY^{+}07, DKS07, KK08a, KK08b, KK08c]. ID [Han00-46, YC09]. ID-based [YC09]. idea [Han00-36]. Identification [MK06, Pal08a, Pal08b, Pal08c, OG07, Rob06, WH04, YWB^{+}04]. identifiers [NTH04]. identifying [Li04]. identities [IKE06]. Identity [Han00-104, DP06, HR08a, HR08b, HR08c, YC04]. identity-based [HR08a, HR08b, HR08c]. IDS [ZHH09]. IDS-estimated [ZHH09]. If [Hin01d]. IFC [Ano06m, Ano06n, Ano06o, Ano07l, Ano07m, Ano07n, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano07t, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o]. IFIP [Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-28, Ano01-29, Ano02u, Ano02v, Ano02w, Ano02x, Ano02y, Ano02z, Ano02-27, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano03w, Ano03x, Ano03y, Ano09p, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u, Ano09v, Fur07d]. IFIPSEC [Ano06e]. II [Hil02b, McK03b, Kov00c]. iii [Ano06p, Ano06q, Ano06r, Ano07u, Ano07v, Ano07w, Ano07x, Ano07y, Ano07z, Ano07-27, Ano07-28, Ano07-29, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano08t, Kov00a]. IIS [Han00-58, HFC^{+}08]. image [GSK09, HLC07, WC05]. imbalanced [Tse07]. imitation [hZxZtZ06]. immunology [Hof04]. Impact [Pou02a, Mat03, Sch07a, Wia05]. impacts [ZHH09]. impaired [BA04]. Implementation [MMZ00, TE01, BFG^{+}05, DS09, HMDC02, Sam09]. Implementing [Bar00a, Dou03, Nos00a]. implications [Ayo06, BLP05, Hof04]. impostor [jLC07]. impostors [Hin01d]. improve [Pou03b]. improved [NTH04]. Improvement [HC04b, HLL05, KPsH^{+}08a, KPsH^{+}08b, KPsH^{+}08c, JPL05, THY04, YRY05]. Improvements [ZC05, HW03]. Improving [Lea03b, Pur04, ASS^{+}06, Fur07e]. inadequacies [PE09]. Incalculable [Hin01e]. Incident [Han00-67, MPD06, Sch04a, Sch05g, ZHH09]. incidents [Kja06, Wia05]. Include [RF01]. Increasing [Han00-41]. Incremental [vSvS01, PW06, SG09, SYL09]. independent [GSK09]. Index [Ano00a, Ano00j, Ano01a, Ano01-48, Ano02b, Ano02-46, Ano03b, Ano03c]. indicators [TC05]. Indicted [Han00-77]. Individual [BCC02]. Industry [Han00-54, KS08a, KS08b, KS08c]. Infection [CY05, hZxZtZ06]. Infects [Han00w]. influence [RKR09]. Information [All01, BvS00, BB05, Ber07, Dhi01, DEE09, EvS00a, EvS00b, Fin00, Fur07d, GvS08a, GvS08b, GvS08c, Han00-35, Han01n, HE02, JS02, KKK05, KMMB09, Kov00b, Kov00c, Kov01a, Kov01b, KJ02, KS08a, KS08b, KS08c, MMZ00, MS00, Nos00a, OKH07, Pou01c, Rat01, RM08a, RM08b, RM08c, SPLS01, Sch03a, SVW00, SE02, TvS05, TM00, TE01, WS02, Web00, Wia05, Woo01, YH01, vS00, vS01a, vSvS01, vS01b, vS05a, vS05b, vS06, vSvS06a, vSvS06b, Alb07, AH09, And03, AC08a, AC08b, AC08c, CZG^{+}05, DF06, Ede03, FvS05, GvS05, HYJ^{+}09, KS05, KJ01, KCC09, KK06, LeV05, MP08a, MP08b, MP08c, MC03a, PK07, PvS04, RKR09, RR06, Sch02d, Sch02g, Sch04e, Sch04f, Sch04h, Sch05b, Sch06b, Sch06d, Sch06g, SS07, SAF09, TKKK05, Trè03]. information [TS05a, VE03, VFMP05, VvS04, vdHvS03, vSvS04a, vSvS05]. infosec [Sch05j, Sch07c, Sch07d, KJ01, KJ02, Sch03e, Sch06f, Sch05i]. Infrastructure [GKL^{+}00, GM00, LCGS03, Hin03e]. Infrastructures [Bar00a, BCR03, Rat01, FL05, Hun02, LOP04, PF09]. inherent [Bra07a, Bra07b, Bra07c]. initiation [YWL05]. injection [MS09]. Insecure [Han00-66]. insecurity [MSZ05]. Insider [MF01, MF05, Sch02a, TKKK05]. Insights [Bla02]. instructional [DSX06]. Insurance [Han00-37, RF01]. integral [Trè03]. integrated [PML09]. Integrating [LLL02]. integration [HC03]. integrity [FvS05, GC09, Kam09, dSSW04]. Intel [Han00-46]. intentions [FGG^{+}04]. Interactive [BC05]. Interesting [Han00-49]. interfaces [JEL03]. interfere [PK07]. intermediary [CSL09]. Internals [FKSF01]. International [AN03, Ano01-33, Ano01-37, Beq02a, Beq02b, HE02, Ano01-30, Ano01-31, Ano01-32, Ano01-34, Ano01-35, Ano01-36, Ano02-32, Ano02-33, Ano02-34, Ano02-35, Ano02-36, Ano03z, Ano03-27, Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano03-32]. Internet [BEP04, Beq00a, Beq00b, Beq01c, CY05, CJC04, DP06, FL05, FBP07, FTP08a, FTP08b, FTP08c, GRJ07, Han00-47, Han00-63, Han01b, KFS^{+}03, Kon09, KB00, Mac05, PF08a, PF08b, PF08c, Phi01a, Phi01b, QW05, Sch07a, Sch03b, Sch05k, Web00]. internetwork [CS09]. internetwork-centric [CS09]. Interoperability [Gup00, KK00, WZ00, LCGS03]. Interpolations [SC02, WL05]. Interpreting [GvS08a, GvS08b, GvS08c]. Introduction [Ano03-41, WZM^{+}00]. Intrusion [Bar01a, BCV01, GFV05, Sch04g, SPP07a, SPP07b, SPP07c, AJS06, Ano03l, ASS^{+}06, Beg04, Bis02, BvS03, CAT05, DGY^{+}05, FGG^{+}04, GBBS08a, GBBS08b, GBBS08c, GTDVMFV09, HC03, Hof05, HFC^{+}08, Ihe04, JSD05, LCK^{+}06, LG07a, LG07b, LG07c, LWT^{+}09, LV02, MP08a, MP08b, MP08c, OL03, SYL09, TC05, VE02, Wan05, WGZY06, YH07a, YH07b, YH07c, YH07d]. intrusions [Beg08a, Beg08b, Beg08c]. intrusive [AMB06, RFR07a, RFR07b, RFR07c]. Investigating [AMB06]. investigation [MK04]. Investigative [Nov07]. investment [Ihe04]. investments [RR06]. Invisible [Zen01a]. IP [ASP03, PF09, SSF^{+}09, YH07a, YH07b, YH07c]. IPSEC [MW00]. IPv6 [JHgJ05]. Iraq [Mat03]. ISAKMP [BM00]. ISC [DeM02]. Isn’t [Han00-49]. ISO [KS06, vS05a]. ISO17799 [TKKK05]. ISPs [Han00-95]. Israeli [Han00y]. ISRAM [KS05]. Issue [WZM^{+}00]. Issues [DFSC09, Han00s, SPLS01, Sch06d, ILW06]. IT-Security [HM04].

JADE [VSR07]. Jailed [Han01c]. JamFest [Han00o]. January [Ano00c, Ano01-39, Ano03-34]. Japan [Han00-50]. Java [HS05, PE06]. Johnny [Her09b]. join [GYLC09]. join-exit-tree [GYLC09]. July [Ano00f, Ano01-41, Ano01-42, Ano03-36, Ano04l, Ano06w, Ano09y]. June [Ano02-37, Ano08-30, Ano04k, Ano05f, Ano06v, Ano07-34, Ano09x]. junk [Hin02f]. Just [vKH00, Hin01d, Pol05]. Justice [Han00-51].

kernel [SPP07a, SPP07b, SPP07c]. Key [BM00, Bar00a, BCR03, CB01, CJ03, Fin00, GKL^{+}00, Gup00, GM00, JL07a, JL07b, JL07c, KMZ00, LCGS03, MK06, MW00, MZ00, SC02, Sho00, SVW00, VBD01, Wil00, WZ00, WZM^{+}00, Zun00, Asl04b, Asl04c, CL06, CMY07, GYLC09, GLMZ08a, GLMZ08b, GLMZ08c, HW03, HLTJ09, HLL05, Jua04, LHL04, LFHT07, Lin01, LC07, PRZ04, RM05, Seb06, SW06, TNG04, Tse07, WL05, WLH06, WH04, YWB^{+}04, YC09, Zho02]. Keyjacking [MSZ05]. keylogging [GRJ07]. Keys [Han00-66, PMRZ00a, LHL04]. KEYSTONE [GKL^{+}00]. Keystroke [sHCP09, YC04, GS03, KPsH^{+}08a, KPsH^{+}08b, KPsH^{+}08c, jLC07]. Kit [Han00-37]. KLEZ [Arm02b]. Klez.H [Nik02b]. knees [Bha03]. KNN [LG07a, LG07b, LG07c]. Know [Hin00a, KJ02, KJ01]. known [THY04]. KRA [WZM^{+}00]. Kwak [KMWD06].

Lab [HM04]. label [BS07a, BS07b, BS07c]. laboratory [Wol03f]. LAN [LFHT07]. language [HC04a]. Laptop [Han00-52]. Large [Han00-53, PF08a, PF08b, PF08c, SYL09]. large-scale [PF08a, PF08b, PF08c, SYL09]. last [BNSZ09]. Laundering [Phi01a]. Law [Eva01, Han00-93, Han00-94, Han01b, Sch02e, BEP04, GRJ07, Hin03d]. Laws [Han00-36, Han00-91]. Layered [KVD06, KVD07]. Leading [CK03]. leakage [HYJ^{+}09]. Learned [Hin01f, PE06]. learning [GRNR09, LG07a, LG07b, LG07c]. Lee [YRY05]. Legal [All01, Han00-45, AN03, BAS09, GvS08a, GvS08b, GvS08c]. Legislated [Han00-45]. Legislation [Pou02a, Hin03f]. legitimate [BA04]. length [MGMSML^{+}09]. Lessons [Hin01f, PE06, Beq02a]. Letter [Car00]. level [Asl05, KMMB09]. LFSR [DS09]. Li [HC04b]. Liable [Han00-95]. libel [Pou03c]. Licence [Cus01]. lies [Pou03c]. Life [Hin00c]. Lifecycle [Ber07, LLL02]. lightweight [LWT^{+}09]. Like [Han00t]. likely [Sch07d]. limit [MGMSML^{+}09]. limitations [HKP02, LG09]. line [BNSZ09, Sch05a]. Link [Hin01j]. Linux [Han00-62]. LIRVA [Fis03]. List [Han01a]. Literature [Ano02a, Ecc01a, Ecc01b, Mey02, Nic00a, Nic00b, Nic00c, Nic00d, Pal00a, Pal00b, Pal00c, PM01, Pal01a, Pal01b, Pal01c, Pal01d, Pal01e, Ecc01c]. local [Ede03]. Location [MS09, JHgJ05, RK06]. location-based [RK06]. Location-specific [MS09]. log [SKM08a, SKM08b, SKM08c]. Logarithms [SC02, LTH03]. logging [RSS02]. logic [BvS03, Nab05]. Logical [Asl04b]. login [GRJ07]. logistic [Wan05]. London [Hil02b, McK03b]. long [LG04]. long-term [LG04]. Look [Han00-42]. Lose [Han00-82]. losing [Sch04l]. loss [BC05]. lossy [Asl04a]. Lotus [Han00-55]. Love [Han00-40, Han00-80, Hin00d]. Low [Han01-27, MFDVGT08a, MFDVGT08b, MFDVGT08c, SK03, WLH06]. low-rate [MFDVGT08a, MFDVGT08b, MFDVGT08c]. LRD [Li04]. Lycos [Sch05a].

M [CMY07]. M-CLIQUES [CMY07]. Mafiaboy [Han00h, Han00-56]. magnetic [HYJ^{+}09]. mail [Han00-92, Han00-95, Han01c, Hin02f, Mac05, Han00s]. mails [LG09]. Making [Fur07e]. malicious [Kum09]. Malware [Ano05o, Ano05p, Ano05m, Ano05n, Ano06s, Ano06-28, Ano06-29, Ano06-30, Ano06-31, Ano06-32, Ano06-33, Ano06-34, Ano07-30, Ano07-31, Ano07-42, Ano07-41, Bra06a, FBB06, Kon09]. Managed [McK02b, Sch01a]. Management [CM03, EvS00a, EvS00b, Fin00, GvS05, GM08a, GM08b, GM08c, Han00-37, MS00, SC02, TMM01, And05, BTY^{+}07, Bra07d, BAS09, DP06, GYLC09, Han02, HLL05, Ihe04, KK02, KS08a, KS08b, KS08c, LSK09, Luo06, MC03b, Pur04, Sch06c, Seb06, SKM08a, SKM08b, SKM08c, Trè03, Ung04, Zho02, Zuc07, vSvS04a, vS05b, Beq00b]. manager [BSC09, HS05]. managers [AH09, Beq03]. Managing [Nos00b]. mandatory [RK06]. Mapping [Hin03b]. March [Ano00d, Ano02-43, Ano04p, Ano05k, Ano06-27, Ano07-40, Ano08u, Ano08v, Ano08w, Ano09w]. Marines [Han01x]. marked [Ano05c]. Market [GKL^{+}00, RF01, HD05]. Market-Insurance [RF01]. Markov [CP03]. Masquerade [JST^{+}07, KC05]. Mass [Han00-85]. massive [BS09]. Matching [RM05]. Mathematical [Coh00]. matrix [GC09]. maximizing [GSK09]. May [Ano01-38, Ano03-35, Ano04j, Ano05e, Ano06u, Ano07-33, Ano08-30, Ano09x]. May-June [Ano09x]. McKenna [Ano02-31, Ano02f]. Mean [Woo01]. measure [SPP07a, SPP07b, SPP07c]. measures [GSK09, GRJ07]. Measuring [AMR07, ZHH09]. mechanism [CL06, JHgJ05, KK07, MGMSML^{+}09, RFR07a, RFR07b, RFR07c]. Mechanisms [Gup00, Pal08a, Pal08b, Pal08c, CLC04, Her09a, LWT^{+}09, Pip03, RM05]. media [CBB05, Sch03a]. media-streaming [CBB05]. meet [Hin01d]. memory [HSKL08a, HSKL08b, HSKL08c]. Merchant [HSH00]. message [ZCW04]. Messages [Han00-95, WC07]. Messaging [CG00]. metamorphosis [Bra06a]. Method [LL02, PMRZ01a, KS05, KS06, OL03, PML09, YZ07]. methodology [HCBLETRG06]. Methods [CFRR02, PZ00, PMRZ01b, PRZ03, SPLS01, Zen01a, Zun00, ES09, HMDC02, Ihe04, KL07, LFHT07, PRZ04]. Metrics [Han00-90]. Microsoft [Bra06b, Han00v, Han00-44, Han00-57, Han00-58, Sch03e, Sch05k, Sch06e, Sch07e]. middleware [LCGS03]. mighty [Bha03]. migration [KN03]. Millennium [Hin00e]. MIME [LG09]. Mining [YH07a, YH07b, YH07c, SG09, SYL09]. Minix [DSX06]. mirroring [TM09]. Mischiefs [Hin00g]. Misdemeanours [Web00]. missed [PE06]. misuse [MF01]. Mitigation [DR04]. mix [JHgJ05]. mix-based [JHgJ05]. Mobile [Bar00b, Bar01b, CFRR02, Han00-38, HS00, JHgJ05, Sch07b, BSC09, BFC09, CF05, CF07, DR04, DP06, sHCP09, KVD06, KVD07, LCC04, MP07, OCB09, SF04, TC05, YC09]. Model [AGAS01, AOG^{+}02, FKSF01, HSH00, KMZ00, KKHR01, RF01, BKPA09, CP03, HC04a, KMMB09, KS08a, KS08b, KS08c, LZXW09, MF05, MPPB04, OCB09, RK06, SLG04, VE04, vdHvS03, vSvS06a]. Modeling [Bra06c, LGW07, CP03, PML09, Wan05]. Modelling [Beg04, Pur01]. models [Ber03, FBV06, HC03, KL07, MP09, OKE09, PdVGA08a, PdVGA08b, PdVGA08c]. Modified [CMY07]. Money [Phi01a, Hin03h]. Monitoring [Sch01a]. Moon [KMWD06]. mouse [Bha03]. moving [DEE09]. Much [Nik02c]. Multi [GKH09, KKHR01, VSR07, Tsa08]. Multi-Agent [VSR07]. Multi-dimensional [GKH09]. Multi-party [KKHR01]. multi-server [Tsa08]. Multiagent [Mou05]. multicast [Asl04a, Asl04c, Asl05, CBB05, CMY07, Seb06, ZCW04]. multicasting [HMDC02]. Multidimensional [vS01b]. multinomial [Wan05]. Multiparty [BCC02, PMRZ00b, CSL09]. Multiple [CZG^{+}05, HSH00, KK00, HC03, SC06]. Multiple-Security-Policies [KK00]. Multiplexer [CL05]. Multiplexer-based [CL05]. Multiprocessor [TM00]. multisignature [CCH03, LWL09]. multisignatures [HR08a, HR08b, HR08c]. mutual [SW06, YC09].

Names [Hin00f, McK03a]. narrowed [Sch04e]. NASA [Han00-59]. National [Eva01, Han00-50, Pou02d]. Nearest [LV02]. Necessarily [Han00-52]. Need [Han00-60, Han00-94, And03, LeV05, Sch04a]. Needed [GM00]. needs [RS04]. negotiation [LL09]. Neighbor [LV02]. net [CH06]. NetHost [AMB06]. NetHost-Sensor [AMB06]. NetSec [McK02a]. Netspionage [Kov00b, Kov00c, Kov00a]. Network [Coh00, Han00-60, Han00-61, Pou01c, Sch01a, AJS06, Bra06c, Cae02, CLK09, DR04, GTDVMFV09, HH05, LG07a, LG07b, LG07c, MW06, MK04, MGMSML^{+}09, OG07, SJ07, SYL09, YZ07, ZLCF08a, ZLCF08b, ZLCF08c]. network-anomaly [OG07]. network-based [AJS06]. networkable [OKH07]. Networks [Har03, AJS06, ASS^{+}06, Asl04a, Beg04, Beg08a, Beg08b, Beg08c, DP06, GBBS08a, GBBS08b, GBBS08c, HFC^{+}08, ILW06, JSD05, KVD06, KVD07, OCB09, TE06, TC05, Tse07, WZM09, WH04, vSM04]. neural [AJS06, Beg08a, Beg08b, Beg08c, HFC^{+}08]. Next [Bar01a, Van03, CH06, Sch07b]. NGSCB [Sch03c]. NIDS [Gra01]. Nimbyism [Hin03e]. No [Han00-102]. Non [Sch05j, RFR07a, RFR07b, RFR07c, PRZ04]. Non-infosec [Sch05j]. non-intrusive [RFR07a, RFR07b, RFR07c]. Non-PKI [PRZ04]. nonrepudiable [THY04]. nonsense [Ano03l]. normal [CL05, OL03]. notarization [LG04]. Note [McK02c]. Nothing [Nik02c, Pol05]. Notification [AGAS01, AOG^{+}02]. Nov [Hil02b, McK03b]. Novel [SC02, YH01, DSGP06, DSX06, JHgJ05, OG07, WC05]. November [Ano00h, Ano02-41, Ano02-42, Ano05j, Ano06z, Ano09-28]. novice [FTP08a, FTP08b, FTP08c]. NSF [Han00-64]. nuisance [Hin03g]. nulla [Hin01g]. number [SWA^{+}04]. Numbers [Han00b].

OAuth [GRNR09]. obedience [TvS05]. Object [GM00]. obligation [Pou02b]. oblique [GC09]. oblique-matrix [GC09]. observing [FGG^{+}04]. Oct [Hil02b, McK03b]. Oct-1 [McK03b]. October [Ano00g, Ano01-44, Ano02-39, Ano02-40, Ano03-38, Ano04n, Ano05i, Ano06y, Ano08x, Ano08y, Ano08z, Ano09-27]. Odyssey [Hin01a, Hin01b]. off [YZPL09]. Office [Han00-65]. Officer [Kov01a, Han01a]. Official [Han00-86]. Old [Han00w, Hin00b, Hin00e, Hin00f]. Omnia [Hin01g]. on/off [YZPL09]. one [HLTJ09, LCLL09, Sch04d, Tsa08, YL04]. one-time [HLTJ09, Sch04d]. one-way [Tsa08, YL04]. Online [CDD^{+}02, Han00q, Han00u, Han00t, Han00-79, Han00-96, Han00-97, Sch03l, Che04, GKS^{+}06, LPF06]. ontology [GLK08a, GLK08b, GLK08c]. ontology-based [GLK08a, GLK08b, GLK08c]. Open [Han00-99, HKP02]. Opening [She00]. operating [DSX06, Hof05, OKH07]. operational [HC04a, vS05b]. Operationalizing [CM03]. opportunities [HKP02]. optimistic [HAGTA08a, HAGTA08b, HAGTA08c]. Opts [Han00-64]. Ordered [Han00-35]. Organisational [RMC07]. Organization [MMZ00, Zun00, Hin00a]. organizational [ASS^{+}06, KMMB09, KCC09]. organizational-level [KMMB09]. Organized [Beq01a]. oriented [GM08a, GM08b, GM08c, MLD08, MSLD09]. OSI [SLG04]. OSNP [HLTJ09]. Other [Hin00g, Hin02a, Hin02f]. overlay [MGMSML^{+}09]. Overlooked [Sch06h]. Overview [Phi01b]. Own [Han00a]. Oxley [Beq03, FN05, Sch04f].

P2P [LGW07, WZM09]. PACA [Art04]. packet [BC05, YH07d, YZPL09]. packets [YH07a, YH07b, YH07c]. Pages [Ano00e, Ano00i, Ano00b, Ano00c, Ano00f, Ano00d, Ano00h, Ano00g, Ano01-46, Ano01-45, Ano01-39, Ano01-41, Ano01-42, Ano01-38, Ano01-44, Ano01-40, Ano01-43, Ano02-38, Ano02-37, Ano02-43, Ano02-41, Ano02-42, Ano02-39, Ano02-40, Ano03-33, Ano03-39, Ano03-40, Ano03-34, Ano03-36, Ano03-35, Ano03-38, Ano03-37, Ano04o, Ano04i, Ano04k, Ano04l, Ano04p, Ano04j, Ano04n, Ano04m, Ano05g, Ano05d, Ano05f, Ano05k, Ano05e, Ano05j, Ano05i, Ano05h, Ano06t, Ano06v, Ano06w, Ano06-27, Ano06u, Ano06z, Ano06y, Ano06x, Ano07-35, Ano07-37, Ano07-38, Ano07-39, Ano07-32, Ano07-34, Ano07-40, Ano07-33, Ano07-36, Ano08-27, Ano08-28, Ano08-29, Ano08-30, Ano08u, Ano08v, Ano08w, Ano08x, Ano08y, Ano08z, Ano09w, Ano09y, Ano09x, Ano09-28, Ano09-27, Ano09z]. PAID [GFV05]. pair [PW06]. pairings [DSGP06]. Palestinians [Han00y]. Palladium [Hil02b]. Pandemonium [Sch03d]. Pandora [Sch03c, Sch07b]. Papers [Ano01-47, Ano02-44, Ano09a, Ano09b, WZM^{+}00, Ano02-45, Ano03-42, Ano03-43, Ano06e]. papers-IFIPSEC [Ano06e]. Paradigm [VBD01, Cae02]. parallel [CLK09]. parameter [CC04, Li06]. Paranoia [Han00r]. Parliament [Han00t]. parsing [BCS03]. Part [Han00-37, Hof05, Kov00b, Kov00c, Kov00a, Col00]. partially [CLX05]. party [GLMZ08a, GLMZ08b, GLMZ08c, KKHR01, LHL04, LC07]. Passes [Han01b]. passive [WZM09]. Password [CC01, PZ00, FLZ02, Fur07a, Fur07b, Fur07c, JPL04, Jua04, LH03, Pha06, SLH03, WLT03]. password-based [Pha06]. Passwords [PPBH00, OBB05]. Pastime [Han00-79]. Patch [Han00-65, And05]. patches [Sch06a]. Patching [Sch03d]. path [LS05]. pathogens [LK07]. paths [MGMSML^{+}09]. Pathways [KCC09]. Pattern [Gra01, Li04]. patterns [GS03, HCS06, jLC07]. Payment [HSH00, CSL09]. payments [TS05b]. PC [Han00-43]. PDA [Han00x]. Peer [Art04]. Peer-assisted [Art04]. peg [Sch04i]. PENET [PML09]. Pentagon [Han00-40]. perception [Ste04]. Perceptions [Hin01c, FBP07, WDCJ09]. Performance [HS05, SK03]. perils [Hin02d]. Personal [Hin00f, Pal08a, Pal08b, Pal08c, Sch05b, BSC09, FBP07]. Personalised [TNG04]. Personnel [Dhi01]. perspective [KKK05, RMC07]. pervasive [DMWS09]. PGP [Han00-66]. Philippine [Han00-67]. Phishing [DCF07]. Phones [Han00-38]. physiological [RFR07a, RFR07b, RFR07c]. PIN [MvO09]. PING [Kum06]. piracy [Sch03l]. PKI [Han00-64, PRZ04]. placing [Sch07a]. plan [DF06]. Plans [Han00q]. Plates [Cus01]. Pleads [Han00h, Han00-59]. please [Ano05c]. points [KLL09]. police [McK03a]. Policies [KK00, WS02, HAGTA08a, HAGTA08b, HAGTA08c, KKK05, PF08a, PF08b, PF08c, SC06, Ung04, vSvS04b]. Policy [Dav02b, Hun02, Pou01c, AP04, DF06, GLK08a, GLK08b, GLK08c, HE02, HSKL08a, HSKL08b, HSKL08c, KMMB09, Pou02c, Wia05]. polymorphic [TXL09]. Polynomial [SC02, WL05]. Popular [Han00-57, Han00-76]. Porngate [Han00-68]. Pornography [Han00-53]. portable [HM04]. Poster [Han00-35]. potential [Hin01e]. Potentially [Han00-66]. power [WLH06]. Practical [CJT02, Chi08a, Chi08b, Chi08c, Han01p, Nos00b, AJS06, BLP05, PML09, YC04, Zho02]. practice [RKR09]. Practices [Han00k, CF05, Fur07a, Fur07b, Fur07c]. Predation [FBB06]. Predicting [FGG^{+}04, Sch06f, AMR07, Sch02a]. Prediction [MF01, Ano03l, MF05]. preliminary [MF05]. Prepares [Han00-88]. Presenting [Ano03-41]. preservation [LG04]. preserving [GSN07, OZ07, WH04]. press [Ano02d, Ano02e, Ano02f, MH02a, MH02b]. prevent [MS09]. Preventing [BS09, PRZ03]. prevention [CH04, FBB06, Hof05, Sch04g]. Preview [Hil02b, McK03b]. Primes [PMRZ01a]. Primitives [Gol01]. Principles [Gri02]. Priority [Han01-27]. Privacy [GSN07, Han00s, Han00-69, Han00-75, Han00-93, Han01q, Han01o, Hin01a, Hin01b, Hin01h, Hin03f, AN04, BEP04, GKH09, HKP02, Hin02d, Hin03a, JHgJ05, KK02, LSK09, OZ07, SKM08a, SKM08b, SKM08c]. Privacy-preserving [GSN07, OZ07]. Private [Har03]. privilege [CP03]. Probabilistic [BKPA09, GFV05, YH07d]. probability [MF01]. Probe [Han00-51]. Problem [AGAS01, AOG^{+}02, Han00-42, Beg04, TC05]. Problems [Han00-87, YC04]. Proceedings [All01]. Process [AGAS01, AOG^{+}02, EvS00b, LLL02, KMMB09, Pur04, Rog03, SKBM06]. Processes [Fin00]. Processing [Pou02a, Ayo06, SPP07a, SPP07b, SPP07c]. Product [EvS00b]. Products [Wil00, HD05]. Profession [Hor01, Woo01]. Professionals [KJ02, KJ01, Sch05j]. profiles [CCC09, vdHvS03]. Profiling [WGZY06, Rog03]. program [WGZY06]. programming [GSN07]. promise [McK03a]. Promote [Han01b]. Promoting [Gup00]. proof [Chi08a, Chi08b, Chi08c]. propagation [BS07a, BS07b, BS07c]. property [WGZY06]. Proposal [MGMSML^{+}09, Pou01c, Ede03]. Proposals [Han00-93]. Proprietary [Beq01b]. Protect [Web00]. Protecting [Fra01, GKH09, Kov01b, PZ00, Rat01, Zen01b, ASS^{+}06, Kam09]. Protection [Hin01c, Hin03d, HLC07, KA09, KJKV09]. Protestor [Han00-85]. Protocol [Gra01, KKHR01, Sho00, Asl04b, Asl04c, Asl05, BFG^{+}05, BC05, CL06, DP06, GKS^{+}06, GLMZ08a, GLMZ08b, GLMZ08c, HAGTA08a, HAGTA08b, HAGTA08c, HLTJ09, Lia04, LBE08a, LBE08b, LBE08c, LC07, MK06, Seb06, Tse07, YWL05, ZC05]. Protocols [GM00, PMRZ00c, BB05, CC03, ESZ06, JPL05, WLH06]. Prototype [SE02, KK06]. Provably [WLH06]. Providing [BNSZ09, PF09, LG06, YWB^{+}04]. Provision [Han01r]. provoke [McK03b]. proximity [HMM09]. proxy [CCH03, HC04b, LTH03, LCC04, SF04, THY04]. Ps [And05]. pseudosolutions [JS03]. Public [Bar00a, BCR03, CB01, CJ03, GKL^{+}00, GM00, LCGS03, LHL04, Lin01, PRZ04]. public-key [Lin01]. Publisher [Ano07a]. Publishers [McK02c]. Puts [Ano01-49].

Quagmire [Beq00b]. qualitative [Alb07, HCS06]. quality [KPsH^{+}08a, KPsH^{+}08b, KPsH^{+}08c]. quantification [BKPA09]. quantitative [BCR03, CZG^{+}05, KS06]. quantization [WC05]. Quarter [Ano01-40]. Queen [Hil02b, McK03b]. Query [OBB05]. Query-directed [OBB05].

radiation [HYJ^{+}09]. Radio [Rob06]. RAID [TM09]. RAID-RMS [TM09]. random [SWA^{+}04]. randomized [CLX05]. ranking [KK08a, KK08b, KK08c]. rate [MFDVGT08a, MFDVGT08b, MFDVGT08c]. Rates [PPBH00]. rational [ZHH09]. raw [CO09]. RBAC [Ber03, LZXW09, MPPB04]. Re [Boo06, Han00-61]. Re-engineering [Boo06]. Re-evaluates [Han00-61]. reaction [HD05]. Real [FvS05, JSD05, LCK^{+}06, AJS06, SG09, SAF09, SYL09]. Real-time [FvS05, JSD05, LCK^{+}06, AJS06, SG09, SAF09, SYL09]. Reality [Hin01c]. Really [Han00-48, Sch04h]. Recall [PPBH00]. receipt [Sch07a]. Recognition [HYJ^{+}09, Li04]. recognizing [HFC^{+}08]. Recommended [Han00-65]. reconfigurable [HSKL08a, HSKL08b, HSKL08c]. Recovery [BM00, Gup00, GM00, KMZ00, MMZ00, MW00, MZ00, Sho00, SVW00, Wil00, WZ00, WZM^{+}00, Zun00, LG06, RM05, Sch05g]. Recruiting [Han00o]. Red [Ano01-49]. reduced [Thi03]. Reducing [MvO09]. Refereed [Ano01-47, Ano02-44, Ano02-45, Ano03-42, Ano03-43]. referees [Ano01-33, Ano01-37, Ano01-30, Ano01-31, Ano01-32, Ano01-34, Ano01-35, Ano01-36, Ano02-32, Ano02-33, Ano02-34, Ano02-35, Ano02-36, Ano03z, Ano03-27, Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano03-32]. Reflecting [BG06]. Reformed [Han00-48]. regression [Wan05]. Regulation [Han00-54, Han00-75]. Regulations [TMM01]. Regulatory [Han00n, GKH09]. rekeying [NCS06]. Related [Dhi01]. relation [RS03]. Relay [HMM09]. Releases [Han00f, Han00-58]. Reliable [OCB09, HM04, Lea04, Sam09]. reliably [Li04]. relic [McK02b]. rely [FL05]. Remainder [CC02]. Remote [CJT02, HS00, DSGP06, FCZ05, MW06, SW06, WC03, YW04, YC09, YRY05]. repairing [dSSW04]. replication [FBB06]. Report [Ano03-41, Eva01, Han00-35, Mou05]. reporting [GKH09, Wia05]. Representing [Sch06g]. reputation [MP09]. requests [LCLL09]. requirement [Zuc04]. Requirements [GvS01, BAS09, GvS08a, GvS08b, GvS08c, GKH09, Gri02, RM05]. Reseacrh [Pip03]. research [Bis02]. Reserve [Han00q]. resource [Tse07]. resource-constrained [Tse07]. resources [BS09]. respect [HKP02]. Response [Han00-67, Han00-78, MPD06, Nov07, RS03, Sch04a, Sch05g, ZHH09]. restitution [LPF06]. Restraining [dSSW04]. Restricted [Han00-97]. Restricting [PK07]. result [Sch07d]. Results [AGAS01, AOG^{+}02]. retorsum [Hin01g]. Retraining [jLC07]. Retrieval [Zun00, KS08a, KS08b, KS08c]. return [Ihe04]. Revenger [Ano02n]. Reverse [GRNR09, SKBM06]. review [CH06, Ecc01d, Her09a]. Reviews [Ecc01e, Ecc01f]. Revisited [Hin01b, HMM09, OHB08]. Revolution [Han00-50]. RFID [Rob06, Ayo06]. rhythms [KPsH^{+}08a, KPsH^{+}08b, KPsH^{+}08c]. Rico [SF04]. Rights [TMM01, Bra07d, KK02, Sch06c]. Ring [Han00-53, KKHR01]. RIP [Han00-86, SSF^{+}09]. Rise [Han00m]. Risk [Ber07, CCC09, CM03, Fin00, GvS01, Han00q, Han00-37, Han01p, LPF06, Nos00b, RF01, GvS05, Hin03d, KS05, Sch05c, Sch06h, Ste04]. Risks [Hin00b, SE02, Wri01, Mat03, vSM04]. RMS [TM09]. Roadmap [KN03]. robots [LCLL09]. Robust [FCZ05, KA09, SSF^{+}09]. rogue [KLL09]. ROI [Pur04, Lea03a]. role [MPPB04, Rog03, TE06]. Romance [Han00-79]. Romancing [Beq00b]. Room [Han00j]. rootkits [BIC08a, BIC08b, BIC08c]. rotations [SK03]. round [Sch04i, YH07d]. round-trip [YH07d]. RSA [CLX05, HR08a, HR08b, HR08c, PMRZ00a, PMRZ01a]. RSA-based [CLX05]. RT [AJS06]. RT-UNNID [AJS06]. rule [HC03]. rule-based [HC03]. Run [BS07a, BS07b, BS07c]. Run-time [BS07a, BS07b, BS07c]. Russian [Han00-81].

S [LG09]. S/MIME [LG09]. SAD [CC04]. Safe [Han01r, HM04]. Safeguards [Beq03, Dhi01]. safely [Her09b]. Safety [Mou05]. Safeway [Han00-70]. same [Hin01d]. Sarbanes [Beq03, FN05, Sch04f]. SASEMAS [Mou05]. SASEMAS’05 [Mou05]. say [HE02]. Says [Han00-75]. SCADA [ILW06]. Scalable [NCS06, Asl04c, Asl05]. scale [PF08a, PF08b, PF08c, SYL09]. Scam [Han00-77, Ede03]. Scamming [Hin00g]. scams [Hin02f]. scanning [CS09]. scenarios [MP09]. schema [Kam09]. Scheme [CC01, SC02, YH01, Asl04a, CLX05, CBB05, CL03, Che04, Chi08a, Chi08b, Chi08c, DSGP06, FLZ02, FCZ05, GYLC09, HLC07, KMWD06, LH03, SC06, SLH03, Tsa08, THY04, WLT03, WC03, WH04, WC05, YW04, YWB^{+}04, YWL05, YC09, YRY05, Zho02, ZL03]. Schemes [Wil00, CCH03, HW03, HC04b, JPL04, Pha06, SP06, WL05]. Schemes/Products [Wil00]. Schneier [Hei03]. Scope [Ano09q, Ano09r, Ano09s]. Screen [Hin02a]. scripting [KJKV09]. SDriver [MS09]. Search [AN04, Gra01, Hin01h, Sch05c, Cae02]. SEAS [BFG^{+}05]. seat [Sch06e]. SEC [Ano09p, BG06]. sectors [Kja06]. Secur [McK04]. Secure [CSL09, HLTJ09, LL02, MK06, McG02, Nab05, SKM08a, SKM08b, SKM08c, VFMP05, YWL05, vKH00, BFG^{+}05, BNSZ09, CJC04, CMY07, GLK08a, GLK08b, GLK08c, GRJ07, Gri02, Hof04, HMDC02, KMWD06, LZXW09, Lia04, LH03, LCC04, LBE08a, LBE08b, LBE08c, Mac05, NCS06, OZ07, WLH06, ZCW04, vSM04]. SecurID [BLP05]. Securing [Car01, GKL^{+}00, Kum09, DFSC09, HC04a]. Security [Ano03b, Ano03c, Ano03-44, Ano04q, Ano04r, Ano04s, Ano04t, Ano05l, Ano05o, Ano05p, Ano05m, Ano05n, Ano06-28, Ano06-29, Ano06-30, Ano06-31, Ano06-32, Ano06-33, Ano06-34, Ano07-42, Ano07-41, Ayo06, Bar00b, Bar01b, BvS00, Ber07, Bla02, BCS03, CDD^{+}02, Dav02a, Dav03, Dhi01, EvS00a, EvS00b, Eva01, FTP08a, FTP08b, FTP08c, GvS01, Han00f, Han00k, Han00i, Han00-42, Han00-58, Han00-60, Han00-61, Han00-71, Han00-72, Han00-73, Han00-74, Han00-84, Han00-89, Han00-90, Han00-96, Han00-102, Han00-98, Han01b, Han01a, Han01s, Han01t, Han01u, Han01v, Han01w, Han01-27, Han02, Hin02e, HS00, Hor01, HM04, ILW06, JS02, JEL03, KK00, LCGS03, Lea03a, LLL02, MP09, MS00, McK02b, Mou05, Nos00a, PW06, Pou01c, Pou02c, Pou02d, Pou03a, Sch01a, Sch01c, SPLS01, Sch02h, Sch02i, Sch02j, Sch02k, Sch02l]. Security [Sch02m, Sch02n, Sch03l, Sch03g, Sch03h, Sch03i, Sch03j, Sch03k, Sch04i, Sch04j, Sch04b, Sch04c, Sch04k, Sch05k, Sch05d, Sch05e, Sch05f, SLH03, TM00, TE01, VSR07, Woo01, vKH00, vS00, vS01a, vSvS01, vS01b, vS05a, vS05b, vS06, vSvS06a, Alb07, AH09, AMR07, And03, AC08a, AC08b, AC08c, And05, Ano02d, Ano02e, Ano02f, AP04, Asl04c, Asl05, BTY^{+}07, BKPA09, BB05, BA04, Boo06, BFC09, Bra06c, Bra06b, Bra07a, Bra07b, Bra07c, BP03, Cae02, Chi08a, Chi08b, Chi08c, CITN07, DEE09, DCF07, DF06, Dou03, DMWS09, DKS07, DSX06, Fur05, FJK06, FBP07, Fur07d, Fur07e, GBBS08a, GBBS08b, GBBS08c, GvS08a, GvS08b, GvS08c, GS03, HCS06, HKP02, Hei03, HS05, Hin02b, Hin03b, HE02, KS05, KKK05, KL07, Kin04, Kja06, KMMB09, KVD06, KVD07]. security [KCC09, KS08a, KS08b, KS08c, KK06, KK08a, KK08b, KK08c, KK07, LeV05, Lea03b, Lea04, MvO09, MH02a, MH02b, MC03a, MC03b, MLD08, MSLD09, NTH04, OKH07, PF09, PE06, Pip03, PK07, PvS04, Pou02b, Pou03b, PML09, Pur04, RS03, RM08a, RM08b, RM08c, RSS02, RKR09, RMC07, RR06, Sch02d, Sch02g, Sch03a, Sch03b, Sch04e, Sch04f, Sch04h, Sch05g, Sch05h, Sch05c, Sch06b, Sch06g, Sch06h, SS07, SF04, SSMJ05, SLG04, TvS05, Trè03, TS05b, TS05a, Van03, VE03, VvS04, WDCJ09, Wia05, YWB^{+}04, Zuc04, Zuc07, vdHvS03, vSvS04a, vSvS05, vSvS06b]. security-by-contract [DMWS09]. see [Ano05c]. seizure [Wol03g]. Selection [BvS00, LWT^{+}09]. Self [Han00-54, Han00-75, RKR09]. Self-efficacy [RKR09]. Self-Regulation [Han00-54, Han00-75]. Sell [Han00-97]. Senate [Han00-94]. sender [Her09a]. sense [Ano03l]. sensibly [Hei03]. sensitivity [GSK09]. Sensor [AMB06]. separation [MPPB04]. September [Ano01-43, Ano03-37, Ano04m, Ano05h, Ano06x, Ano07-36, Ano09z]. sequences [FGG^{+}04]. Serious [Han00k, Sch04h]. Server [Fra01, LHL04, Tsa08]. servers [MFDVGT08a, MFDVGT08b, MFDVGT08c, Sch03f]. Service [CBR06, CE04, Han00-63, CLC04, ZC05]. Services [McK02b, DMWS09, GLK08a, GLK08b, GLK08c, GVL09, KHKK06, MW06, McK03b, MC03b]. session [CC04, OHB08, WC07, YWL05]. session-aware [OHB08]. set [McK03b]. Sets [Han00l]. Setting [Han00-90, Wol03f]. Shieh [McK04, CZ03]. Should [Han00-48, Kov00b, KJ02, KJ01]. Shut [Han00-70]. side [KLL09, KJKV09, MSZ05, Pou03a, Sch07c]. sign [GRNR09]. sign-on [GRNR09]. signal [SBS09]. Signature [Han00-64, CLX05, ES09, HAGTA08a, HAGTA08b, HAGTA08c, HC04b, LTH03, LCC04, SC06, THY04, ZCW04]. signature-based [LTH03]. signatures [LG04, LG09, MS09, TXL09]. signcryption [KMWD06]. signers [THY04]. significant [LSK09]. signing [SC06]. signs [Li04]. SIKA [MK06]. similarities [LK07]. similarity [SPP07a, SPP07b, SPP07c]. Simple [Coh00, Hin00c, LC07, Pur01, And05, GT06, GLMZ08a, GLMZ08b, GLMZ08c, YRY05]. Simulation [AGAS01, AOG^{+}02]. single [GRNR09]. Sinkhole [TC05]. sins [vSvS04a]. SIP [GLK08a, GLK08b, GLK08c, GVL09, GM08a, GM08b, GM08c]. SIP-based [GLK08a, GLK08b, GLK08c]. SIP-oriented [GM08a, GM08b, GM08c]. Sister [Han00l]. Site [Han00e, Han00-44, Han00-49, Han00c, Han00-76, KJKV09]. Slandering [Hin00g]. Smaller [Han01z]. Smart [Ano05q, CJT02, FCZ05, HMM09, Jua04, Pha06, SLH03, WC03, YW04]. Smarter [Han01x]. smartphones [BSC09]. SMS [LBE08a, LBE08b, LBE08c]. SMSSec [LBE08a, LBE08b, LBE08c]. SMTP [GT06]. Smurfing [Hin00g]. Snooping [Han00-86]. society [Bha03]. Socio [TE01]. Socio-ethical [TE01]. Software [Beq01c, CG00, Han00-77, LLL02, McG02, AMR07, HCS06, KA09, PRZ03]. Solution [CJT02, AJS06, GRNR09]. Solutions [Han00-61, JS03, OKH07, YC04]. solving [Beg04]. Some [Han00-44, TC05, Wol03e]. sophistication [MF05]. Sorry [Han00v, Sch02e]. Source [Han00-99, HKP02, CBB05, RS03]. sources [Sch06h]. Space [Beq01b]. Spam [Hin02f, Hin03g, CH06, CH04, Ede03, GT06, Sch07a, Han01e]. Spammers [Han01c]. Spamming [Hin00g, WC07]. spatial [TE06]. Special [Sch06h, WZM^{+}00]. specific [MS09]. specifications [HSKL08a, HSKL08b, HSKL08c]. specificity [GSK09]. Specified [PMRZ01b]. Specifying [SBS09]. spectantia [Hin01g]. speed [JSD05]. Speedy [Han00-78]. SPIT [GM08a, GM08b, GM08c]. spoofed [LS05]. Spoofing [Han00-76, Hin00g]. spread [LGW07]. spreading [PF08a, PF08b, PF08c]. spring [Hin02e]. Spying [Han00-79]. spyware [Sch03c]. SQL [MS09]. square [Sch04i]. Squatting [Hin00g]. SSL [ESZ06, Har03, MSZ05, OHB08, Sho00]. SSL/TLS [ESZ06, OHB08, Sho00]. st [Kov01b, Nos00a]. staff [Beq03]. stand [Sch02f]. Standard [Han00-98, MS00]. Standardising [VEL08]. Standards [LLL02, HE02, Kin04]. State [Sch02e, MPD06]. state-of-the-art [MPD06]. States [BEP04]. statistical [GSK09]. stay [Sch03e]. steganalysis [GSK09]. Steganographic [LL02]. Steganography [HCBLETRG06, CO09]. Step [Bar01a]. stepping [YH07a, YH07b, YH07c]. stepping-stone [YH07a, YH07b, YH07c]. sthreats [McK03b]. Still [Han00-40]. stone [YH07a, YH07b, YH07c]. Stones [Han00y]. Stored [Zun00]. strategic [DF06]. strategies [AC08a, AC08b, AC08c]. Strategy [Pou02d, And05, LL09]. Stream [BC05, Gol01, Han00w, DS09]. streaming [CBB05]. streams [CO09, ZCW04]. Street [Han00y]. Streets [Han00-83]. stripped [TM09]. Structure [Coh00, SG09]. Struggles [Sch07c]. studies [LFHT07]. study [Alb07, BTY^{+}07, Beg08a, Beg08b, Beg08c, BCR03, Dou03, IKE06, KL07, Kin04, KK08a, KK08b, KK08c, LCLL09, LSK09, RM08a, RM08b, RM08c, YZPL09]. stumps [JST^{+}07]. subgroup [Asl04c]. subgroup-key [Asl04c]. Subject [Ano00j, Ano01-48, Ano02-46, Ano03c]. subjects [SV07]. Subscriber [CFRR02]. subscription [MW06]. subscription-based [MW06]. substitution [MP07]. Success [Han00-78]. suitable [ÁP03]. Summit [Han00-98]. supervised [LG07a, LG07b, LG07c]. Supervision [FDIR00]. support [HC04a, Ung04]. Supporting [Han00-99, HSH00]. Supposedly [Han01b]. Supreme [Han00-95]. Surf [Hin00f, Her09b]. Surfing [Hin00g]. surprising [MSZ05]. Survey [AGAS01, AOG^{+}02, FDIR00, CF05, ES09, FJK06, LOP04, QW05, RS04, VFMP05]. surveys [Hin02e]. surviving [Hil02b]. Suspect [Han00-80]. Suspected [Han00-81]. SVision [OG07]. SVM [HYJ^{+}09, KC05]. SVM-based [KC05]. Swamping [Hin00g]. Symmetric [MPPB04, CC02, YZ07]. Sympathy [Han00-44]. Syndicated [Beq02a]. System [VSR07, YH01, BP03, CJC04, DGY^{+}05, DSX06, FGG^{+}04, FvS05, GSK09, GFV05, Gri02, Hof05, LWT^{+}09, LG06, OKH07, dSSW04, SYL09]. systematic [MLD08]. Systems [All01, Bar01a, BE01, CDD^{+}02, Fin00, JS02, Mou05, TM00, WS02, AMR07, Ano03l, ASS^{+}06, BCV01, CAT05, DF06, GTDVMFV09, GRNR09, Hof04, KKK05, KK02, KL07, Lea04, LGW07, Mac05, MF05, MP09, MK04, Nab05, RSS02, Sam09, Sch06h, SS07, SK03, TKKK05, TM09, Trè03, VE02, VFMP05]. systems/ciphers [SK03]. sython [GSN07]. table [Tsa08]. Tactics [KJ02, KJ01]. tailers [Han01d]. Take [SJ07]. takes [MPPB04]. Taking [Sch02f]. tampering [DR04]. Tang [YRY05]. tangled [Pou03c]. Target [Han00-38, Han00-57, DEE09]. Targeted [Han00-43, Han00-52]. Targets [KB00]. tasks [PK07]. Taxes [Han00-85]. taxonomy [ÁP03, HH05, Kja06, SP06, VE03]. TBSE [Lea04]. TC11 [Ano09q, Ano09r]. TCII [Ano09s]. TCM [LG07a, LG07b, LG07c]. TCM-KNN [LG07a, LG07b, LG07c]. TCP [YH07a, YH07b, YH07c, YH07d]. TCP/IP [YH07a, YH07b, YH07c]. te [Hin01g]. Teaching [SS07]. Team [Han00-67]. teams [Sch04a]. tech [Beq02b]. Technical [Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-28, Ano02u, Ano02x, Ano03x, Ano09t, Ano09u, Ano09v, Ano01-29, Ano02v, Ano02w, Ano02y, Ano02z, Ano02-27, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano03w, Ano03y]. technicians [BTY^{+}07]. technique [GC09, OG07, WC05]. Techniques [GTDVMFV09, PPBH00, SPP07a, SPP07b, SPP07c]. technologies [FBV06, VE03]. Technology [LeV05, SE02, WS02, WZM^{+}00, Kum09, Ja02]. Teen [Han00h]. Tele [HM04]. Tele-Lab [HM04]. Telecom [Pol05]. Telecommunications [Col00]. telephones [CF05]. Telephony [CFRR02]. Tells [Han00-92]. temporal [TE06]. Ten [Han00-96, Han01a]. term [LG04]. Terrorism [Pou02a, Beq02a, Beq02b, Hil02b, Hin01e, Hin03c]. Terrorist [Woo01]. Terrorists [KJ02, KJ01]. testimony [Wol03e]. tests [JPL04]. Texas [Han00-53]. text [SPP07a, SPP07b, SPP07c]. Theft [Han00-52]. Their [Han00y, Han00-92, BLP05]. Them [Han00-48]. Theorem [CC02]. theoretic [ZHH09]. theorizing [DM01]. There [Han00-36, And05]. Thing [Han01o]. things [Fur07e]. thinking [Hei03]. Thinks [Han00-34]. Third [vS00]. those [Hin01d]. thoughts [Wol03e]. Threat [Beq00a, Kov00b, Kov00c, MF01, CZG^{+}05, MF05, TKKK05]. threats [BKPA09, MvO09, MP09]. three [GLMZ08a, GLMZ08b, GLMZ08c, LHL04, LC07]. three-party [GLMZ08a, GLMZ08b, GLMZ08c, LHL04, LC07]. threshold [SC06, THY04]. Ticket [PMRZ00c]. Tightening [CH06]. Time [Han00-34, Hin03h, Kov01b, AJS06, BS07a, BS07b, BS07c, FvS05, HLTJ09, JSD05, LCK^{+}06, SG09, Sch04d, Sch05b, SAF09, SYL09, YH07d]. timely [RS03]. Times [Hin01i, PMRZ01b]. Timestamp [CC01, FLZ02, SLH03, WLT03]. Timestamp-Based [CC01, FLZ02, SLH03, WLT03]. Timestamping [PMRZ00c, PMRZ01b]. Timing [FVA05, YZPL09]. TLS [ESZ06, OHB08, Sho00]. Today [CDD^{+}02]. token [HMM09]. tokens [WDCJ09]. tolerant [Seb06, TM09]. Too [Han00-34]. Tool [Han00-58, Han00-83, MF01, Pur01, Ber07, PML09]. tools [CH06, CH04]. Top [Han01a, Hof05]. traceable [SC06]. Traceback [ASP03]. tradeoffs [PK07]. traffic [Li04, Li06, PF08a, PF08b, PF08c, YZ07]. Tragedy [Ano02n]. Trails [All01]. Training [Hor01, CITN07, HM04, Sch04i]. Traits [Wil00]. transaction [Che04]. Transactions [HSH00]. transition [WGZY06]. Transmission [PZ00]. treat [Hin01d]. Treaty [Han00-47]. tree [GYLC09, HLL05]. tree-based [HLL05]. trees [BP03]. trend [BvS03, Li06]. Trends [Bis02, Phi01b, Ano03m, Ber03, QW05]. Trinity [Han00-83]. trip [YH07d]. triumph [Hin01d]. Trojan [BNSZ09]. true [PF09]. Truly [Han00-84]. Trust [CB01, DeM02, Pur01, LPF06, LL09, MP09, OCB09, TS05b]. Trusted [Cae02, Dhi01]. trustworthy [Cae02, SK03]. tunnel [IKE06]. turn [Sch05i]. TV [Han00-49]. Two [Asl05, Hin01d, HW03, Pha06, WL05]. Two-level [Asl05].

U.S. [Ano01-49, Sch04h, Sch05b]. Ubiquity [Han00-57, Han00-103]. UK [Han00-70, Han00-85, Han00-86, Hil02b, McK03a, McK03b]. unauthorized [PRZ03]. uncertain [Hei03]. Uncovering [IKE06]. Understanding [AN03, Dhi01, FKSF01, GS03, LG09, FJK06, Sch02a]. unfortunate [Pou03a]. unintended [GRJ07]. Union [Pou01c, BEP04]. Unique [PMRZ00a, PMRZ01a]. Unit [Han00q]. United [BEP04]. University [RSS02]. UNIX [JST^{+}07, KC05, Han00-62]. Unix/Linux [Han00-62]. unknown [Zen01a]. UNNID [AJS06]. Unofficial [Sch06a]. unstructured [WZM09]. unsupervised [AJS06]. Update [Ano05p, Ano06s, Ano06-29, Ano06-30, Ano06-31, Ano06-32, Ano06-33, Ano07-30, Ano07-42, Han00-84, Ano05c, Ano05o, Ano05m, Ano05n, Ano06c, Ano06-28, Ano06-34, Ano07-31, Pou02c]. updating [LH03]. Usability [SPLS01, WDCJ09, Woo04]. usable [Fur07e]. usage [BCS03, PE09]. Use [Han00-44, LV02, Fur05, Woo04]. used [CH04]. User [FDIR00, PMRZ00a, PMRZ01b, WDCJ09, Alb07, Chi08a, Chi08b, Chi08c, CF07, DSGP06, DCF07, GS03, HW03, Lea03b, MF05, MK06, MP07, OL03, OHB08, PK07, RKR09, RMC07, SSMJ05, WC03, WH04, YW04, YWB^{+}04, YRY05, Zho02]. User-Specified [PMRZ01b]. Users [Han00-92, AH09, BA04, CF05, Fur05, FJK06, FBP07, FTP08a, FTP08b, FTP08c, Her09b, LPF06]. Uses [Kov00a]. Using [FKSF01, KKHR01, PMRZ01b, SVW00, TXL09, WC07, YH01, ASP03, AJS06, Asl04c, CP03, DSGP06, DR04, DS09, DSX06, FJK06, GSN07, GKH09, HLTJ09, JST^{+}07, Jua04, KLL09, KC05, KL07, Kum09, LWT^{+}09, Pha06, SBS09, SPP07a, SPP07b, SPP07c, SLH03, YL04]. Utilising [BvS03, RFR07a, RFR07b, RFR07c]. Utilizing [GVL09].

v3 [Han00-83]. V5.0 [Han00-66]. Value [DKS07]. Value-focused [DKS07]. Various [EvS00a]. vector [WC05]. vectors [Rob07]. veiled [GKH09]. verification [Tsa08, YC04]. verification-its [YC04]. via [LCK^{+}06, Zun00]. video [CO09, CITN07]. view [Alb07]. Views [Ano03-44, Ano06-29, Ano06-30, Ano06-31, Ano06-32, Ano06-33, Ano06-34, Ano07-42, Ano07-41, Han00-71, Han00-72, Han00-73, Han00-74, Han01s, Han01t, Han01u, Han01v, Han01w, Sch01c, Sch02h, Sch02i, Sch02k, Sch02l, Sch02m, Sch02n, Sch03h, Sch03j, Ano04q, Ano04r, Ano04s, Ano04t, Ano05l, Ano05o, Ano05p, Ano05m, Ano05n, Ano06-28, Sch02j, Sch03l, Sch03g, Sch03i, Sch03k, Sch04j, Sch04b, Sch04c, Sch04k, Sch05d, Sch05e, Sch05f]. Violation [Dhi01]. Virii [Han01m]. Virtual [Han00y, Har03, LG06]. Virus [Ano03m, AGAS01, AOG^{+}02, Han00x, Han00w, Han00-57, Fis03, Zen01a]. Viruses [Zen01b, HD05, PdVGA08a, PdVGA08b, PdVGA08c, Sch04l, hZxZtZ06]. Visa [Han00-96]. Vista [Sch07e]. visual [OG07]. visualization [CLK09]. voice [PF09]. VoIP [Bra07a, Bra07b, Bra07c, DFSC09, GLK08a, GLK08b, GLK08c]. Volume [Ano00a, Ano00j, Ano01a, Ano01-48, Ano02b, Ano02-46, Ano03b, Ano03c]. Voodoo [Han01y]. voting [BFG^{+}05, CL06, CJC04, Gri02, Lia04, SP06]. VPN [IKE06]. vs [Gra01, vS05b]. VTG [SJ07]. vu [Hin02c]. Vulnerabilities [FKSF01, VE02, AMR07, Bis02, JS03, KFS^{+}03, KCC09, RS03, SJ07]. Vulnerability [VE04, Sch06d, VEL08, SJ07].

wake [Sch05b]. Wakes [Han00-94]. Walls [McG02]. Wants [Han00-64, Han00-54, Han00-69]. war [Mat03]. Warfare [Han01n, KJ02, Ede03, KJ01]. Was [Hin00c, Hin02c]. watermarking [HLC07, KA09, WC05]. Wave [vS00, vS06, McK02b]. Wavelet [CBR06]. Way [Han00w, Han01m, And05, Sch05i, Tsa08, YL04]. Weakest [Hin01j]. Web [Han00-69, BS09, CC04, Her09b, LCLL09, ÁP03, CE04, Fra01, Han00-44, Hin00g, McK03b, Pou03c, Sch03f]. website [Fur07a, Fur07b, Fur07c, Han00-70]. Well [Han01y]. went [Pol05]. Westminster [Hil02b, McK03b]. Which [Sch05i]. White [Beq02b, Han00-99, Han00-98]. Who [Kov00a, RSS02, Han00-100]. Wide [Hin00g]. Win32.Perrun [Nik02c]. window [Bra06b]. Windows [OKH07, Sch07e]. Windows-based [OKH07]. winds [Sch06b]. wired [vSM04]. Wireless [Han00-101, Han00-102, Han00-103, Han01z, HLTJ09, LFHT07, TE06, Tse07, Van03, vSM04]. Wiretap [Han00-91]. within [DM01]. without [CJ03, LHL04, Lin01, Tsa08]. Word [PPBH00]. Work [Han00-75]. Workflow [BE01]. workplace [Dav02b]. workshop [Fur07d, Mou05]. World [Bar00b, Bar01b, Hin00g, Hei03, Sch07e]. Worm [Han00-80, Ano03m, Nik02b, PF08a, PF08b, PF08c]. Worm.Lentin [Nik02a]. Worms [Sch04l, CS09, LK07, LGW07, QW05, TXL09, WZM09]. Worsening [AGAS01, AOG^{+}02]. Worst [Han00a]. wrapped [HLC07]. wrapper [LWT^{+}09]. wrapper-based [LWT^{+}09]. writers [Bra06a]. WTLS [ZC05].

X86 [TM00]. XML [AP04, BCS03, HC04a, JCKL04]. XPath [JCKL04].

Y2K [Han01-27]. Yaha [Nik02a]. Yahoo [Han00-104]. Yang [McK04, CZ03]. Year [Han01-27]. Yields [Han00-78].

Zealand [JS02]. Zone [Kov00c, Kov00a].

References

Anderson:2008:EISa

[AC08a] Evan E. Anderson and Joobin Choobineh. Enterprise information security strategies. Computers & Security, 27(1–2):22– 29, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000060.

Anderson:2008:EISb

[AC08b] Evan E. Anderson and Joobin Choobineh. Enterprise information security strategies. Computers & Security, 27(1–2):22– 29, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000060.

Anderson:2008:EISc

[AC08c] Evan E. Anderson and Joobin Choobineh. Enterprise information security strategies. Computers & Security, 27(1–2):22– 29, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000060.

Aron:2001:BNP

[AGAS01] Joan L. Aron, Ronald A. Gove, Shiva Azadegan, and M. Christina Schneider. The benefits of a notification process in addressing the worsening computer virus problem: Results of a survey and a simulation model. Computers & Security, 20(8):693–714, December 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801008124.

Albrechtsen:2009:ISD

[AH09] Eirik Albrechtsen and Jan Hovden. The information security digital divide between information security managers and users. Computers & Security, 28(6):476–490, September 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809000029.

Amini:2006:RUP

[AJS06] Morteza Amini, Rasool Jalili, and Hamid Reza Shahriari. RT-UNNID: a practical solution to real-time network-based intrusion detection using unsupervised neural networks. Computers & Security, 25(6):459–468, September 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806000782.

Albrechtsen:2007:QSU

[Alb07] Eirik Albrechtsen. A qualitative study of users’ view on information security. Computers & Security, 26(4):276–289, June 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404806002033.

Allinson:2001:ISA

[All01] Caroline Allinson. Information systems audit trails in legal proceedings as evidence. Computers & Security, 20(5):409– 421, July 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404801005132.

Abimbola:2006:NSI

[AMB06] A. A. Abimbola, J. M. Munoz, and W. J. Buchanan. NetHost-Sensor: Investigating the capture of end-to-end encrypted intrusive data. Computers & Security, 25(6):445–451, September 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404806000605.

Alhazmi:2007:MAP

[AMR07] O. H. Alhazmi, Y. K. Malaiya, and I. Ray. Measuring, analyzing and predicting security vulnerabilities in software systems. Computers & Security, 26(3):219–228, May 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001520.

Aljifri:2003:ILA

[AN03] Hassan Aljifri and Diego Sánchez Navarro. International legal aspects of cryptography: Understanding cryptography.

Computers & Security, 22(3):196–203, April 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803003055.

Aljifri:2004:SEP

[AN04] Hassan Aljifri and Diego Sánchez Navarro. Search engines and privacy. Computers & Security, 23(5):379–388, July 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804000240.

Anderson:2003:WWN

[And03] James M. Anderson. Why we need a new definition of information security. Computers & Security, 22(4):308–313, May 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404803004073.

Andrew:2005:FPP

[And05] C. Andrew. The five ps of patch management: Is there a simple way for businesses to develop and deploy an advanced security patch management strategy? Computers & Security, 24(5):362–363, August 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805000994.

Anonymous:2000:AIV

[Ano00a] Anonymous. Author index for volume 18. Computers & Security, 19(2):175, February 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480087827X.

Anonymous:2000:PF

[Ano00b] Anonymous. Pages 106–194 (1 February 2000). Computers & Security, 19(2):??, February 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2000:PJa

[Ano00c] Anonymous. Pages 2–104 (1 January 2000). Computers & Security, 19(1):??, January 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2000:PM

[Ano00d] Anonymous. Pages 200–288 (1 March 2000). Computers & Security, 19(3):??, March 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2000:PA

[Ano00e] Anonymous. Pages 294–374 (1 April 2000). Computers & Security, 19(4):??, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2000:PJb

[Ano00f] Anonymous. Pages 380–469 (1 July 2000). Computers & Security, 19(5):??, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2000:PO

[Ano00g] Anonymous. Pages 474–563 (1 October 2000). Computers & Security, 19(6):??, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2000:PN

[Ano00h] Anonymous. Pages 568–656 (1 November 2000). Computers & Security, 19(7):??, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2000:PD

[Ano00i] Anonymous. Pages 662–746 (1 December 2000). Computers & Security, 19(8):??, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2000:SIV

[Ano00j] Anonymous. Subject index for volume 18. Computers & Security, 19(2):176, February 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800878281.

Anonymous:2001:AIV

[Ano01a] Anonymous. Author index for volume 19. Computers & Security, 20(3):246, May 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801003108.

Anonymous:2001:Ca

[Ano01b] Anonymous. Calendar. Computers & Security, 20(1):75–76, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404801010239.

Anonymous:2001:Cb

[Ano01c] Anonymous. Calendar. Computers & Security, 20(2):161–162, April 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404801002073.

Anonymous:2001:Cc

[Ano01d] Anonymous. Calendar. Computers & Security, 20(3):244–245, May 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404801003091.

Anonymous:2001:Cd

[Ano01e] Anonymous. Calendar. Computers & Security, 20(4):327–328, July 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404801004096.

Anonymous:2001:Ce

[Ano01f] Anonymous. Calendar. Computers & Security, 20(5):405–406, July 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404801005119.

Anonymous:2001:Cf

[Ano01g]  Anonymous.  Calendar.  Computers & Security, 20

(6):510–511, September 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801006101.

Anonymous:2001:Cg

[Ano01h] Anonymous. Calendar. Computers & Security, 20(7):589, October 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404801007088.

Anonymous:2001:Ch

[Ano01i]  Anonymous.  Calendar.  Computers & Security, 20

(8):689–690, December 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801008100.

Anonymous:2001:CFC

[Ano01j] Anonymous. Calendar of forthcoming conferences and events. Computers & Security, 21(1):46, First Quarter 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802001062.

Anonymous:2001:CEa

[Ano01k] Anonymous. COMPSEC 2001 Exhibitors 2001. Computers & Security, 20(6):512–522, September 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801006113.

Anonymous:2001:E

[Ano01l] Anonymous. From the Editor. Computers & Security, 20 (4):276, July 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404801004011.

Anonymous:2001:GAa

[Ano01m] Anonymous. Guide for authors. Computers & Security, 20(1):94, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801010276.

Anonymous:2001:GAb

[Ano01n] Anonymous. Guide for authors. Computers & Security, 20(2):184, April 1, 2001. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801002127.

Anonymous:2001:GAc

[Ano01o] Anonymous. Guide for authors. Computers & Security, 20(3): 274, May 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404801003169.

Anonymous:2001:GAd

[Ano01p] Anonymous. Guide for authors. Computers & Security, 20(5): 446, July 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404801005168.

Anonymous:2001:GAe

[Ano01q] Anonymous. Guide for authors. Computers & Security, 20(6):542, September 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801006162.

Anonymous:2001:GAf

[Ano01r] Anonymous. Guide for authors. Computers & Security, 20(7):638, October 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801007143.

Anonymous:2001:GAg

[Ano01s] Anonymous. Guide for authors. Computers & Security, 20(8):742, December 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801008161.

Anonymous:2001:GAh

[Ano01t] Anonymous. Guide for authors. Computers & Security, 21(1):61, First Quarter 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480200113X.

Anonymous:2001:ITCa

[Ano01u] Anonymous. IFIP technical committee 11. Computers & Security, 20(1):91–93, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801010264.

Anonymous:2001:ITCb

[Ano01v] Anonymous. IFIP technical committee 11. Computers & Security, 20(2):181–183, April 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801002115.

Anonymous:2001:ITCc

[Ano01w] Anonymous. IFIP technical committee 11. Computers & Security, 20(3):271–273, May 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801003157.

Anonymous:2001:ITCd

[Ano01x] Anonymous. IFIP technical committee 11. Computers & Security, 20(4):341–343, July 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801004126.

Anonymous:2001:ITCe

[Ano01y] Anonymous. IFIP technical committee 11. Computers & Security, 20(5):441–443, July 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801005156.

Anonymous:2001:ITCf

[Ano01z] Anonymous. IFIP technical committee 11. Computers & Security, 20(6):539–541, September 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801006150.

Anonymous:2001:ITCg

[Ano01-27] Anonymous. IFIP technical committee 11. Computers & Security, 20(7):635–637, October 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801007131.

Anonymous:2001:ITCh

[Ano01-28] Anonymous. IFIP technical committee 11. Computers & Security, 20(8):739–741, December 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480100815X.

Anonymous:2001:ITCi

[Ano01-29] Anonymous. IFIP technical committee 11. Computers & Security, 21(1):93–95, First Quarter 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802001128.

Anonymous:2001:IBRa

[Ano01-30] Anonymous. International Board of Referees. Computers & Security, 20(1):77–78, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801010240.

Anonymous:2001:IBRb

[Ano01-31] Anonymous. International Board of Referees. Computers & Security, 20(3):253–254, May 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801003121.

Anonymous:2001:IBRc

[Ano01-32] Anonymous. International Board of Referees. Computers & Security, 20(4):329–330, July 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801004102.

Anonymous:2001:IBRd

[Ano01-33] Anonymous. International board of referees. Computers & Security, 20(5):407–408, July 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801005120.

Anonymous:2001:IBRe

[Ano01-34] Anonymous. International Board of Referees. Computers & Security, 20(6):523–524, September 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801006125.

Anonymous:2001:IBRf

[Ano01-35] Anonymous. International Board of Referees. Computers & Security, 20(7):590–591, October 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480100709X.

Anonymous:2001:IBRg

[Ano01-36] Anonymous. International Board of Referees. Computers & Security, 20(8):691–692, December 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801008112.

Anonymous:2001:IBRh

[Ano01-37] Anonymous. International board of referees. Computers & Security, 21(1):61, First Quarter 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802001086.

Anonymous:2001:PM

[Ano01-38] Anonymous. Pages 186–274 (1 May 2001). Computers & Security, 20(3):??, May 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2001:PJa

[Ano01-39] Anonymous. Pages 2–94 (31 January 2001). Computers & Security, 20(1):??, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2001:PQ

[Ano01-40] Anonymous. Pages 2–95 (1st Quarter 2001). Computers & Security, 21(1):??, First Quarter 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2001:PJb

[Ano01-41] Anonymous. Pages 276–343 (31 July 2001). Computers & Security, 20(4):??, July 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2001:PJc

[Ano01-42] Anonymous. Pages 346–446 (1 July 2001). Computers & Security, 20(5):??, July 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2001:PS

[Ano01-43] Anonymous. Pages 448–542 (1 September 2001). Computers & Security, 20(6):??, September 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2001:PO

[Ano01-44] Anonymous. Pages 544–638 (31 October 2001). Computers & Security, 20(7):??, October 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2001:PD

[Ano01-45] Anonymous. Pages 640–742 (1 December 2001). Computers & Security, 20(8):??, December 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2001:PA

[Ano01-46] Anonymous. Pages 96–184 (1 April 2001). Computers & Security, 20(2):??, April 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2001:RP

[Ano01-47]  Anonymous.  Refereed papers.  Computers & Secu-

rity, 20(2):163–164, April 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801002085.

Anonymous:2001:SIV

[Ano01-48] Anonymous. Subject index for volume 19. Computers & Security, 20(3):247–252, May 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480100311X.

Anonymous:2001:UDP

[Ano01-49] Anonymous. U.S. DoD puts up blocks to Code Red. Computers & Security, 20(6):451–467, September 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801006022.

Anonymous:2002:ARA

[Ano02a] Anonymous. Abstracts of recent articles and literature. Computers & Security, 21(2):126–131, March 31, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802002055.

Anonymous:2002:AIV

[Ano02b] Anonymous. Author index for volume 20. Computers & Security, 21(2):133, March 31, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802002079.

Anonymous:2002:CFC

[Ano02c] Anonymous. Calendar of forthcoming conferences and events. Computers & Security, 21(4):343, August 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802004108.

Anonymous:2002:DRIa

[Ano02d] Anonymous. Digest of recent IT security press coverage. Computers & Security, 21(6):532–536, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802010106.

Anonymous:2002:DRIb

[Ano02e] Anonymous. Digest of recent IT security press coverage. Computers & Security, 21(7):629–633, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802011112.

Anonymous:2002:DRIc

[Ano02f] Anonymous. Digest of recent IT security press coverage: Compiled by Bill McKenna. Computers & Security, 21(8):712–716, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802008088.

Anonymous:2002:Eb

[Ano02g] Anonymous. Erratum. Computers & Security, 21(2):143, March 31, 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802002158.

Anonymous:2002:Ea

[Ano02h] Anonymous. Events. Computers & Security, 21(2):132, March 31, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404802002067.

Anonymous:2002:Ec

[Ano02i] Anonymous. Events. Computers & Security, 21(3):251, June 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404802003103.

Anonymous:2002:Ed

[Ano02j] Anonymous. Events. Computers & Security, 21(5):437, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802005114.

Anonymous:2002:Ee

[Ano02k] Anonymous. Events. Computers & Security, 21(6):537, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740480201012X.

Anonymous:2002:Ef

[Ano02l] Anonymous. Events. Computers & Security, 21(7):634, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802011136.

Anonymous:2002:Eg

[Ano02m] Anonymous. Events. Computers & Security, 21(8):717, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802008106.

Anonymous:2002:ECR

[Ano02n] Anonymous. From the editor-in-chief: the Revenger’s tragedy. Computers & Security, 21(6):478–480, October 1, 2002. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404802010015.

Anonymous:2002:GAa

[Ano02o] Anonymous. Guide for authors. Computers & Security, 21(2):141, March 31, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480200216X.

Anonymous:2002:GAb

[Ano02p] Anonymous. Guide for authors. Computers & Security, 21(3): 252, June 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802003115.

Anonymous:2002:GAc

[Ano02q] Anonymous. Guide for authors. Computers & Security, 21(4):344, August 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480200411X.

Anonymous:2002:GAd

[Ano02r] Anonymous. Guide for authors. Computers & Security, 21(5):438, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802005126.

Anonymous:2002:GAe

[Ano02s] Anonymous. Guide for authors. Computers & Security, 21(7):635, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480201115X.

Anonymous:2002:GAf

[Ano02t] Anonymous. Guide for authors. Computers & Security, 21(8):718, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802008118.

Anonymous:2002:ITCa

[Ano02u] Anonymous. IFIP technical committee. Computers & Security, 21(2):193–195, March 31, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802002134.

Anonymous:2002:ITCb

[Ano02v] Anonymous. IFIP technical committee 11. Computers & Security, 21(3):285–287, June 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802003140.

Anonymous:2002:ITCc

[Ano02w] Anonymous. IFIP technical committee 11. Computers & Security, 21(4):377–379, August 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802004169.

Anonymous:2002:ITCd

[Ano02x] Anonymous. IFIP technical committee 11. Computers & Security, 21(5):473–475, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802005175.

Anonymous:2002:ITCe

[Ano02y] Anonymous. IFIP technical committee 11. Computers & Security, 21(6):573–575, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802010179.

Anonymous:2002:ITCf

[Ano02z] Anonymous. IFIP technical committee 11. Computers & Security, 21(7):669–671, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802011197.

Anonymous:2002:ITCg

[Ano02-27] Anonymous. IFIP technical committee 11. Computers & Security, 21(8):761–763, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802008167.

Anonymous:2002:Ba

[Ano02-28] Anonymous. In brief. Computers & Security, 21(4):338–341, August 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802004091.

Anonymous:2002:Bb

[Ano02-29] Anonymous. in brief. Computers & Security, 21(6):532–536, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802010118.

Anonymous:2002:Bc

[Ano02-30] Anonymous. In brief. Computers & Security, 21(7):629–633, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802011124.

Anonymous:2002:BCB

[Ano02-31] Anonymous. In brief: Compiled by Bill McKenna. Computers & Security, 21(8):712–716, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480200809X.

Anonymous:2002:IBRa

[Ano02-32] Anonymous. International Board of Referees. Computers & Security, 21(3):252, June 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802003061.

Anonymous:2002:IBRb

[Ano02-33] Anonymous. International Board of Referees. Computers & Security, 21(4):344, August 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802004121.

Anonymous:2002:IBRc

[Ano02-34] Anonymous. International Board of Referees. Computers & Security, 21(5):438, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802005138.

Anonymous:2002:IBRd

[Ano02-35] Anonymous. International Board of Referees. Computers & Security, 21(7):635, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802011148.

Anonymous:2002:IBRe

[Ano02-36] Anonymous. International Board of Referees. Computers & Security, 21(8):718, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480200812X.

Anonymous:2002:PJ

[Ano02-37] Anonymous. Pages 198–287 (1 June 2002). Computers & Security, 21(3):??, June 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2002:PA

[Ano02-38] Anonymous. Pages 290–379 (1 August 2002). Computers & Security, 21(4):??, August 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2002:POa

[Ano02-39] Anonymous. Pages 382–475 (1 October 2002). Computers & Security, 21(5):??, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2002:POb

[Ano02-40] Anonymous. Pages 478–575 (1 October 2002). Computers & Security, 21(6):??, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2002:PNa

[Ano02-41] Anonymous. Pages 578–671 (November 2002). Computers & Security, 21(7):??, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2002:PNb

[Ano02-42] Anonymous. Pages 674–763 (November 2002). Computers & Security, 21(8):??, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2002:PM

[Ano02-43] Anonymous. Pages 98–195 (31 March 2002). Computers & Security, 21(2):??, March 31, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2002:RPa

[Ano02-44]  Anonymous.  Refereed papers.  Computers & Secu-

rity, 21(2):141, March 31, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802002092.

Anonymous:2002:RPb

[Ano02-45]  Anonymous.  Refereed papers.  Computers & Secu-

rity, 21(6):538, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802010131.

Anonymous:2002:SIV

[Ano02-46] Anonymous. Subject index for volume 20. Computers & Security, 21(2):134–140, March 31, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802002080.

Anonymous:2003:CFC

[Ano03a] Anonymous. Calendar of forthcoming conferences and events. Computers & Security, 22(8):693, December 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803000087.

Anonymous:2003:CSVa

[Ano03b] Anonymous. Computers and security volume 21 — author index. Computers & Security, 22(2):129, February 2003. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404803002104.

Anonymous:2003:CSVb

[Ano03c] Anonymous. Computers and security volume 21 — subject index. Computers & Security, 22(2):130–134, February 2003.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404803002165.

Anonymous:2003:Ea

[Ano03d] Anonymous. Events. Computers & Security, 22(1):43, January 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740480300110X.

Anonymous:2003:Eb

[Ano03e] Anonymous. Events. Computers & Security, 22(2):128, February 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404803002098.

Anonymous:2003:Ed

[Ano03f] Anonymous. Events. Computers & Security, 22(3):212, April 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404803003080.

Anonymous:2003:Ee

[Ano03g] Anonymous. Events. Computers & Security, 22(4):314, May 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404803004085.

Anonymous:2003:Ef

[Ano03h] Anonymous. Events. Computers & Security, 22(5):421, July 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404803005091.

Anonymous:2003:Eg

[Ano03i] Anonymous. Events. Computers & Security, 22(6):521, September 2003. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404803006114.

Anonymous:2003:Eh

[Ano03j] Anonymous. Events. Computers & Security, 22(7):589, October 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404803007077.

Anonymous:2003:EC

[Ano03k] Anonymous. From the editor-in-chief. Computers & Security, 22(1):2–3, January 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803001019.

Anonymous:2003:ECG

[Ano03l] Anonymous. From the editor-in-chief: Gartner ’s prediction concerning intrusion detection systems: sense or nonsense? Computers & Security, 22(6):462–463, September 2003. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404803006011.

Anonymous:2003:ECV

[Ano03m] Anonymous. From the editor-in-chief: Virus and worm trends. Computers & Security, 22(3):174–175, April 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803003018.

Anonymous:2003:GAa

[Ano03n] Anonymous. Guide for authors. Computers & Security, 22(1): 44, January 2003. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404803001111.

Anonymous:2003:GAb

[Ano03o] Anonymous. Guide for authors. Computers & Security, 22(2):135, February 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803002116.

Anonymous:2003:GAc

[Ano03p] Anonymous. Guide for authors. Computers & Security, 22(3): 213, April 2003. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404803003158.

Anonymous:2003:GAd

[Ano03q] Anonymous. Guide for authors. Computers & Security, 22(4): 315, May 2003. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404803004103.

Anonymous:2003:ITCa

[Ano03r] Anonymous. IFIP technical committee 11. Computers & Security, 22(1):73–75, January 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803001159.

Anonymous:2003:ITCb

[Ano03s] Anonymous. IFIP technical committee 11. Computers & Security, 22(2):167–169, February 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803002153.

Anonymous:2003:ITCc

[Ano03t] Anonymous. IFIP technical committee 11. Computers & Security, 22(3):265–267, April 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803003146.

Anonymous:2003:ITCd

[Ano03u] Anonymous. IFIP technical committee 11. Computers & Security, 22(4):361–363, May 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803004140.

Anonymous:2003:ITCe

[Ano03v] Anonymous. IFIP technical committee 11. Computers & Security, 22(5):457–459, July 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803005157.

Anonymous:2003:ITCf

[Ano03w] Anonymous. IFIP technical committee 11. Computers & Security, 22(6):551–553, September 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803006175.

Anonymous:2003:ITCg

[Ano03x] Anonymous. IFIP technical committee 11. Computers & Security, 22(7):647–649, October 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803007144.

Anonymous:2003:ITCh

[Ano03y] Anonymous. IFIP technical committee 11. Computers & Security, 22(8):735–737, December 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803000142.

Anonymous:2003:IBRa

[Ano03z] Anonymous. International Board of Referees. Computers & Security, 22(1):44, January 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803001160.

Anonymous:2003:IBRb

[Ano03-27] Anonymous. International Board of Referees. Computers & Security, 22(2):135, February 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803002177.

Anonymous:2003:IBRc

[Ano03-28] Anonymous. International Board of Referees. Computers & Security, 22(3):213, April 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803003092.

Anonymous:2003:IBRd

[Ano03-29] Anonymous. International Board of Referees. Computers & Security, 22(4):315, May 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803004097.

Anonymous:2003:IBRe

[Ano03-30] Anonymous. International Board of Referees. Computers & Security, 22(5):422, July 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803005108.

Anonymous:2003:IBRf

[Ano03-31] Anonymous. International Board of Referees. Computers & Security, 22(7):590, October 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803007089.

Anonymous:2003:IBRg

[Ano03-32] Anonymous. International Board of Referees. Computers & Security, 22(8):694, December 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803000099.

Anonymous:2003:PA

[Ano03-33] Anonymous. Pages 174–267 (April 2003). Computers & Security, 22(3):??, April 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2003:PJa

[Ano03-34] Anonymous. Pages 2–75 (January 2003). Computers & Security, 22(1):??, January 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2003:PM

[Ano03-35] Anonymous. Pages 270–363 (May 2003). Computers & Security, 22(4):??, May 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2003:PJb

[Ano03-36] Anonymous. Pages 366–459 (July 2003). Computers & Security, 22(5):??, July 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2003:PS

[Ano03-37] Anonymous. Pages 462–553 (September 2003). Computers & Security, 22(6):??, September 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2003:PO

[Ano03-38] Anonymous. Pages 555–650 (October 2003). Computers & Security, 22(7):??, October 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2003:PD

[Ano03-39] Anonymous. Pages 652–737 (December 2003). Computers & Security, 22(8):??, December 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2003:PF

[Ano03-40] Anonymous. Pages 78–169 (February 2003). Computers & Security, 22(2):??, February 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2003:PER

[Ano03-41] Anonymous. Presenting the evidence report: Introduction. Computers & Security, 22(6):479–481, September 2003. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404803006047.

Anonymous:2003:RP

[Ano03-42]  Anonymous.  Refereed papers.  Computers & Secu-

rity, 22(6):522, September 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803006126.

Anonymous:2003:RPG

[Ano03-43] Anonymous. Refereed papers — guide for authors. Computers & Security, 22(5):459, July 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803005169.

Anonymous:2003:SV

[Ano03-44] Anonymous. Security views. Computers & Security, 22(7): 559–569, October 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404803007028.

Anonymous:2004:ECd

[Ano04a] Anonymous. Events calendar. Computers & Security, 23(4): 280, June 2004. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404804001294.

Anonymous:2004:ECf

[Ano04b]  Anonymous.  Events calendar.  Computers & Secu-

rity, 23(6):457, September 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804001993.

Anonymous:2004:ECg

[Ano04c] Anonymous. Events calendar. Computers & Security, 23 (7):547, October 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404804002470.

Anonymous:2004:ECh

[Ano04d]  Anonymous.  Events calendar.  Computers & Secu-

rity, 23(8):647, December 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804002718.

Anonymous:2004:ECa

[Ano04e] Anonymous. Events calender. Computers & Security, 23 (1):29, February 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404804000513.

Anonymous:2004:ECb

[Ano04f] Anonymous. Events calender. Computers & Security, 23(2): 105, March 2004. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404804000781.

Anonymous:2004:ECc

[Ano04g] Anonymous. Events calender. Computers & Security, 23(3): EX1, May 2004. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404804001087.

Anonymous:2004:ECe

[Ano04h] Anonymous. Events calender. Computers & Security, 23(5): 377, July 2004. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404804001592.

Anonymous:2004:PF

[Ano04i] Anonymous. Pages 1–86 (February 2004). Computers & Security, 23(1):??, February 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2004:PMb

[Ano04j] Anonymous. Pages 179–264 (May 2004). Computers & Security, 23(3):??, May 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2004:PJa

[Ano04k] Anonymous. Pages 265–351 (June 2004). Computers & Security, 23(4):??, June 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2004:PJb

[Ano04l] Anonymous. Pages 353–440 (July 2004). Computers & Security, 23(5):??, July 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2004:PS

[Ano04m] Anonymous. Pages 441–530 (September 2004). Computers & Security, 23(6):??, September 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2004:PO

[Ano04n] Anonymous. Pages 531–620 (October 2004). Computers & Security, 23(7):??, October 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2004:PD

[Ano04o] Anonymous. Pages 621–714 (December 2004). Computers & Security, 23(8):??, December 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2004:PMa

[Ano04p] Anonymous. Pages 87–178 (March 2004). Computers & Security, 23(2):??, March 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2004:SVa

[Ano04q] Anonymous. Security views. Computers & Security, 23 (1):3–11, February 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404804000112.

Anonymous:2004:SVb

[Ano04r] Anonymous. Security views. Computers & Security, 23(2):89– 96, March 2004. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404804000306.

Anonymous:2004:SVc

[Ano04s] Anonymous. Security views. Computers & Security, 23(6): 443–452, September 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404804001889.

Anonymous:2004:SVd

[Ano04t] Anonymous. Security views. Computers & Security, 23(8): 623–632, December 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404804002652.

Anonymous:2005:CE

[Ano05a] Anonymous. Calendar of events. Computers & Security, 24(8):EX1, November 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404805001896.

Anonymous:2005:EC

[Ano05b]  Anonymous.  Events calendar.  Computers & Secu-

rity, 24(1):EX1, February 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404805000064.

Anonymous:2005:ECP

[Ano05c] Anonymous. Events calendar, please update, see marked copy. Computers & Security, 24(7):EX1, October 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805001689.

Anonymous:2005:PF

[Ano05d] Anonymous. Pages 1–86 (February 2005). Computers & Security, 24(1):??, February 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2005:PMb

[Ano05e] Anonymous. Pages 175–260 (May 2005). Computers & Security, 24(3):??, May 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2005:PJ

[Ano05f] Anonymous. Pages 261–346 (June 2005). Computers & Security, 24(4):??, June 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2005:PA

[Ano05g] Anonymous. Pages 347–424 (August 2005). Computers & Security, 24(5):??, August 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2005:PS

[Ano05h] Anonymous. Pages 425–504 (September 2005). Computers & Security, 24(6):??, September 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2005:PO

[Ano05i] Anonymous. Pages 505–586 (October 2005). Computers & Security, 24(7):??, October 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2005:PN

[Ano05j] Anonymous. Pages 587–674 (November 2005). Computers & Security, 24(8):??, November 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2005:PMa

[Ano05k] Anonymous. Pages 87–174 (March 2005). Computers & Security, 24(2):??, March 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2005:SV

[Ano05l] Anonymous. Security views. Computers & Security, 24 (4):263–270, June 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404805000520.

Anonymous:2005:SVMc

[Ano05m] Anonymous. Security views — malware update. Computers & Security, 24(3):177–187, May 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805000441.

Anonymous:2005:SVMd

[Ano05n] Anonymous. Security views — malware update. Computers & Security, 24(7):507–515, October 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805001434.

Anonymous:2005:SVMa

[Ano05o] Anonymous. Security views: Malware update. Computers & Security, 24(1):1–9, February 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804002809.

Anonymous:2005:SVMb

[Ano05p] Anonymous. Security views: Malware update. Computers & Security, 24(2):91–98, March 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805000155.

Anonymous:2005:SCB

[Ano05q] Anonymous. Smart card based authentication — any future? Computers & Security, 24(3):188–191, May 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480500043X.

Anonymous:2006:CEa

[Ano06a] Anonymous. Calendar of events. Computers & Security, 25(2):EX1, March 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806000459.

Anonymous:2006:CEb

[Ano06b] Anonymous. Calendar of events. Computers & Security, 25(4): EX1, June 2006. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404806000836.

Anonymous:2006:CEU

[Ano06c] Anonymous. Calendar of events — update. Computers & Security, 25(3):EX1, May 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480600068X.

Anonymous:2006:CEc

[Ano06d] Anonymous. Calender of events. Computers & Security, 25(5): EX1, July 2006. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404806001052.

Anonymous:2006:CPI

[Ano06e] Anonymous. Call for papers-IFIPSEC 2007. Computers & Security, 25(7):551, October 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806001477.

Anonymous:2006:E

[Ano06f] Anonymous. Events. Computers & Security, 25(1):EX1, February 2006. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404806000162.

Anonymous:2006:GAa

[Ano06g] Anonymous. Guide for authors. Computers & Security, 25(6):iv, September 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806001337.

Anonymous:2006:GAb

[Ano06h] Anonymous. Guide for authors. Computers & Security, 25(7):iv–v, October 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806001660.

Anonymous:2006:GAc

[Ano06i] Anonymous. Guide for authors. Computers & Security, 25(8):iv–v, November 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806001969.

Anonymous:2006:ICEa

[Ano06j] Anonymous. IBC — calendar of events. Computers & Security, 25(6):CO3, September 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806001349.

Anonymous:2006:ICEb

[Ano06k] Anonymous. IBC — calendar of events. Computers & Security, 25(7):CO3, October 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806001817.

Anonymous:2006:ICEc

[Ano06l] Anonymous. IBC — calendar of events. Computers & Security, 25(8):CO3, November 2006. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806001970.

Anonymous:2006:IEBa

[Ano06m] Anonymous. IFC — editorial board. Computers & Security, 25(6):CO2, September 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806001271.

Anonymous:2006:IEBb

[Ano06n] Anonymous. IFC — editorial board. Computers & Security, 25(7):CO2, October 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806001611.

Anonymous:2006:IEBc

[Ano06o] Anonymous. IFC — editorial board. Computers & Security, 25(8):CO2, November 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806001908.

Anonymous:2006:ICa

[Ano06p] Anonymous. (iii) contents. Computers & Security, 25(6):iii, September 2006. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404806001301.

Anonymous:2006:ICb

[Ano06q] Anonymous. (iii) contents. Computers & Security, 25(7):iii, October 2006. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404806001647.

Anonymous:2006:ICc

[Ano06r] Anonymous. (iii) contents. Computers & Security, 25(8):iii, November 2006. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404806001933.

Anonymous:2006:MU

[Ano06s] Anonymous. Malware update. Computers & Security, 25 (4):238–246, June 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404806000629.

Anonymous:2006:PF

[Ano06t] Anonymous. Pages 1–78 (February 2006). Computers & Security, 25(1):??, February 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2006:PMb

[Ano06u] Anonymous. Pages 155–236 (May 2006). Computers & Security, 25(3):??, May 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2006:PJa

[Ano06v] Anonymous. Pages 237–314 (June 2006). Computers & Security, 25(4):??, June 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2006:PJb

[Ano06w] Anonymous. Pages 315–392 (July 2006). Computers & Security, 25(5):??, July 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2006:PS

[Ano06x] Anonymous. Pages 393–474 (September 2006). Computers & Security, 25(6):??, September 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2006:PO

[Ano06y] Anonymous. Pages 475–552 (October 2006). Computers & Security, 25(7):??, October 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2006:PN

[Ano06z] Anonymous. Pages 553–616 (November 2006). Computers & Security, 25(8):??, November 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2006:PMa

[Ano06-27] Anonymous. Pages 79–154 (March 2006). Computers & Security, 25(2):??, March 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2006:SVMa

[Ano06-28] Anonymous. Security views — malware update. Computers & Security, 25(1):3–12, February 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805002075.

Anonymous:2006:SVMb

[Ano06-29] Anonymous. Security views — malware update. Computers & Security, 25(2):81–88, March 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806000204.

Anonymous:2006:SVMc

[Ano06-30] Anonymous. Security views — malware update. Computers & Security, 25(3):156–162, May 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806000599.

Anonymous:2006:SVMd

[Ano06-31] Anonymous. Security views — malware update. Computers & Security, 25(5):317–324, July 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806000952.

Anonymous:2006:SVMe

[Ano06-32] Anonymous. Security views — malware update. Computers & Security, 25(6):395–404, September 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001143.

Anonymous:2006:SVMf

[Ano06-33] Anonymous. Security views — malware update. Computers & Security, 25(7):477–485, October 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001465.

Anonymous:2006:SVMg

[Ano06-34] Anonymous. Security views — malware update. Computers & Security, 25(8):555–565, November 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001842.

Anonymous:2007:P

[Ano07a] Anonymous. From the publisher. Computers & Security, 26(6):419, September 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807000776.

Anonymous:2007:GAa

[Ano07b] Anonymous. Guide for authors. Computers & Security, 26(1):V–VI, February 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807000144.

Anonymous:2007:GAb

[Ano07c] Anonymous. Guide for authors. Computers & Security, 26(2): iv–v, March 2007. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404807000338.

Anonymous:2007:GAc

[Ano07d] Anonymous. Guide for authors. Computers & Security, 26(3): iv–v, May 2007. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404807000491.

Anonymous:2007:GAd

[Ano07e] Anonymous. Guide for authors. Computers & Security, 26(4): iv–v, June 2007. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740480700065X.

Anonymous:2007:GAe

[Ano07f] Anonymous. Guide for authors. Computers & Security, 26(5):iv–v, August 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807000879.

Anonymous:2007:GAf

[Ano07g] Anonymous. Guide for authors. Computers & Security, 26(6):iv–v, September 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807001022.

Anonymous:2007:GAg

[Ano07h] Anonymous. Guide for authors. Computers & Security, 26(7–8):iv–v, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807001216.

Anonymous:2007:GAh

[Ano07i] Anonymous. Guide for authors. Computers & Security, 26(7–8):iv–v, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807001216.

Anonymous:2007:GAi

[Ano07j] Anonymous. Guide for authors. Computers & Security, 26(7–8):iv–v, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807001216.

Anonymous:2007:ICE

[Ano07k] Anonymous. IBC — calendar of events. Computers & Security, 26(1):CO3, February 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807000156.

Anonymous:2007:IEBa

[Ano07l] Anonymous. IFC — editorial board. Computers & Security, 26(1):CO2, February 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807000089.

Anonymous:2007:IEBb

[Ano07m] Anonymous. IFC — editorial board. Computers & Security, 26(2):ifc, March 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807000260.

Anonymous:2007:IEBc

[Ano07n] Anonymous. IFC — editorial board. Computers & Security, 26(3):ifc, May 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404807000429.

Anonymous:2007:IEBd

[Ano07o] Anonymous. IFC — editorial board. Computers & Security, 26(4):ifc, June 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404807000582.

Anonymous:2007:IEBe

[Ano07p] Anonymous. IFC — editorial board. Computers & Security, 26(5):ifc, August 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807000806.

Anonymous:2007:IEBf

[Ano07q] Anonymous. IFC — editorial board. Computers & Security, 26(6):ifc, September 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480700096X.

Anonymous:2007:IEBg

[Ano07r] Anonymous. IFC — editorial board. Computers & Security, 26(7–8):ifc, December 2007. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807001149.

Anonymous:2007:IEBh

[Ano07s] Anonymous. IFC — editorial board. Computers & Security, 26(7–8):ifc, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807001149.

Anonymous:2007:IEBi

[Ano07t] Anonymous. IFC — editorial board. Computers & Security, 26(7–8):ifc, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807001149.

Anonymous:2007:ICa

[Ano07u] Anonymous. (iii) contents. Computers & Security, 26(1):iii, February 2007. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404807000119.

Anonymous:2007:ICb

[Ano07v] Anonymous. (iii) contents. Computers & Security, 26(2):iii, March 2007. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404807000296.

Anonymous:2007:ICc

[Ano07w] Anonymous. (iii) contents. Computers & Security, 26(3):iii, May 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404807000454.

Anonymous:2007:ICd

[Ano07x] Anonymous. (iii) contents. Computers & Security, 26(4):iii, June 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404807000612.

Anonymous:2007:ICf

[Ano07y] Anonymous. (iii) contents. Computers & Security, 26(5):iii, August 2007. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404807000831.

Anonymous:2007:ICg

[Ano07z] Anonymous. (iii) contents. Computers & Security, 26(6):iii, September 2007. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404807000995.

Anonymous:2007:ICh

[Ano07-27] Anonymous. (iii) contents. Computers & Security, 26(7–8):iii, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404807001174.

Anonymous:2007:ICi

[Ano07-28] Anonymous. (iii) contents. Computers & Security, 26(7–8):iii, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404807001174.

Anonymous:2007:ICj

[Ano07-29] Anonymous. (iii) contents. Computers & Security, 26(7–8):iii, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404807001174.

Anonymous:2007:MUa

[Ano07-30] Anonymous. Malware update. Computers & Security, 26(2):100–108, March 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807000223.

Anonymous:2007:MUb

[Ano07-31] Anonymous. Malware update. Computers & Security, 26 (3):188–200, May 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404807000387.

Anonymous:2007:PF

[Ano07-32] Anonymous. Pages 1–98 (February 2007). Computers & Security, 26(1):??, February 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2007:PMb

[Ano07-33] Anonymous. Pages 187–266 (May 2007). Computers & Security, 26(3):??, May 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2007:PJ

[Ano07-34] Anonymous. Pages 267–348 (June 2007). Computers & Security, 26(4):??, June 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2007:PA

[Ano07-35] Anonymous. Pages 349–418 (August 2007). Computers & Security, 26(5):??, August 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2007:PS

[Ano07-36] Anonymous. Pages 419–444 (September 2007). Computers & Security, 26(6):??, September 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2007:PDa

[Ano07-37] Anonymous. Pages 445–514 (December 2007). Computers & Security, 26(7–8):??, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2007:PDb

[Ano07-38] Anonymous. Pages 445–514 (December 2007). Computers & Security, 26(7–8):??, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2007:PDc

[Ano07-39] Anonymous. Pages 445–514 (December 2007). Computers & Security, 26(7–8):??, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2007:PMa

[Ano07-40] Anonymous. Pages 99–186 (March 2007). Computers & Security, 26(2):??, March 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2007:SVMb

[Ano07-41] Anonymous. Security views — malware. Computers & Security, 26(4):269–275, June 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807000557.

Anonymous:2007:SVMa

[Ano07-42] Anonymous. Security views — malware update. Computers & Security, 26(1):3–13, February 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806002112.

Anonymous:2008:IEBa

[Ano08a] Anonymous. IFC — editorial board. Computers & Security, 27(1–2):ifc, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000175.

Anonymous:2008:IEBb

[Ano08b] Anonymous. IFC — editorial board. Computers & Security, 27(1–2):ifc, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000175.

Anonymous:2008:IEBc

[Ano08c] Anonymous. IFC — editorial board. Computers & Security, 27(1–2):ifc, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000175.

Anonymous:2008:IEBd

[Ano08d] Anonymous. IFC — editorial board. Computers & Security, 27(3–4):ifc, May/June 2008. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000515.

Anonymous:2008:IEBg

[Ano08e] Anonymous. IFC — editorial board. Computers & Security, 27(5–6):ifc, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000722.

Anonymous:2008:IEBh

[Ano08f] Anonymous. IFC — editorial board. Computers & Security, 27(5–6):ifc, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000722.

Anonymous:2008:IEBi

[Ano08g] Anonymous. IFC — editorial board. Computers & Security, 27(5–6):ifc, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000722.

Anonymous:2008:IEBj

[Ano08h] Anonymous. IFC — editorial board. Computers & Security, 27(7–8):ifc, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000990.

Anonymous:2008:IEBk

[Ano08i] Anonymous. IFC — editorial board. Computers & Security, 27(7–8):ifc, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000990.

Anonymous:2008:IEBl

[Ano08j] Anonymous. IFC — editorial board. Computers & Security, 27(7–8):ifc, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000990.

Anonymous:2008:ICa

[Ano08k] Anonymous. (iii) contents. Computers & Security, 27(1–2): iii, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000199.

Anonymous:2008:ICb

[Ano08l] Anonymous. (iii) contents. Computers & Security, 27(1–2): iii, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000199.

Anonymous:2008:ICc

[Ano08m] Anonymous. (iii) contents. Computers & Security, 27(1–2): iii, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000199.

Anonymous:2008:ICd

[Ano08n] Anonymous. (iii) contents. Computers & Security, 27(3–4):iii, May/June 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000539.

Anonymous:2008:ICg

[Ano08o] Anonymous. (iii) contents. Computers & Security, 27(5–6): iii, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000746.

Anonymous:2008:ICh

[Ano08p] Anonymous. (iii) contents. Computers & Security, 27(5–6): iii, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000746.

Anonymous:2008:ICi

[Ano08q] Anonymous. (iii) contents. Computers & Security, 27(5–6): iii, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000746.

Anonymous:2008:ICj

[Ano08r] Anonymous. (iii) contents. Computers & Security, 27(7–8):iii, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808001016.

Anonymous:2008:ICk

[Ano08s] Anonymous. (iii) contents. Computers & Security, 27(7–8):iii, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808001016.

Anonymous:2008:ICl

[Ano08t] Anonymous. (iii) contents. Computers & Security, 27(7–8):iii, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808001016.

Anonymous:2008:PMa

[Ano08u] Anonymous. Pages 1–62 (March 2008). Computers & Security, 27(1–2):??, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2008:PMb

[Ano08v] Anonymous. Pages 1–62 (March 2008). Computers & Security, 27(1–2):??, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2008:PMc

[Ano08w] Anonymous. Pages 1–62 (March 2008). Computers & Security, 27(1–2):??, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2008:POa

[Ano08x] Anonymous. Pages 123–232 (October 2008). Computers & Security, 27(5–6):??, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2008:POb

[Ano08y] Anonymous. Pages 123–232 (October 2008). Computers & Security, 27(5–6):??, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2008:POc

[Ano08z] Anonymous. Pages 123–232 (October 2008). Computers & Security, 27(5–6):??, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2008:PDa

[Ano08-27] Anonymous. Pages 233–360 (December 2008). Computers & Security, 27(7–8):??, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2008:PDb

[Ano08-28] Anonymous. Pages 233–360 (December 2008). Computers & Security, 27(7–8):??, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2008:PDc

[Ano08-29] Anonymous. Pages 233–360 (December 2008). Computers & Security, 27(7–8):??, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2008:PMJ

[Ano08-30] Anonymous. Pages 63–122 (May–June 2008). Computers & Security, 27(3–4):??, May/June 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2009:CPa

[Ano09a] Anonymous. Call for papers. Computers & Security, 28 (7):728, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404809001035.

Anonymous:2009:CPb

[Ano09b]  Anonymous.  Call for papers.  Computers & Secu-

rity, 28(8):862–863, November 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809001230.

Anonymous:2009:Ca

[Ano09c] Anonymous. Contents. Computers & Security, 28(1–2): iii, February/March 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404808001259.

Anonymous:2009:Cb

[Ano09d] Anonymous. Contents. Computers & Security, 28(3–4):iii, May/June 2009. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404809000212.

Anonymous:2009:Cc

[Ano09e] Anonymous. Contents. Computers & Security, 28(5):iii, July 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404809000340.

Anonymous:2009:Cd

[Ano09f] Anonymous. Contents. Computers & Security, 28(6):iii, September 2009. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404809000649.

Anonymous:2009:Ce

[Ano09g] Anonymous. Contents. Computers & Security, 28(7):iii, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404809000972.

Anonymous:2009:Cf

[Ano09h] Anonymous. Contents. Computers & Security, 28(8):iii, November 2009. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404809001187.

Anonymous:2009:GA

[Ano09i] Anonymous. Guide for authors. Computers & Security, 28(1–2):V–VI, February/March 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808001272.

Anonymous:2009:IEBa

[Ano09j] Anonymous. IFC — editorial board. Computers & Security, 28(1–2):ifc, February/March 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808001223.

Anonymous:2009:IEBb

[Ano09k] Anonymous. IFC — editorial board. Computers & Security, 28(3–4):ifc, May/June 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000182.

Anonymous:2009:IEBc

[Ano09l] Anonymous. IFC — editorial board. Computers & Security, 28(5):ifc, July 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404809000315.

Anonymous:2009:IEBd

[Ano09m] Anonymous. IFC — editorial board. Computers & Security, 28(6):ifc, September 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000613.

Anonymous:2009:IEBe

[Ano09n] Anonymous. IFC — editorial board. Computers & Security, 28(7):ifc, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000947.

Anonymous:2009:IEBf

[Ano09o] Anonymous. IFC — editorial board. Computers & Security, 28(8):ifc, November 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809001151.

Anonymous:2009:ISA

[Ano09p] Anonymous. IFIP SEC 2009 advert. Computers & Security, 28(3–4):246, May/June 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480900025X.

Anonymous:2009:ITAa

[Ano09q] Anonymous. IFIP TC11 — aims and scope. Computers & Security, 28(6):i, September 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000686.

Anonymous:2009:ITAb

[Ano09r] Anonymous. IFIP TC11 — aims and scope. Computers & Security, 28(7):723, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809001023.

Anonymous:2009:ITAc

[Ano09s] Anonymous. IFIP TCII — aims and scope. Computers & Security, 28(8):857, November 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809001229.

Anonymous:2009:ITCa

[Ano09t] Anonymous. IFIP technical committee. Computers & Security, 28(6):ii–v, September 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000650.

Anonymous:2009:ITCb

[Ano09u] Anonymous. IFIP technical committee. Computers & Security, 28(7):724–727, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000984.

Anonymous:2009:ITCc

[Ano09v] Anonymous. IFIP technical committee. Computers & Security, 28(8):858–861, November 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809001199.

Anonymous:2009:PFM

[Ano09w] Anonymous. Pages 1–104 (February-March 2009). Computers & Security, 28(1–2):??, February/March 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2009:PMJ

[Ano09x] Anonymous. Pages 105–246 (May-June 2009). Computers & Security, 28(3–4):??, May/June 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2009:PJ

[Ano09y] Anonymous. Pages 247–324 (July 2009). Computers & Security, 28(5):??, July 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2009:PS

[Ano09z] Anonymous. Pages 325–490 (September 2009). Computers & Security, 28(6):??, September 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2009:PO

[Ano09-27] Anonymous. Pages 491–728 (October 2009). Computers & Security, 28(7):??, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2009:PN

[Ano09-28] Anonymous. Pages 729–864 (November 2009). Computers & Security, 28(8):??, November 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Aron:2002:BNP

[AOG^{+}02] Joan L. Aron, Michael O’Leary, Ronald A. Gove, Shiva Azadegan, and M. Cristina Schneider. The benefits of a notification process in addressing the worsening computer virus problem: Results of a survey and a simulation model. Computers & Security, 21(2):142–163, March 31, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802002109.

Alvarez:2003:NTW

[ÁP03] Gonzalo Álvarez and Slobodan Petrović. A new taxonomy of Web attacks suitable for efficient encoding. Computers & Security, 22(5):435–449, July 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803005121.

Apvrille:2004:XDS

[AP04] A. Apvrille and M. Pourzandi. XML distributed security policy for clusters. Computers & Security, 23(8):649–658, December 2004. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404804002342.

Armstrong:2002:B

[Arm02a]  Andrew Armstrong.  Bugbear.  Computers & Secu-

rity, 21(7):607–608, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802011057.

Armstrong:2002:K

[Arm02b] Andrew Armstrong. KLEZ H. Computers & Security, 21(8):694–695, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802008040.

Artail:2004:PAC

[Art04] Hassan A. Artail. Peer-assisted carrying authentication (PACA). Computers & Security, 23(6):478–488, September 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404804001191.

Aslan:2004:HSM

[Asl04a] Heba K. Aslan. A hybrid scheme for multicast authentication over lossy networks. Computers & Security, 23(8):705–713, December 2004. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404804001798.

Aslan:2004:LAA

[Asl04b] Heba K. Aslan. Logical analysis of AUTHMAC DH: a new protocol for authentication and key distribution. Computers & Security, 23(4):290–299, June 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804000215.

Aslan:2004:SDM

[Asl04c] Heba K. Aslan. A scalable and distributed multicast security protocol using a subgroup-key hierarchy. Computers & Security, 23(4):320–329, June 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804000227.

Aslan:2005:TLC

[Asl05] Heba K. Aslan. Two-level controllers hierarchy for a scalable and distributed multicast security protocol. Computers & Security, 24(5):399–408, August 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805000180.

Aljifri:2003:ITU

[ASP03] Hassan Aljifri, Marcel Smets, and Alexander Pons. IP traceback using header compression. Computers & Security, 22(2):136–151, February 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803002128.

Artail:2006:HHF

[ASS^{+}06] Hassan Artail, Haidar Safa, Malek Sraj, Iyad Kuwatly, and Zaid Al-Masri. A hybrid honeypot framework for improving intrusion detection systems in protecting organizational networks. Computers & Security, 25(4):274–288, June 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404806000587.

Ayoade:2006:SIR

[Ayo06] John Ayoade. Security implications in RFID and authentication processing framework. Computers & Security, 25(3):207– 212, May 2006. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404805001951.

Besnard:2004:CSI

[BA04] Denis Besnard and Budi Arief. Computer security impaired by legitimate users. Computers & Security, 23(3):253–264, May 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404804000161.

Barber:2000:IPK

[Bar00a] Richard Barber. Implementing public key infrastructures in a dynamic business environment. Computers & Security, 19(3):230–233, March 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800886113.

Barber:2000:SMW

[Bar00b] Richard Barber. Security in a mobile world — is Bluetooth the answer? Computers & Security, 19(4):321–325, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404800040190.

Barber:2001:EID

[Bar01a] Richard Barber. The evolution of intrusion detection systems — the next step. Computers & Security, 20(2):132–145, April 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404801002048.

Barber:2001:SMW

[Bar01b] Richard Barber. Security in a mobile world — is Bluetooth the answer? Computers & Security, 20(5):374–379, July 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404801005053.

Breaux:2009:DRM

[BAS09] Travis D. Breaux, Annie I. Antón, and Eugene H. Spafford. A distributed requirements management framework for legal compliance and accountability. Computers & Security, 28(1–2):8–17, February/March 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000679.

Bella:2005:IAS

[BB05] Giampaolo Bella and Stefano Bistarelli. Information assurance for security protocols. Computers & Security, 24(4):322–333, June 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404804002615.

Bergadano:2005:DPL

[BC05] Francesco Bergadano and Davide Cavagnino. Dealing with packet loss in the interactive chained stream authentication protocol. Computers & Security, 24(2):139–146, March 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804001877.

Bergadano:2002:IAM

[BCC02] F. Bergadano, D. Cavagnino, and B. Crispo. Individual authentication in multiparty communications. Computers & Security, 21(8):719–735, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802008131.

Bruschi:2003:QSP

[BCR03] D. Bruschi, A. Curti, and E. Rosti. A quantitative study of public key infrastructures. Computers & Security, 22(1):56– 67, January 2003. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404803001135.

Blyth:2003:SAX

[BCS03] Dr. Andrew Blyth, Dr. Daniel Cunliffe, and Dr. Iain Sutherland. Security analysis of XML usage and XML parsing. Computers & Security, 22(6):494–505, September 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803006072.

Biermann:2001:CID

[BCV01] E. Biermann, E. Cloete, and L. M. Venter. A comparison of intrusion detection systems. Computers & Security, 20(8):676–683, December 1, 2001. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801008069.

Botha:2001:ACD

[BE01] Reinhardt A. Botha and Jan H. P. Eloff. Access control in document-centric workflow systems — an agent-based approach. Computers & Security, 20(6):525–532, September 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404801006137.

Beghdad:2004:MSI

[Beg04] Rachid Beghdad. Modelling and solving the intrusion detection problem in computer networks. Computers & Security, 23(8):687–696, December 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804001725.

Beghdad:2008:CSNa

[Beg08a] Rachid Beghdad. Critical study of neural networks in detecting intrusions. Computers & Security, 27(5–6):168–175, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000357.

Beghdad:2008:CSNb

[Beg08b] Rachid Beghdad. Critical study of neural networks in detecting intrusions. Computers & Security, 27(5–6):168–175, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000357.

Beghdad:2008:CSNc

[Beg08c] Rachid Beghdad. Critical study of neural networks in detecting intrusions. Computers & Security, 27(5–6):168–175, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000357.

Baumer:2004:IPL

[BEP04] David L. Baumer, Julia B. Earp, and J. C. Poindexter. Internet privacy law: a comparison between the United States and the European Union. Computers & Security, 23(5):400– 412, July 2004. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404804000136.

Bequai:2000:AIC

[Beq00a] August Bequai. America’s Internet commerce and the threat of fraud. Computers & Security, 19(8):688–691, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404800080172.

Bequai:2000:RIM

[Beq00b] August Bequai. Romancing the Internet and Management’s quagmire. Computers & Security, 19(7):591–595, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740480007019X.

Bequai:2001:OCG

[Beq01a] August Bequai. Organized crime goes cyber. Computers & Security, 20(6):475–478, September 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801006046.

Bequai:2001:PDE

[Beq01b] August Bequai. Proprietary data: Eldorado in cyber-space. Computers & Security, 20(5):372–373, July 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801005041.

Bequai:2001:SCI

[Beq01c] August Bequai. Software concerns and the Internet. Computers & Security, 20(1):28–30, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801010173.

Bequai:2002:SCI

[Beq02a] August Bequai. Syndicated crime and international terrorism: the lessons of ‘9-11’. Computers & Security, 21(4):333–337, August 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802004078.

Bequai:2002:WCC

[Beq02b] August Bequai. White collar crime: a handmaiden of international tech terrorism. Computers & Security, 21(6):514–519, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802010076.

Bequai:2003:SIM

[Beq03] August Bequai. Safeguards for IT managers and staff under the Sarbanes–Oxley act. Computers & Security, 22(2):124–127, February 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803002086.

Bertino:2003:RMC

[Ber03] Elisa Bertino. RBAC models — concepts and trends. Computers & Security, 22(6):511–514, September 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803006096.

Bernard:2007:ILS

[Ber07] Ray Bernard. Information lifecycle security risk assessment: a tool for closing security gaps. Computers & Security, 26(1):26–30, February 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806002124.

Botha:2009:DME

[BFC09] Reinhardt A. Botha, Steven M. Furnell, and Nathan L. Clarke. From desktop to mobile: Examining the security experience. Computers & Security, 28(3–4):130–137, May/June 2009. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404808001089.

Baiardi:2005:SSV

[BFG^{+}05] Fabrizio Baiardi, Alessandro Falleni, Riccardo Granchi, Fabio Martinelli, Marinella Petrocchi, and Anna Vaccarelli. SEAS, a secure e-voting protocol: Design and implementation. Computers & Security, 24(8):642–652, November 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805001100.

Botha:2006:RSC

[BG06] Reinhardt A. Botha and Tshepo G. Gaadingwe. Reflecting on 20 SEC conferences. Computers & Security, 25(4):247– 256, June 2006. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404806000617.

Bhalla:2003:MCM

[Bha03] Neelam Bhalla. Is the mouse click mighty enough to bring society to its knees? Computers & Security, 22(4):322–336, May 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404803004127.

Baliga:2008:ACRa

[BIC08a] Arati Baliga, Liviu Iftode, and Xiaoxin Chen. Automated containment of rootkits attacks. Computers & Security, 27(7–8):323–334, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000382.

Baliga:2008:ACRb

[BIC08b] Arati Baliga, Liviu Iftode, and Xiaoxin Chen. Automated containment of rootkits attacks. Computers & Security, 27(7–8):323–334, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000382.

Baliga:2008:ACRc

[BIC08c] Arati Baliga, Liviu Iftode, and Xiaoxin Chen. Automated containment of rootkits attacks. Computers & Security, 27(7–8):323–334, December 2008. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000382.

Bishop:2002:TAR

[Bis02] Matt Bishop. Trends in academic research: vulnerabilities analysis and intrusion detection. Computers & Security, 21(7):609–612, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802011069.

Basagiannis:2009:PMC

[BKPA09] Stylianos Basagiannis, Panagiotis Katsaros, Andrew Pombortsis, and Nikolaos Alexiou. Probabilistic model checking for the quantification of DoS security threats. Computers & Security, 28(6):450–465, September 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000042.

Blakley:2002:SI

[Bla02] Bob Blakley. Security insights. Computers & Security, 21(2):120–125, March 31, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802002043.

Biryukov:2005:RAA

[BLP05] Alex Biryukov, Joseph Lano, and Bart Preneel. Recent attacks on alleged SecurID and their practical implications. Computers & Security, 24(5):364–370, August 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805000660.

Balenson:2000:IKR

[BM00] David Balenson and Tom Markham. ISAKMP key recovery extensions. Computers & Security, 19(1):91–99, January 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404800863683.

Bloom:2009:PSE

[BNSZ09] Gedare Bloom, Bhagirath Narahari, Rahul Simha, and Joseph Zambreno. Providing secure execution environments with a last line of defense against Trojan circuit attacks. Computers & Security, 28(7):660–669, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000145.

Booker:2006:REE

[Boo06] Robert Booker. Re-engineering enterprise security. Computers & Security, 25(1):13–17, February 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805002051.

Brooke:2003:FTS

[BP03] Phillip J. Brooke and Richard F. Paige. Fault trees for security system design and analysis. Computers & Security, 22(3):256– 264, April 2003. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404803003134.

Bradbury:2006:MMW

[Bra06a] Danny Bradbury. The metamorphosis of malware writers. Computers & Security, 25(2):89–90, March 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806000198.

Bradbury:2006:MNW

[Bra06b] Danny Bradbury. Microsoft’s new window on security. Computers & Security, 25(6):405–407, September 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001155.

Bradbury:2006:MNS

[Bra06c] Danny Bradbury. Modeling network security. Computers & Security, 25(3):163–164, May 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806000551.

Bradbury:2007:SCIa

[Bra07a] D. Bradbury. The security challenges inherent in VoIP. Computers & Security, 26(7–8):485–487, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807000235.

Bradbury:2007:SCIb

[Bra07b] D. Bradbury. The security challenges inherent in VoIP. Computers & Security, 26(7–8):485–487, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807000235.

Bradbury:2007:SCIc

[Bra07c] D. Bradbury. The security challenges inherent in VoIP. Computers & Security, 26(7–8):485–487, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807000235.

Bradbury:2007:DDR

[Bra07d] Danny Bradbury. Decoding digital rights management. Computers & Security, 26(1):31–33, February 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806002136.

Buchholz:2007:RTLa

[BS07a] Florian Buchholz and Eugene H. Spafford. Run-time label propagation for forensic audit data. Computers & Security, 26(7–8):496–513, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807000922.

Buchholz:2007:RTLb

[BS07b] Florian Buchholz and Eugene H. Spafford. Run-time label propagation for forensic audit data. Computers & Security, 26(7–8):496–513, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807000922.

Buchholz:2007:RTLc

[BS07c] Florian Buchholz and Eugene H. Spafford. Run-time label propagation for forensic audit data. Computers & Security, 26(7–8):496–513, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807000922.

Basso:2009:PMA

[BS09] Alessandro Basso and Stefano Sicco. Preventing massive automated access to web resources. Computers & Security, 28(3–4):174–188, May/June 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808001156.

Bhatt:2009:PMD

[BSC09] Siddharth Bhatt, Radu Sion, and Bogdan Carbunar. A personal mobile DRM manager for smartphones. Computers & Security, 28(6):327–340, September 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000121.

Bakari:2007:BGB

[BTY^{+}07] Jabiri Kuwe Bakari, Charles N. Tarimo, Louise Yngström, Christer Magnusson, and Stewart Kowalski. Bridging the gap between general management and technicians — a case study on ICT security in a developing country. Computers & Security, 26(1):44–55, February 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001568.

Barnard:2000:FAE

[BvS00] Lynette Barnard and Rossouw von Solms. A formalized approach to the effective selection and evaluation of information security controls. Computers & Security, 19(2):185–194, February 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404800878293.

Botha:2003:UFL

[BvS03] Martin Botha and Rossouw von Solms. Utilising fuzzy logic and trend analysis for effective intrusion detection. Computers & Security, 22(5):423–434, July 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480300511X.

Caelli:2002:TLL

[Cae02] William J. Caelli. Trusted … or … trustworthy: the search for a new paradigm for computer and network security. Computers & Security, 21(5):413–420, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802005060.

Carroll:2000:LE

[Car00] John Carroll. Letter to the Editor. Computers & Security, 19(4):357–358, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800040232.

Carter:2001:SYB

[Car01] Bob Carter. Securing your brand in cyberspace. Computers & Security, 20(3):227–238, May 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801003078.

Chebrolu:2005:FDE

[CAT05] Srilatha Chebrolu, Ajith Abraham, and Johnson P. Thomas. Feature deduction and ensemble design of intrusion detection systems. Computers & Security, 24(4):295–307, June 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740480400238X.

Chadwick:2001:ETP

[CB01] David W. Chadwick and Andrew Basden. Evaluating trust in a public key certification authority. Computers & Security, 20(7):592–611, October 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801007106.

Challal:2005:HHC

[CBB05] Yacine Challal, Abdelmadjid Bouabdallah, and Hatem Bettahar. H_{2}A: Hybrid hash-chaining scheme for adaptive multicast source authentication of media-streaming. Computers & Security, 24(1):57–68, February 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804001816.

Carl:2006:WBD

[CBR06] Glenn Carl, Richard R. Brooks, and Suresh Rai. Wavelet based denial-of-service detection. Computers & Security, 25(8):600–615, November 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806001210.

Chan:2001:CTB

[CC01] Chi-Kwong Chan and L. M. Cheng. Cryptanalysis of a timestamp-based password authentication scheme. Computers & Security, 21(1):74–76, First Quarter 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802001104.

Chen:2002:HAC

[CC02] Tzer-Shyong Chen and Yu-Fang Chung. Hierarchical access control based on Chinese remainder theorem and symmetric algorithm. Computers & Security, 21(6):565–570, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404802010167.

Chang:2003:EAA

[CC03] Chin-Chen Chang and Ya-Fen Chang. Efficient anonymous auction protocols with freewheeling bids. Computers & Security, 22(8):728–734, December 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803000130.

Cho:2004:SWS

[CC04] Sanghyun Cho and Sungdeok Cha. SAD: web session anomaly detection based on parameter estimation. Computers & Security, 23(4):312–319, June 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804000264.

Chivers:2009:RPD

[CCC09] Howard Chivers, John A. Clark, and Pau-Chen Cheng. Risk profiles and distributed risk assessment. Computers & Security, 28(7):521–535, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000455.

Chen:2003:EPM

[CCH03] Tzer-Shyong Chen, Yu-Fang Chung, and Gwo-Shiuan Huang. Efficient proxy multisignature schemes based on the elliptic curve cryptosystem. Computers & Security, 22(6):527–534, September 2003. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740480300614X.

Claessens:2002:STO

[CDD^{+}02] Joris Claessens, Valentin Dem, Danny De Cock, Bart Preneel, and Joos Vandewalle. On the security of Today’s online electronic banking systems. Computers & Security, 21(3):253– 265, June 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802003127.

Coetzee:2004:TWS

[CE04] M. Coetzee and J. H. P. Eloff. Towards Web service access control. Computers & Security, 23(7):559–570, October 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804001695.

Clarke:2005:AUM

[CF05] N. L. Clarke and S. M. Furnell. Authentication of users on mobile telephones — a survey of attitudes and practices. Computers & Security, 24(7):519–527, October 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805001446.

Clarke:2007:AUA

[CF07] N. L. Clarke and S. M. Furnell. Advanced user authentication for mobile devices. Computers & Security, 26(2):109–119, March 2007. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404806001428.

Clarke:2002:ASA

[CFRR02] N. L. Clarke, S. M. Furnell, P. M. Rodwell, and P. L. Reynolds. Acceptance of subscriber authentication methods for mobile telephony devices. Computers & Security, 21(3):220–228, June 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802003048.

Clarke:2000:SCH

[CG00] Andy Clarke and Vince Gallo. The software colander — holes in messaging. Computers & Security, 19(8):692–697, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404800080184.

Cournane:2004:ATU

[CH04] Allister Cournane and Ray Hunt. An analysis of the tools used for the generation and prevention of spam. Computers & Security, 23(2):154–166, March 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804000148.

Carpinter:2006:TNR

[CH06] James Carpinter and Ray Hunt. Tightening the net: a review of current and next generation spam filtering tools. Computers & Security, 25(8):566–578, November 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806000939.

Chen:2004:EAS

[Che04] Tzer-Shyong Chen. An English auction scheme in the online transaction environment. Computers & Security, 23(5):389– 399, July 2004. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740480400094X.

Chien:2008:PAUa

[Chi08a] Hung-Yu Chien. Practical anonymous user authentication scheme with security proof. Computers & Security, 27(5–6):216–223, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000291.

Chien:2008:PAUb

[Chi08b] Hung-Yu Chien. Practical anonymous user authentication scheme with security proof. Computers & Security, 27(5–6):216–223, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000291.

Chien:2008:PAUc

[Chi08c] Hung-Yu Chien. Practical anonymous user authentication scheme with security proof. Computers & Security, 27(5–6):216–223, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000291.

Cone:2007:VGC

[CITN07] Benjamin D. Cone, Cynthia E. Irvine, Michael F. Thompson, and Thuy D. Nguyen. A video game for cyber security training and awareness. Computers & Security, 26(1):63–72, February 2007. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404806001556.

Chien:2003:NHA

[CJ03] Hung-Yu Chien and Jinn-Ke Jan. New hierarchical assignment without public key cryptography. Computers & Security, 22(6):523–526, September 2003. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803006138.

Chen:2004:DSA

[CJC04] Yu-Yi Chen, Jinn-Ke Jan, and Chin-Ling Chen. The design of a secure anonymous Internet voting system. Computers & Security, 23(4):330–337, June 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804000720.

Chien:2002:EPS

[CJT02] Hung-Yu Chien, Jinn-Ke Jan, and Yuh-Min Tseng. An efficient and practical solution to remote authentication: Smart card. Computers & Security, 21(4):372–375, August 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404802004157.

Cohen:2003:LAT

[CK03] Fred Cohen and Deanna Koike. Leading attackers through attack graphs with deceptions. Computers & Security, 22 (5):402–411, July 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404803005066.

Chang:2003:FDA

[CL03] Chin-Chen Chang and Yeu-Pong Lai. A flexible dateattachment scheme on e-cash. Computers & Security, 22(2): 160–166, February 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404803002141.

Chiou:2005:MBD

[CL05] Che Wun Chiou and Chiou-Yng Lee. Multiplexer-based double-exponentiation for normal basis of GF(2^{m}). Computers & Security, 24(1):83–86, February 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804002421.

Chang:2006:AVM

[CL06] Chin-Chen Chang and Jung-San Lee. An anonymous voting mechanism based on the key exchange protocol. Computers & Security, 25(4):307–314, June 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806000320.

Chen:2004:CDM

[CLC04] Li-Chiou Chen, Thomas A. Longstaff, and Kathleen M. Carley. Characterization of defense mechanisms against distributed denial of service attacks. Computers & Security, 23(8):665–678, December 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804001774.

Choi:2009:FDV

[CLK09] Hyunsang Choi, Heejo Lee, and Hyogon Kim. Fast detection and visualization of network attacks on parallel coordinates. Computers & Security, 28(5):276–288, July 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808001363.

Cao:2005:RRB

[CLX05] Tianjie Cao, Dongdai Lin, and Rui Xue. A randomized RSAbased partially blind signature scheme for electronic cash. Computers & Security, 24(1):44–49, February 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804001737.

Coles:2003:OIR

[CM03] Robert S. Coles and Rolf Moulton. Operationalizing IT risk management. Computers & Security, 22(6):487–493, September 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404803006060.

Chen:2007:MCM

[CMY07] Xiaoyan Chen, Bobby N. W. Ma, and Cungang Yang. M-CLIQUES: Modified CLIQUES key agreement for secure multicast. Computers & Security, 26(3):238–245, May 2007. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404806001866.

Cetin:2009:NSA

[CO09] Ozdemir Cetin and A. Turan Ozcerit. A new steganography algorithm based on color histograms for data embedding into raw video streams. Computers & Security, 28(7):670–682, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740480900042X.

Cohen:2000:MSS

[Coh00] Fred Cohen. A mathematical structure of simple defensive network deceptions. Computers & Security, 19(6):520–528, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404800060260.

Collins:2000:TCP

[Col00] Michael Collins. Telecommunications crime — Part 3. Computers & Security, 19(2):141–148, February 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800878244.

Cho:2003:EAD

[CP03] Sung-Bae Cho and Hyuk-Jang Park. Efficient anomaly detection by modeling privilege flows using hidden Markov model. Computers & Security, 22(1):45–55, January 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803001123.

Curtis-Raleigh:2002:B

[CR02] Mark Curtis-Raleigh. In brief. Computers & Security, 21(3):246–250, June 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802003085.

Coull:2009:DIC

[CS09] Scott E. Coull and Boleslaw K. Szymanski. On the development of an internetwork-centric defense for scanning worms. Computers & Security, 28(7):637–647, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000777.

Carbonell:2009:SMP

[CSL09] Mildrey Carbonell, José María Sierra, and Javier Lopez. Secure multiparty payment with an intermediary entity. Computers & Security, 28(5):289–300, July 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808001351.

Cusick:2001:CLP

[Cus01] Thomas W. Cusick. Computer licence plates. Computers & Security, 20(5):392–394, July 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801005089.

Chang:2005:IDI

[CY05] David B. Chang and Carl S. Young. Infection dynamics on the Internet. Computers & Security, 24(4):280–286, June 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404805000453.

Chen:2003:AEY

[CZ03] Ke-Fei Chen and Sheng Zhong. Attacks on the (enhanced) Yang–Shieh authentication. Computers & Security, 22(8):725–727, December 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/

S0167404803000129. See erratum [McK04].

Chen:2005:MBI

[CZG^{+}05] Xiu-Zhen Chen, Qing-Hua Zheng, Xiao-Hong Guan, Chen-Guang Lin, and Jie Sun. Multiple behavior information fusion based quantitative threat evaluation. Computers & Security, 24(3):218–231, May 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804002135.

David:2002:GS

[Dav02a] Jon David. Giga security. Computers & Security, 21(8): 696–700, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404802008052.

David:2002:PEW

[Dav02b] Jon David. Policy enforcement in the workplace. Computers & Security, 21(6):506–513, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802010064.

David:2003:SF

[Dav03] Jon David. Security in a flash. Computers & Security, 22(1):29–33, January 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803001068.

Dodge:2007:PUS

[DCF07] Ronald C. Dodge, Curtis Carver, and Aaron J. Ferguson. Phishing for user security awareness. Computers & Security, 26(1):73–80, February 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001581.

Dlamini:2009:ISM

[DEE09] M. T. Dlamini, J. H. P. Eloff, and M. M. Eloff. Information security: the moving target. Computers & Security, 28(3–4):189–198, May/June 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808001168.

DeMaio:2002:GTC

[DeM02] Harry DeMaio. Global trust, certification and (ISC). Computers & Security, 21(8):701–704, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802008064.

Doherty:2006:AIS

[DF06] Neil F. Doherty and Heather Fulford. Aligning the information security policy with the strategic information systems plan. Computers & Security, 25(1):55–63, February 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805001720.

Dantu:2009:ICS

[DFSC09] Ram Dantu, Sonia Fahmy, Henning Schulzrinne, and João Cangussu. Issues and challenges in securing VoIP. Computers & Security, 28(8):743–753, November 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000480.

Dasgupta:2005:CAB

[DGY^{+}05] D. Dasgupta, F. Gonzalez, K. Yallapu, J. Gomez, and R. Yarramsettii. CIDS: an agent-based intrusion detection system. Computers & Security, 24(5):387–398, August 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404805000179.

Dhillon:2001:VST

[Dhi01] Gurpreet Dhillon. Violation of safeguards by trusted personnel and understanding related information security concerns. Computers & Security, 20(2):165–172, April 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801002097.

Drevin:2007:VFA

[DKS07] L. Drevin, H. A. Kruger, and T. Steyn. Value-focused assessment of ICT security awareness in an academic environment. Computers & Security, 26(1):36–43, February 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001544.

Dhillon:2001:CCT

[DM01] Gurpreet Dhillon and Steve Moores. Computer crimes: theorizing about the enemy within. Computers & Security, 20(8):715–723, December 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801008136.

Dragoni:2009:WHA

[DMWS09] N. Dragoni, F. Massacci, T. Walter, and C. Schaefer. What the heck is this application doing? — a securityby-contract architecture for pervasive services. Computers & Security, 28(7):566–577, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000716.

Doughty:2003:IES

[Dou03] Ken Doughty. Implementing enterprise security: a case study. Computers & Security, 22(2):99–114, February 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803002050.

Dimitriadis:2006:IMP

[DP06] Christos K. Dimitriadis and Despina Polemi. An identity management protocol for Internet applications over 3G mobile networks. Computers & Security, 25(1):45–51, February 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404805001768.

DeMara:2004:MNT

[DR04] Ronald F. DeMara and Adam J. Rocke. Mitigation of network tampering using dynamic dispatch of mobile agents. Computers & Security, 23(1):31–42, February 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804000689.

Deepthi:2009:DIA

[DS09] P. P. Deepthi and P. S. Sathidevi. Design, implementation and analysis of hardware efficient stream ciphers using LFSR based hash functions. Computers & Security, 28(3–4):229–241, May/June 2009. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808001193.

Das:2006:NRU

[DSGP06] Manik Lal Das, Ashutosh Saxena, Ved P. Gulati, and Deepak B. Phatak. A novel remote user authentication scheme using bilinear pairings. Computers & Security, 25(3):184–189, May 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404805001525.

Serafim:2004:RRF

[dSSW04] Vinícius da Silveira Serafim and Raul Fernando Weber. Restraining and repairing file system damage through file integrity control. Computers & Security, 23(1):52–62, February 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404804000665.

Du:2006:NAC

[DSX06] Wenliang Du, Mingdong Shang, and Haizhi Xu. A novel approach for computer security education using Minix instructional operating system. Computers & Security, 25(3):190– 200, May 2006. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404805001550.

Ecclestone:2001:ARAa

[Ecc01a] Rachel Ecclestone. Abstracts of recent articles and literature. Computers & Security, 20(7):585–588, October 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404801007076.

Ecclestone:2001:ARAb

[Ecc01b] Rachel Ecclestone. Abstracts of recent articles and literature. Computers & Security, 20(8):684–687, December 1, 2001. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404801008070.

Ecclestone:2001:ARAc

[Ecc01c] Rachel Ecclestone. Abstracts of recent articles and literature. Computers & Security, 21(1):42–45, First Quarter 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404802001050.

Ecclestone:2001:AR

[Ecc01d] Rachel Ecclestone. ACSAC 2001 review. Computers & Security, 21(1):47–60, First Quarter 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802001074.

Ecclestone:2001:BRa

[Ecc01e] Rachel Ecclestone. Book reviews. Computers & Security, 20(8):688, December 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801008082.

Ecclestone:2001:BRb

[Ecc01f] Rachel Ecclestone. Book reviews. Computers & Security, 20(8):688, December 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801008094.

Ecclestone:2001:B

[Ecc01g] Rachel Ecclestone. In brief. Computers & Security, 21(1):42–44, First Quarter 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802001141.

Edelson:2003:SIW

[Ede03] Eve Edelson. The 419 scam: information warfare on the spam front and a proposal for local filtering. Computers & Security, 22(5):392–401, July 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803005054.

Edge:2009:SSB

[ES09] Michael Edward Edge and Pedro R. Falcone Sampaio. A survey of signature based methods for financial fraud detection. Computers & Security, 28(6):381–394, September 2009. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809000091.

Elgohary:2006:DES

[ESZ06] Ashraf Elgohary, Tarek S. Sobh, and M. Zaki. Design of an enhancement for SSL/TLS protocols. Computers & Security, 25(4):297–306, June 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806000319.

Evans:2001:CRG

[Eva01] Paul Evans. Conference report — the global forum for law enforcement and national security, Edinburgh, 2001. Computers & Security, 20(5):395–399, July 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801005090.

Eloff:2000:ISMa

[EvS00a] M. M. Eloff and S. H. von Solms. Information security management: a hierarchical framework for various approaches. Computers & Security, 19(3):243–256, March 1, 2000. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800886137.

Eloff:2000:ISMb

[EvS00b] M. M. Eloff and S. H. von Solms. Information security management: an approach to combine process certification and product evaluation. Computers & Security, 19(8):698–709, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800080196.

Ford:2006:PCR

[FBB06] Richard Ford, Mark Bush, and Alexander Bulatov. Predation and the cost of replication: New approaches to malware prevention? Computers & Security, 25(4):257–264, June 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404806000289.

Furnell:2007:ASP

[FBP07] S. M. Furnell, P. Bryant, and A. D. Phippen. Assessing the security perceptions of personal Internet users. Computers & Security, 26(5):410–417, August 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807000363.

Flowerday:2006:CAT

[FBV06] S. Flowerday, A. W. Blundell, and R. Von Solms. Continuous auditing technologies and models: a discussion. Computers & Security, 25(5):325–331, July 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806000964.

Fan:2005:RRA

[FCZ05] Chun-I Fan, Yung-Cheng Chan, and Zhi-Kai Zhang. Robust remote authentication scheme with smart cards. Computers & Security, 24(8):619–628, November 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805000477.

Furnell:2000:ASS

[FDIR00] S. M. Furnell, P. S. Dowland, H. M. Illingworth, and P. L. Reynolds. Authentication and supervision: a survey of user attitudes. Computers & Security, 19(6):529–539, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404800060272.

Feng:2004:PII

[FGG^{+}04] Li Feng, Xiaohong Guan, Sangang Guo, Yan Gao, and Peini Liu. Predicting the intrusion intentions by observing system call sequences. Computers & Security, 23(3):241–252, May 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404804000732.

Finne:2000:ISR

[Fin00] Thomas Finne. Information systems risk management: Key concepts and business processes. Computers & Security, 19(3):234–242, March 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800886125.

Fisher:2003:LV

[Fis03] Mark Fisher. LIRVA virus. Computers & Security, 22(1):41– 42, January 2003. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404803001093.

Furnell:2006:CUU

[FJK06] S. M. Furnell, A. Jusoh, and D. Katsabas. The challenges of understanding and using security: a survey of end-users. Computers & Security, 25(1):27–35, February 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805002038.

Frantzen:2001:FUV

[FKSF01] Mike Frantzen, Florian Kerschbaum, E. Eugene Schultz, and Sonia Fahmy. A framework for understanding vulnerabilities in firewalls using a dataflow model of firewall internals. Computers & Security, 20(3):263–270, May 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801003145.

Fischer:2005:CCI

[FL05] W. Fischer and N. Lepperhoff. Can critical infrastructures rely on the Internet? Computers & Security, 24(6):485–491, September 2005. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404805000507.

Fan:2002:ETB

[FLZ02] Lei Fan, Jian-Hua Li, and Hong-Wen Zhu. An enhancement of timestamp-based password authentication scheme. Computers & Security, 21(7):665–667, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802011185.

Foote:2005:BSO

[FN05] P. Foote and T. Neudenberger. Beyond Sarbanes–Oxley compliance. Computers & Security, 24(7):516–518, October 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404805001070.

Franklin:2001:PWS

[Fra01] Iain Franklin. Protecting the Web server and applications. Computers & Security, 20(1):31–35, January 31, 2001. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404801010185.

Furnell:2008:SBBa

[FTP08a] Steven Furnell, Valleria Tsaganidi, and Andy Phippen. Security beliefs and barriers for novice Internet users. Computers & Security, 27(7–8):235–240, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000370.

Furnell:2008:SBBb

[FTP08b] Steven Furnell, Valleria Tsaganidi, and Andy Phippen. Security beliefs and barriers for novice Internet users. Computers & Security, 27(7–8):235–240, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000370.

Furnell:2008:SBBc

[FTP08c] Steven Furnell, Valleria Tsaganidi, and Andy Phippen. Security beliefs and barriers for novice Internet users. Computers & Security, 27(7–8):235–240, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000370.

Furnell:2005:WUC

[Fur05] Steven Furnell. Why users cannot use security. Computers & Security, 24(4):274–279, June 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805000532.

Furnell:2007:AWPa

[Fur07a] Steven Furnell. An assessment of website password practices. Computers & Security, 26(7–8):445–451, December 2007. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404807001083.

Furnell:2007:AWPb

[Fur07b] Steven Furnell. An assessment of website password practices. Computers & Security, 26(7–8):445–451, December 2007. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404807001083.

Furnell:2007:AWPc

[Fur07c] Steven Furnell. An assessment of website password practices. Computers & Security, 26(7–8):445–451, December 2007. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404807001083.

Furnell:2007:IWI

[Fur07d] Steven Furnell. IFIP workshop — information security culture. Computers & Security, 26(1):35, February 2007. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404806001854.

Furnell:2007:MSU

[Fur07e] Steven Furnell. Making security usable: Are things improving? Computers & Security, 26(6):434–443, September 2007.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404807000788.

Friess:2005:TE

[FVA05] Nathan Friess, Ryan Vogt, and John Aycock. Timing is everything. Computers & Security, 24(8):599–603, November 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404805001707.

Flowerday:2005:RTI

[FvS05] Stephen Flowerday and Rossouw von Solms. Real-time information integrity = system integrity + data integrity + continuous assurances. Computers & Security, 24(8):604–613, November 2005. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404805001458.

Ganame:2008:GSAa

[GBBS08a] Abdoul Karim Ganame, Julien Bourgeois, Renaud Bidou, and Francois Spies. A global security architecture for intrusion detection on computer networks. Computers & Security, 27(1–2):30–47, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000047.

Ganame:2008:GSAb

[GBBS08b] Abdoul Karim Ganame, Julien Bourgeois, Renaud Bidou, and Francois Spies. A global security architecture for intrusion detection on computer networks. Computers & Security, 27(1–2):30–47, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000047.

Ganame:2008:GSAc

[GBBS08c] Abdoul Karim Ganame, Julien Bourgeois, Renaud Bidou, and Francois Spies. A global security architecture for intrusion detection on computer networks. Computers & Security, 27(1–2):30–47, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000047.

Ghaeb:2009:OMT

[GC09] J. A. Ghaeb and J. Chebil. An oblique-matrix technique for data integrity assurance. Computers & Security, 28 (1–2):94–99, February/March 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000977.

Gordon:2002:C

[GF02] Sarah Gordon and Richard Ford. Cyberterrorism? Computers & Security, 21(7):636–647, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802011161.

Gritzalis:2009:Ec

[GF09] Dimitris Gritzalis and Steven Furnell. Editorial. Computers & Security, 28(7):491–492, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480900087X.

Gowadia:2005:PPA

[GFV05] Vaibhav Gowadia, Csilla Farkas, and Marco Valtorta. PAID: a probabilistic agent-based intrusion detection system. Computers & Security, 24(7):529–545, October 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805001033.

Gritzalis:2009:Ed

[GK09a] Dimitris Gritzalis and Tom Karygiannis. Editorial. Computers & Security, 28(8):729–730, November 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809001114.

Gritzalis:2009:Ea

[GK09b] Dimitris Gritzalis and Sokratis Katsikas. Editorial. Computers & Security, 28(5):247, July 2009. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000388.

Gerdes:2009:MDC

[GKH09] John H. Gerdes, Joakim Kalvenes, and Chin-Tser Huang. Multi-dimensional credentialing using veiled certificates: Protecting privacy in the face of regulatory reporting requirements. Computers & Security, 28(5):248–259, July 2009. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404808001387.

Gritzalis:2000:SEM

[GKL^{+}00] Stefanos Gritzalis, Sokratis K. Katsikas, Dimitrios Lekkas, Konstantinos Moulinos, and Eleni Polydorou. Securing the electronic market: the KEYSTONE public key infrastructure architecture. Computers & Security, 19(8):731–746, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404800080226.

Goyal:2006:NPC

[GKS^{+}06] Vipul Goyal, Virendra Kumar, Mayank Singh, Ajith Abraham, and Sugata Sanyal. A new protocol to counter online dictionary attacks. Computers & Security, 25(2):114–120, March 2006. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404805001537.

Gritzalis:2009:Eb

[GL09] Dimitris Gritzalis and Javier Lopez. Editorial. Computers & Security, 28(6):325–326, September 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000431.

Geneiatakis:2008:OBPa

[GLK08a] Dimitris Geneiatakis, Costas Lambrinoudakis, and Georgios Kambourakis. An ontology-based policy for deploying secure SIP-based VoIP services. Computers & Security, 27(7–8):285–297, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000424.

Geneiatakis:2008:OBPb

[GLK08b] Dimitris Geneiatakis, Costas Lambrinoudakis, and Georgios Kambourakis. An ontology-based policy for deploying secure SIP-based VoIP services. Computers & Security, 27(7–8):285–297, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000424.

Geneiatakis:2008:OBPc

[GLK08c] Dimitris Geneiatakis, Costas Lambrinoudakis, and Georgios Kambourakis. An ontology-based policy for deploying secure SIP-based VoIP services. Computers & Security, 27(7–8):285–297, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000424.

Guo:2008:CSTa

[GLMZ08a] Hua Guo, Zhoujun Li, Yi Mu, and Xiyong Zhang. Cryptanalysis of simple three-party key exchange protocol. Computers & Security, 27(1–2):16–21, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000035.

Guo:2008:CSTb

[GLMZ08b] Hua Guo, Zhoujun Li, Yi Mu, and Xiyong Zhang. Cryptanalysis of simple three-party key exchange protocol. Computers & Security, 27(1–2):16–21, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000035.

Guo:2008:CSTc

[GLMZ08c] Hua Guo, Zhoujun Li, Yi Mu, and Xiyong Zhang. Cryptanalysis of simple three-party key exchange protocol. Computers & Security, 27(1–2):16–21, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000035.

Gupta:2000:PKI

[GM00] Sarbari Gupta and Stephen M. Matyas. Public key infrastructure: Analysis of existing and needed protocols and object formats for key recovery. Computers & Security, 19(1):56–68, January 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404800863646.

Gritzalis:2008:SOSa

[GM08a] Dimitris Gritzalis and Yannis Mallios. A SIP-oriented SPIT management framework. Computers & Security, 27(5–6): 136–153, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404808000333.

Gritzalis:2008:SOSb

[GM08b] Dimitris Gritzalis and Yannis Mallios. A SIP-oriented SPIT management framework. Computers & Security, 27(5–6): 136–153, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404808000333.

Gritzalis:2008:SOSc

[GM08c] Dimitris Gritzalis and Yannis Mallios. A SIP-oriented SPIT management framework. Computers & Security, 27(5–6): 136–153, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404808000333.

Golic:2001:HCC

[Gol01] Jovan Dj. Golić. How to construct cryptographic primitives from stream ciphers. Computers & Security, 20(1):79–89, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404801010252.

Graham:2001:NPS

[Gra01] Robert Graham. NIDS — pattern search vs. protocol decode. Computers & Security, 20(1):37–41, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404801010197.

Gritzalis:2002:PRS

[Gri02] Dimitris A. Gritzalis. Principles and requirements for a secure e-voting system. Computers & Security, 21(6):539–556, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802010143.

Goring:2007:AKM

[GRJ07] Stuart P. Goring, Joseph R. Rabaiotti, and Antonia J. Jones. Anti-keylogging measures for secure Internet login: an example of the law of unintended consequences. Computers & Security, 26(6):421–426, September 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807000569.

Gonzalez:2009:ROS

[GRNR09] Jorge Fontenla González, Manuel Caeiro Rodríguez, Martín Llamas Nistal, and Luis Anido Rifón. Reverse OAuth: a solution to achieve delegated authorizations in single sign-on e-learning systems. Computers & Security, 28(8):843–856, November 2009. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404809000571.

Guven:2003:UUK

[GS03] Aykut Guven and Ibrahim Sogukpinar. Understanding users’ keystroke patterns for computer access security. Computers & Security, 22(8):695–706, December 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803000105.

Geetha:2009:BIS

[GSK09] S. Geetha, Siva S. Sivatha Sindhu, and N. Kamaraj. Blind image steganalysis based on content independent statistical measures maximizing the specificity and sensitivity of the system. Computers & Security, 28(7):683–697, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000285.

Gaiman:2007:PPP

[GSN07] Michael Gaiman, Rahul Simha, and Bhagirath Narahari. Privacy-preserving programming using sython. Computers & Security, 26(2):130–136, March 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001386.

Gonzalez-Talavan:2006:SCS

[GT06] Guillermo González-Talaván. A simple, configurable SMTP anti-spam filter: Greylists. Computers & Security, 25(3):229– 236, May 2006. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404806000332.

Garcia-Teodoro:2009:ABN

[GTDVMFV09] P. García-Teodoro, J. Díaz-Verdejo, G. Maciá-Fernández, and E. Vázquez. Anomaly-based network intrusion detection: Techniques, systems and challenges. Computers & Security, 28(1–2):18–28, February/March 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000692.

Gupta:2000:CKR

[Gup00] Sarbari Gupta. A common key recovery block format: Promoting interoperability between dissimilar key recovery mechanisms. Computers & Security, 19(1):41–47, January 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800863622.

Geneiatakis:2009:UBF

[GVL09] Dimitris Geneiatakis, Nikos Vrakas, and Costas Lambrinoudakis. Utilizing bloom filters for detecting flooding attacks against SIP based services. Computers & Security, 28(7):578–591, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000443.

Gerber:2001:RAS

[GvS01] Mariana Gerber and Rossouw von Solms. From risk analysis to security requirements. Computers & Security, 20(7):577–584, October 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404801007064.

Gerber:2005:MRI

[GvS05] Mariana Gerber and Rossouw von Solms. Management of risk in the information age. Computers & Security, 24(1):16–30, February 2005. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404804002780.

Gerber:2008:ISRa

[GvS08a] Mariana Gerber and Rossouw von Solms. Information security requirements — interpreting the legal aspects. Computers & Security, 27(5–6):124–135, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000461.

Gerber:2008:ISRb

[GvS08b] Mariana Gerber and Rossouw von Solms. Information security requirements — interpreting the legal aspects. Computers & Security, 27(5–6):124–135, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000461.

Gerber:2008:ISRc

[GvS08c] Mariana Gerber and Rossouw von Solms. Information security requirements — interpreting the legal aspects. Computers & Security, 27(5–6):124–135, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000461.

Gu:2009:HBJ

[GYLC09] Xiaozhuo Gu, Jianzu Yang, Julong Lan, and Zhenhuan Cao. Huffman-based join-exit-tree scheme for contributory key management. Computers & Security, 28(1–2):29–39, February/March 2009. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000680.

Hernandez-Ardieta:2008:OFEa

[HAGTA08a] Jorge L. Hernandez-Ardieta, Ana I. Gonzalez-Tablas, and Benjamin Ramos Alvarez. An optimistic fair exchange protocol based on signature policies. Computers & Security, 27(7–8):309–322, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000436.

Hernandez-Ardieta:2008:OFEb

[HAGTA08b] Jorge L. Hernandez-Ardieta, Ana I. Gonzalez-Tablas, and Benjamin Ramos Alvarez. An optimistic fair exchange protocol based on signature policies. Computers & Security, 27(7–8):309–322, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000436.

Hernandez-Ardieta:2008:OFEc

[HAGTA08c] Jorge L. Hernandez-Ardieta, Ana I. Gonzalez-Tablas, and Benjamin Ramos Alvarez. An optimistic fair exchange protocol based on signature policies. Computers & Security, 27(7–8):309–322, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000436.

Hancock:2000:AOW

[Han00a] Bill Hancock. Agencies are “Own worst Enemy”. Computers & Security, 19(4):298, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800040062.

Hancock:2000:AEC

[Han00b] Bill Hancock. American express creates disposable credit card numbers. Computers & Security, 19(7):571–572, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404800070036.

Hancock:2000:NBY

[Han00c] Bill Hancock. And now, the bug-in-your-site award goes to …. Computers & Security, 19(5):404, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800050240.

Hancock:2000:AGB

[Han00d] Bill Hancock. AOL gets busted — again. Computers & Security, 19(5):402, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800050215.

Hancock:2000:ASD

[Han00e] Bill Hancock. Apache site defaced. Computers & Security, 19(4):303–305, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800040116.

Hancock:2000:BRS

[Han00f] Bill Hancock. Baltimore releases security developer freeware. Computers & Security, 19(7):572–573, November 1, 2000. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740480007005X.

Hancock:2000:CCC

[Han00g] Bill Hancock. Canadian credit card conundrum caused by a cracker. Computers & Security, 19(5):397–398, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800050161.

Hancock:2000:CTM

[Han00h] Bill Hancock. Canadian teen Mafiaboy pleads guilty. Computers & Security, 19(8):669, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480008007X.

Hancock:2000:CSH

[Han00i] Bill Hancock. Cellular security hazards. Computers & Security, 19(7):579, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800070127.

Hancock:2000:CCR

[Han00j] Bill Hancock. CIA chat room causes grief. Computers & Security, 19(8):674–676, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800080135.

Hancock:2000:CGS

[Han00k] Bill Hancock. CIO’s get serious about best security practices. Computers & Security, 19(5):388, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800050069.

Hancock:2000:CSS

[Han00l] Bill Hancock. CMU sets up a sister to CERT. Computers & Security, 19(5):400–401, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800050197.

Hancock:2000:CR

[Han00m] Bill Hancock. Cyberstalking on the rise. Computers & Security, 19(4):307–308, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480004013X.

Hancock:2000:DDG

[Han00n] Bill Hancock. DDoS defence gets regulatory. Computers & Security, 19(6):483–485, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800060107.

Hancock:2000:DRJ

[Han00o] Bill Hancock. DefCon recruiting JamFest. Computers & Security, 19(6):493–494, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480006017X.

Hancock:2000:DCG

[Han00p] Bill Hancock. Digital certificates get creative. Computers & Security, 19(6):480–487, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800060065.

Hancock:2000:DFP

[Han00q] Bill Hancock. DOD finds plans online — reserve unit assesses risk. Computers & Security, 19(4):297–298, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800040050.

Hancock:2000:CPF

[Han00r] Bill Hancock. E-commerce paranoia: Flaws in code. Computers & Security, 19(6):477–479, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800060041.

Hancock:2000:MPI

[Han00s] Bill Hancock. E-Mail privacy issues escalate. Computers & Security, 19(6):495, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800060193.

Hancock:2000:EPD

[Han00t] Bill Hancock. European Parliament doesn’t like anonymity online. Computers & Security, 19(4):297, April 1, 2000. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800040049.

Hancock:2000:EGO

[Han00u] Bill Hancock. Extortion going online. Computers & Security, 19(5):392–393, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800050100.

Hancock:2000:FSM

[Han00v] Bill Hancock. Feeling sorry for Microsoft? Computers & Security, 19(8):665–666, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800080032.

Hancock:2000:FSV

[Han00w] Bill Hancock. ‘files streams’ virus infects in an old and new way. Computers & Security, 19(7):573–574, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800070061.

Hancock:2000:FPV

[Han00x] Bill Hancock. First PDA virus hits the airwaves. Computers & Security, 19(7):583–584, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800070176.

Hancock:2000:SSC

[Han00y] Bill Hancock. From stones in the street to cyberspace — Israeli and Palestinians continue their virtual battles 24 × 7. Computers & Security, 19(8):664–665, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800080020.

Hancock:2000:Ea

[Han00z] Bill Hancock. From the Editor. Computers & Security, 19(1):2–5, January 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800863555.

Hancock:2000:Eb

[Han00-27] Bill Hancock. From the Editor. Computers & Security, 19(2):106–107, February 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800878190.

Hancock:2000:Ec

[Han00-28] Bill Hancock. From the Editor. Computers & Security, 19(3):200–201, March 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800886083.

Hancock:2000:Ed

[Han00-29] Bill Hancock. From the Editor. Computers & Security, 19(4):294–295, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800040013.

Hancock:2000:Ee

[Han00-30] Bill Hancock. From the Editor. Computers & Security, 19(5):380–381, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480005001X.

Hancock:2000:Ef

[Han00-31] Bill Hancock. From the Editor. Computers & Security, 19(6):474–475, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800060016.

Hancock:2000:Eg

[Han00-32] Bill Hancock. From the Editor. Computers & Security, 19(7):568–569, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800070012.

Hancock:2000:Eh

[Han00-33] Bill Hancock. From the Editor. Computers & Security, 19(8):662–663, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800080019.

Hancock:2000:GTA

[Han00-34] Bill Hancock. G8 thinks about cybercrime (it’s about time, too). Computers & Security, 19(5):405–407, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800050264.

Hancock:2000:GOR

[Han00-35] Bill Hancock. GeoCities ordered to report on information poster. Computers & Security, 19(6):495–496, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740480006020X.

Hancock:2000:GLH

[Han00-36] Bill Hancock. Getting the laws to help combat cybercrime (there’s a grand idea). Computers & Security, 19(8):669–670, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800080081.

Hancock:2000:HIN

[Han00-37] Bill Hancock. Hacker insurance now part of the business risk management kit. Computers & Security, 19 (7):574–575, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800070085.

Hancock:2000:HTM

[Han00-38] Bill Hancock. Hacker target: Mobile phones. Computers & Security, 19(6):494–495, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800060181.

Hancock:2000:HBF

[Han00-39] Bill Hancock. Hackers breach firewall-1. Computers & Security, 19(6):496–497, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800060223.

Hancock:2000:HSL

[Han00-40] Bill Hancock. Hackers still love to attack the Pentagon. Computers & Security, 19(6):488, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800060144.

Hancock:2000:HI

[Han00-41] Bill Hancock. Hacktivism increasing. Computers & Security, 19(8):673–674, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800080111.

Hancock:2000:HCS

[Han00-42] Bill Hancock. Health care security: a hard look at a growing problem. Computers & Security, 19(6):482, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800060077.

Hancock:2000:HPT

[Han00-43] Bill Hancock. Home PC’s targeted by hackers (duh!). Computers & Security, 19(5):387–388, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800050057.

Hancock:2000:HMC

[Han00-44] Bill Hancock. However, Microsoft could use some sympathy — Dutch hacker busts Microsoft Web site — again. Computers & Security, 19(8):666–667, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800080044.

Hancock:2000:ALA

[Han00-45] Bill Hancock. “I Agree” legal agreements get legislated. Computers & Security, 19(5):383–384, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800050033.

Hancock:2000:IEI

[Han00-46] Bill Hancock. Intel eliminates ID from new chips. Computers & Security, 19(4):296–297, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800040037.

Hancock:2000:IBG

[Han00-47] Bill Hancock. Internet biz group calls for cybercrime treaty delay. Computers & Security, 19(8):671, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800080093.

Hancock:2000:CHR

[Han00-48] Bill Hancock. Is a convicted hacker really reformed and should you hire them? Computers & Security, 19(6):491–493, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404800060168.

Hancock:2000:III

[Han00-49] Bill Hancock. Isn’t it interesting what you can buy at an auction site? TV’s, computers, drugs …. Computers & Security, 19(5):404–405, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800050252.

Hancock:2000:JCC

[Han00-50] Bill Hancock. Japan creates cyberconditions for IT national revolution. Computers & Security, 19(8):676–677, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404800080147.

Hancock:2000:JDC

[Han00-51] Bill Hancock. Justice Department conducting criminal probe in former CIA Director activities. Computers & Security, 19(4):302–303, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800040104.

Hancock:2000:LTN

[Han00-52] Bill Hancock. Laptop theft now targeted towards data and not necessarily hardware. Computers & Security, 19(8):671–673, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480008010X.

Hancock:2000:LCP

[Han00-53] Bill Hancock. Large child pornography ring busted in Texas. Computers & Security, 19(4):308–309, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800040141.

Hancock:2000:LIW

[Han00-54] Bill Hancock. … But the industry wants self-regulation. Computers & Security, 19(5):394–395, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800050136.

Hancock:2000:LDH

[Han00-55] Bill Hancock. Lotus domino holes up. Computers & Security, 19(6):497–498, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800060235.

Hancock:2000:MHN

[Han00-56] Bill Hancock. ‘Mafiaboy’ hit with 64 new charges. Computers & Security, 19(6):496, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800060211.

Hancock:2000:MPV

[Han00-57] Bill Hancock. Microsoft a popular virus target due to ubiquity. Computers & Security, 19(5):389–391, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800050082.

Hancock:2000:MRN

[Han00-58] Bill Hancock. Microsoft releases new IIS security tool. Computers & Security, 19(7):572, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800070048.

Hancock:2000:NHP

[Han00-59] Bill Hancock. NASA hacker pleads guilty. Computers & Security, 19(8):668–669, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800080068.

Hancock:2000:NNS

[Han00-60] Bill Hancock. A need for a network security czar. Computers & Security, 19(6):476–477, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800060028.

Hancock:2000:NSR

[Han00-61] Bill Hancock. Network solutions Re-evaluates security. Computers & Security, 19(5):398–400, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800050185.

Hancock:2000:NCU

[Han00-62] Bill Hancock. New classes of Unix/Linux attacks. Computers & Security, 19(7):570–571, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800070024.

Hancock:2000:NDS

[Han00-63] Bill Hancock. New denial of service attack on Internet. Computers & Security, 19(4):309–310, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800040153.

Hancock:2000:EWP

[Han00-64] Bill Hancock. Not everyone wants PKI — NSF opts for digital signature alternative. Computers & Security, 19(4):301–302, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404800040098.

Hancock:2000:OPR

[Han00-65] Bill Hancock. Office 2000 patch recommended by CERT. Computers & Security, 19(5):393, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800050112.

Hancock:2000:PVK

[Han00-66] Bill Hancock. PGP v5.0 keys potentially insecure. Computers & Security, 19(5):396–397, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480005015X.

Hancock:2000:PGC

[Han00-67] Bill Hancock. Philippine government creates incident response team. Computers & Security, 19(6):483, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800060090.

Hancock:2000:P

[Han00-68]  Bill Hancock. ‘Porngate’?  Computers & Security,

19(6):486–487, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800060120.

Hancock:2000:PGW

[Han00-69] Bill Hancock. Privacy group wants ‘web bugs’ disclosures. Computers & Security, 19(7):581, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800070152.

Hancock:2000:SUW

[Han00-70] Bill Hancock. Safeway UK’s Website shut down. Computers & Security, 19(6):483, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800060089.

Hancock:2000:SVa

[Han00-71] Bill Hancock. Security views. Computers & Security, 19(1):6–17, January 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800863567.

Hancock:2000:SVb

[Han00-72] Bill Hancock. Security views. Computers & Security, 19(2):108–118, February 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800878207.

Hancock:2000:SVc

[Han00-73] Bill Hancock. Security views. Computers & Security, 19(3):202–221, March 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800886095.

Hancock:2000:SVd

[Han00-74] Bill Hancock. Security views. Computers & Security, 19(5):382–383, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800050021.

Hancock:2000:SRP

[Han00-75] Bill Hancock. Self-regulation for privacy doesn’t work says FTC. Computers & Security, 19(5):393–394, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800050124.

Hancock:2000:SSB

[Han00-76] Bill Hancock. Site spoofing becomes more popular. Computers & Security, 19(7):581–583, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800070164.

Hancock:2000:SSI

[Han00-77] Bill Hancock. Software scam — 17 indicted. Computers & Security, 19(4):310–311, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800040177.

Hancock:2000:SRY

[Han00-78] Bill Hancock. Speedy response yields success at DOE. Computers & Security, 19(5):401, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800050203.

Hancock:2000:SHN

[Han00-79] Bill Hancock. Spying at home: a new pastime to detect online romance. Computers & Security, 19(6):488–491, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404800060156.

Hancock:2000:SCL

[Han00-80] Bill Hancock. Suspect charged in ‘love bug’ worm has charges dropped. Computers & Security, 19(5):398, July 1, 2000. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800050173.

Hancock:2000:SHA

[Han00-81] Bill Hancock. Suspected hackers arrested in Russian creditcard fraud. Computers & Security, 19(4):296, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800040025.

Hancock:2000:WHW

[Han00-82] Bill Hancock. This is what happens when you lose your computer with classified data on it …. Computers & Security, 19(4):300–301, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800040086.

Hancock:2000:TVD

[Han00-83] Bill Hancock. Trinity v3, a DDoS tool, hits the streets. Computers & Security, 19(7):574, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800070073.

Hancock:2000:TCS

[Han00-84] Bill Hancock. Truly certified: Security certifications update. Computers & Security, 19(6):479–480, October 1, 2000. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800060053.

Hancock:2000:UFT

[Han00-85] Bill Hancock. UK fuel taxes protestor accused of mass hacking. Computers & Security, 19(7):580, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800070139.

Hancock:2000:URB

[Han00-86] Bill Hancock. UK’s RIP bill creates official snooping. Computers & Security, 19(5):391–392, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800050094.

Hancock:2000:UEC

[Han00-87] Bill Hancock. US and Europe cybercrime agreement problems. Computers & Security, 19(4):306–307, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800040128.

Hancock:2000:UDD

[Han00-88] Bill Hancock. US department of defense prepares cybercrime database. Computers & Security, 19(8):674, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404800080123.

Hancock:2000:UDE

[Han00-89] Bill Hancock. US department of energy security criticized — again. Computers & Security, 19(4):298–300, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800040074.

Hancock:2000:UGB

[Han00-90] Bill Hancock. US Government board setting up security metrics. Computers & Security, 19(7):580, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800070140.

Hancock:2000:UGF

[Han00-91] Bill Hancock. US Government fighting for expanded wiretap laws. Computers & Security, 19(6):477, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740480006003X.

Hancock:2000:UGT

[Han00-92] Bill Hancock. The US Government tells its users to clean up their e-mail. Computers & Security, 19(5):385–387, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404800050045.

Hancock:2000:UPL

[Han00-93] Bill Hancock. US privacy law proposals cause concerns. Computers & Security, 19(8):677–678, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800080159.

Hancock:2000:USW

[Han00-94] Bill Hancock. US Senate wakes up to need for cybertools for law enforcers. Computers & Security, 19(5):395–396, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404800050148.

Hancock:2000:USC

[Han00-95] Bill Hancock. US supreme court confirms ISPs not liable in e-mail messages. Computers & Security, 19(4):310, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404800040165.

Hancock:2000:VTC

[Han00-96] Bill Hancock. Visa’s ten commandments for e-security online. Computers & Security, 19(6):485–486, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800060119.

Hancock:2000:WYS

[Han00-97] Bill Hancock. What you sell online in France could be restricted …. Computers & Security, 19(6):487, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404800060132.

Hancock:2000:WHSb

[Han00-98] Bill Hancock. White House Security Summit builds towards a B2B security standard. Computers & Security, 19(7):578–579, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800070115.

Hancock:2000:WHSa

[Han00-99] Bill Hancock. White House supporting open source code. Computers & Security, 19(7):577–578, November 1, 2000. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800070103.

Hancock:2000:WB

[Han00-100]  Bill Hancock. Who’s to blame?  Computers & Secu-

rity, 19(5):402–403, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800050227.

Hancock:2000:WBB

[Han00-101] Bill Hancock. Wireless big brother. Computers & Security, 19(8):667–668, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800080056.

Hancock:2000:WCN

[Han00-102] Bill Hancock. Wireless crazed? No security for you! Computers & Security, 19(7):575–577, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800070097.

Hancock:2000:WUB

[Han00-103] Bill Hancock. Wireless ubiquity for e-biz. Computers & Security, 19(5):403, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800050239.

Hancock:2000:YFI

[Han00-104] Bill Hancock. Yahoo faces identity crisis. Computers & Security, 19(5):389, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800050070.

Hancock:2001:CSO

[Han01a] Bill Hancock. The Chief security Officer’s top ten list for 2001. Computers & Security, 20(1):10–14, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404801010100.

Hancock:2001:CPL

[Han01b] Bill Hancock. China passes law to promote Internet security — supposedly. Computers & Security, 20(1):7, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404801010057.

Hancock:2001:MSJ

[Han01c] Bill Hancock. E-mail spammers jailed. Computers & Security, 20(1):9–10, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801010094.

Hancock:2001:TBH

[Han01d] Bill Hancock. E-tailers batten down the hatches. Computers & Security, 20(1):8, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801010070.

Hancock:2001:FSE

[Han01e] Bill Hancock. Fighting SPAM in Europe. Computers & Security, 20(1):18–19, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801010148.

Hancock:2001:Ea

[Han01f] Bill Hancock. From the Editor. Computers & Security, 20(1):2–3, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480101001X.

Hancock:2001:Eb

[Han01g] Bill Hancock. From the Editor. Computers & Security, 20(2):96–97, April 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801002012.

Hancock:2001:Ec

[Han01h] Bill Hancock. From the Editor. Computers & Security, 20(3):186–187, May 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801003017.

Hancock:2001:Ed

[Han01i] Bill Hancock. From the Editor. Computers & Security, 20(5):346–347, July 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801005016.

Hancock:2001:Ee

[Han01j] Bill Hancock. From the Editor. Computers & Security, 20(6):448–450, September 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801006010.

Hancock:2001:Ef

[Han01k] Bill Hancock. From the Editor. Computers & Security, 20(7):544–545, October 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801007015.

Hancock:2001:Eg

[Han01l] Bill Hancock. From the Editor. Computers & Security, 20(8):640–642, December 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480100801X.

Hancock:2001:HHV

[Han01m] Bill Hancock. Hand-held virii on the way. Computers & Security, 20(1):6–7, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801010045.

Hancock:2001:IWH

[Han01n] Bill Hancock. Information warfare highlighted as a concern by US Government. Computers & Security, 20(1):8–9, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404801010082.

Hancock:2001:PT

[Han01o] Bill Hancock. It’s a privacy thing in 2001. Computers & Security, 20(1):17–18, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801010136.

Hancock:2001:PR

[Han01p] Bill Hancock. A practical risk. Computers & Security, 20(2):98–126, April 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801002024.

Hancock:2001:PAG

[Han01q] Bill Hancock. Privacy advocates grow in Europe. Computers & Security, 20(1):19–20, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480101015X.

Hancock:2001:SHP

[Han01r] Bill Hancock. ‘safe harbor’ provision adopters few and far between. Computers & Security, 20(1):14–15, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404801010112.

Hancock:2001:SVa

[Han01s] Bill Hancock. Security views. Computers & Security, 20(3):188–201, May 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801003029.

Hancock:2001:SVb

[Han01t] Bill Hancock. Security views. Computers & Security, 20(4):278–294, July 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801004023.

Hancock:2001:SVc

[Han01u] Bill Hancock. Security views. Computers & Security, 20(5):348–363, July 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801005028.

Hancock:2001:SVd

[Han01v] Bill Hancock. Security views. Computers & Security, 20(7):546–560, October 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801007027.

Hancock:2001:SVe

[Han01w] Bill Hancock. Security views. Computers & Security, 20(8):643–656, December 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801008021.

Hancock:2001:SM

[Han01x] Bill Hancock. Smarter marines. Computers & Security, 20(1):7, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801010069.

Hancock:2001:VAW

[Han01y] Bill Hancock. Voodoo is alive and well on your computer. Computers & Security, 20(1):5–6, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801010033.

Hancock:2001:WGD

[Han01z] Bill Hancock. The wireless great divide — but getting smaller. Computers & Security, 20(1):15–17, January 31, 2001. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404801010124.

Hancock:2001:YSL

[Han01-27] Bill Hancock. Y2K and security — a low priority year. Computers & Security, 20(1):4–5, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801010021.

Hancock:2002:SCM

[Han02] Bill Hancock. Security crisis management — the basics. Computers & Security, 21(5):397–401, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802005035.

Harding:2003:SVP

[Har03] Andrew Harding. SSL virtual private networks. Computers & Security, 22(5):416–420, July 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480300508X.

Han:2003:DIR

[HC03] Sang-Jun Han and Sung-Bae Cho. Detecting intrusion with rule-based integration of multiple models. Computers & Security, 22(7):613–623, October 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803007119.

Hwang:2004:OML

[HC04a] Gwan-Hwan Hwang and Tao-Ku Chang. An operational model and language support for securing XML documents. Computers & Security, 23(6):498–529, September 2004. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804001026.

Hwang:2004:ILA

[HC04b] Shin-Jia Hwang and Ching-Chung Chan. Improvement on Li et al.’s generalization of proxy signature schemes. Computers & Security, 23(7):615–619, October 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804001713.

Hernandez-Castro:2006:SGG

[HCBLETRG06] Julio C. Hernandez-Castro, Ignacio Blasco-Lopez, Juan M. Estevez-Tapiador, and Arturo Ribagorda-Garnacho. Steganography in games: a general methodology and its application to the game of go. Computers & Security, 25(1):64–71, February 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404805002002.

Halkidis:2006:QAS

[HCS06] Spyros T. Halkidis, Alexander Chatzigeorgiou, and George Stephanides. A qualitative analysis of software security patterns. Computers & Security, 25(5):379–392, July 2006. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404806000526.

Hovav:2005:CMR

[HD05] Anat Hovav and John D’Arcy. Capital market reaction to defective IT products: the case of computer viruses. Computers & Security, 24(5):409–424, August 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805000374.

Hone:2002:ISP

[HE02] Karin Höne and J. H. P. Eloff. Information security policy — what do international information security standards say? Computers & Security, 21(5):402–409, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404802005047.

Heiser:2003:BCB

[Hei03] Jay G. Heiser. Beyond cryptography: Bruce Schneier’s beyond fear: thinking sensibly about security in an uncertain world. Computers & Security, 22(8):673–674, December 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404803000051.

Herzberg:2009:DBE

[Her09a] Amir Herzberg. DNS-based email sender authentication mechanisms: a critical review. Computers & Security, 28(8):731–742, November 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000492.

Herzberg:2009:WJC

[Her09b] Amir Herzberg. Why Johnny can’t surf (safely)? Attacks and defenses for web users. Computers & Security, 28(1–2):63–71, February/March 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000916.

Horng:2008:FID

[HFC^{+}08] Shi-Jinn Horng, Pingzhi Fan, Yao-Ping Chou, Yen-Cheng Chang, and Yi Pan. A feasible intrusion detector for recognizing IIS attacks based on neural networks. Computers & Security, 27(3–4):84–100, May/June 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000114.

Hansman:2005:TNC

[HH05] Simon Hansman and Ray Hunt. A taxonomy of network and computer attacks. Computers & Security, 24(1):31–43, February 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404804001804.

Hilley:2002:B

[Hil02a] Sarah Hilley. In brief. Computers & Security, 21(2):126–130, March 31, 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802002146.

Hilley:2002:PFS

[Hil02b] Sarah Hilley. Palladium, fraud, and surviving terrorism — compsec 2002: Preview of Compsec 2002, 30 Oct–1 Nov, Queen Elizabeth II Conference Centre, Westminster, London, UK. Computers & Security, 21(6):520–521, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404802010088.

Hinde:2000:DYK

[Hin00a] Stephen Hinde. Do you know your Organization’s Achilles heel? Computers & Security, 19(7):585–590, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404800070188.

Hinde:2000:FBO

[Hin00b] Stephen Hinde. Fireworks, beer and old halfpennies — the risks of assumption. Computers & Security, 19(6):499–504, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404800060247.

Hinde:2000:LWS

[Hin00c] Stephen Hinde. Life was simple then. Computers & Security, 19(3):222–229, March 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800886101.

Hinde:2000:LCA

[Hin00d] Stephen Hinde. Love conquers all? Computers & Security, 19(5):408–420, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800050276.

Hinde:2000:NMO

[Hin00e] Stephen Hinde. New Millennium, old failures. Computers & Security, 19(2):119–127, February 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800878219.

Hinde:2000:NNO

[Hin00f] Stephen Hinde. New names for old — a personal surf through compsec 2000. Computers & Security, 19(8):679–687, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404800080160.

Hinde:2000:SSS

[Hin00g] Stephen Hinde. Smurfing, swamping, spamming, spoofing, squatting, slandering, surfing, scamming and other mischiefs of the world wide Web. Computers & Security, 19(4):312–320, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404800040189.

Hinde:2001:PO

[Hin01a] Stephen Hinde. 2001: a privacy odyssey. Computers & Security, 20(1):21–27, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801010161.

Hinde:2001:POR

[Hin01b] Stephen Hinde. 2001: a privacy odyssey revisited. Computers & Security, 21(1):16–34, First Quarter 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802001037.

Hinde:2001:CPR

[Hin01c] Stephen Hinde. Cyberthreats: Perceptions, reality and protection. Computers & Security, 20(5):364–371, July 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740480100503X.

Hinde:2001:IYC

[Hin01d] Stephen Hinde. If you can meet with triumph and disaster and treat those two impostors just the same …. Computers & Security, 20(8):657–666, December 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801008033.

Hinde:2001:IPD

[Hin01e] Stephen Hinde. Incalculable potential for damage by cyberterrorism. Computers & Security, 20(7):568–572, October 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404801007040.

Hinde:2001:LL

[Hin01f] Stephen Hinde. Lessons learned. Computers & Security, 20(7):561–567, October 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801007039.

Hinde:2001:OTA

[Hin01g] Stephen Hinde. Omnia te adversum spectantia, nulla retorsum. Computers & Security, 20(6):468–474, September 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404801006034.

Hinde:2001:SP

[Hin01h] Stephen Hinde. The search for privacy. Computers & Security, 20(2):127–131, April 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801002036.

Hinde:2001:TTC

[Hin01i] Stephen Hinde. The times they are a-changin’. Computers & Security, 20(3):202–206, May 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801003030.

Hinde:2001:WL

[Hin01j] Stephen Hinde. The weakest link. Computers & Security, 20(4):295–301, July 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801004035.

Hinde:2002:BSD

[Hin02a] Stephen Hinde. The blue screen of death and other deadly bugs. Computers & Security, 21(6):491–496, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404802010039.

Hinde:2002:CCS

[Hin02b] Stephen Hinde. Compsec 2002: the complete security circle. Computers & Security, 21(8):689–693, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404802008039.

Hinde:2002:IWD

[Hin02c] Stephen Hinde. It was Déjà vu all over again. Computers & Security, 21(3):212–219, June 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802003036.

Hinde:2002:PP

[Hin02d] Stephen Hinde. The perils of privacy. Computers & Security, 21(5):424–432, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802005084.

Hinde:2002:SSS

[Hin02e] Stephen Hinde. Security surveys spring crop. Computers & Security, 21(4):310–321, August 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802004042.

Hinde:2002:SSC

[Hin02f] Stephen Hinde. Spam, scams, chains, hoaxes and other junk mail. Computers & Security, 21(7):592–606, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404802011045.

Hinde:2003:CAP

[Hin03a] Stephen Hinde. Careless about privacy. Computers & Security, 22(4):284–288, May 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803004036.

Hinde:2003:CSM

[Hin03b] Stephen Hinde. Computer security: Mapping the future. Computers & Security, 22(8):664–669, December 2003. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404803000038.

Hinde:2003:CTC

[Hin03c] Stephen Hinde. Cyber-terrorism in context. Computers & Security, 22(3):188–192, April 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803003031.

Hinde:2003:LCR

[Hin03d] Stephen Hinde. The law, cybercrime, risk assessment and cyber protection. Computers & Security, 22(2):90–95, February 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404803002037.

Hinde:2003:NDC

[Hin03e] Stephen Hinde. Nimbyism, dominoes and creaking infrastructure. Computers & Security, 22(7):570–576, October 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740480300703X.

Hinde:2003:PLC

[Hin03f] Stephen Hinde. Privacy legislation: a comparison of the US and European approaches. Computers & Security, 22(5):378– 387, July 2003. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404803005030.

Hinde:2003:SEN

[Hin03g] Stephen Hinde. Spam: the evolution of a nuisance. Computers & Security, 22(6):474–478, September 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803006035.

Hinde:2003:TCM

[Hin03h] Stephen Hinde. Time cost$ money. Computers & Security, 22(1):14–21, January 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803001032.

Hansen:2002:OSA

[HKP02] Marit Hansen, Kristian Köhntopp, and Andreas Pfitzmann. The open source approach — opportunities and limitations with respect to security and privacy. Computers & Security, 21(5):461–471, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802005163.

Hu:2007:DWD

[HLC07] Ming-Chiang Hu, Der-Chyuan Lou, and Ming-Chang Chang. Dual-wrapped digital watermarking scheme for image copyright protection. Computers & Security, 26(4):319–330, June 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404806002070.

Hwang:2005:IFT

[HLL05] Min-Shiang Hwang, Jung-Wen Lo, and Chia-Hsin Liu. Improvement on the flexible tree-based key management framework. Computers & Security, 24(6):500–504, September 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804002391.

Huang:2009:OSW

[HLTJ09] Y. L. Huang, P. H. Lu, J. D. Tygar, and A. D. Joseph. OSNP: Secure wireless authentication protocol using one-time key. Computers & Security, 28(8):803–815, November 2009. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809000558.

Hu:2004:TLI

[HM04] Ji Hu and Christoph Meinel. Tele-lab “IT-security” on CD: portable, reliable and safe IT security training. Computers & Security, 23(4):282–289, June 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804000975.

Holeman:2002:DSM

[HMDC02] S. Holeman, G. Manimaran, J. Davis, and A. Chakrabarti. Differentially secure multicasting and its implementation methods. Computers & Security, 21(8):736–749, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404802008143.

Hancke:2009:CST

[HMM09] G. P. Hancke, K. E. Mayes, and K. Markantonakis. Confidence in smart token proximity: Relay attacks revisited. Computers & Security, 28(7):615–627, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000595.

Hofmeyr:2004:IIS

[Hof04] S. Hofmeyr. The implications of immunology for secure systems design. Computers & Security, 23(6):453–455, September 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740480400166X.

Hofmeyr:2005:HIP

[Hof05] S. Hofmeyr. Host intrusion prevention: Part of the operating system or on top of the operating system? Computers & Security, 24(6):440–442, September 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805001069.

Horrocks:2001:STE

[Hor01] Ivan Horrocks. Security training: Education for an emerging profession? Computers & Security, 20(3):219–226, May 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404801003066.

Harn:2008:EIBa

[HR08a] Lein Harn and Jian Ren. Efficient identity-based RSA multisignatures. Computers & Security, 27(1–2):12–15, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404808000059.

Harn:2008:EIBb

[HR08b] Lein Harn and Jian Ren. Efficient identity-based RSA multisignatures. Computers & Security, 27(1–2):12–15, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404808000059.

Harn:2008:EIBc

[HR08c] Lein Harn and Jian Ren. Efficient identity-based RSA multisignatures. Computers & Security, 27(1–2):12–15, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404808000059.

Hoogenboom:2000:SRA

[HS00] Mark Hoogenboom and Patrick Steemers. Security for remote access and mobile applications. Computers & Security, 19(2):149–163, February 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800878256.

Herzog:2005:PJS

[HS05] Almut Herzog and Nahid Shahmehri. Performance of the Java security manager. Computers & Security, 24(3):192–207, May 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404804001932.

Huang:2000:GEP

[HSH00] Yu-Lun Huang, Shiuh-Pyng Shieh, and Fu-Shen Ho. A generic electronic payment model supporting multiple merchant transactions. Computers & Security, 19(5):453–465, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404800050318.

Huffmire:2008:EMPa

[HSKL08a] Ted Huffmire, Timothy Sherwood, Ryan Kastner, and Timothy Levin. Enforcing memory policy specifications in reconfigurable hardware. Computers & Security, 27(5–6):197–215, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000138.

Huffmire:2008:EMPb

[HSKL08b] Ted Huffmire, Timothy Sherwood, Ryan Kastner, and Timothy Levin. Enforcing memory policy specifications in reconfigurable hardware. Computers & Security, 27(5–6):197–215, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000138.

Huffmire:2008:EMPc

[HSKL08c] Ted Huffmire, Timothy Sherwood, Ryan Kastner, and Timothy Levin. Enforcing memory policy specifications in reconfigurable hardware. Computers & Security, 27(5–6):197–215, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000138.

Hunker:2002:PCB

[Hun02] Jeffrey Hunker. Policy challenges in building dependability in global infrastructures. Computers & Security, 21(8):705–711, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802008076.

Hsu:2003:CIT

[HW03] Chien-Lung Hsu and Tzong-Sun Wu. Cryptanalyses and improvements of two cryptographic key assignment schemes for dynamic access control in a user hierarchy. Computers & Security, 22(5):453–456, July 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803005145.

Hongxin:2009:REM

[HYJ^{+}09] Zhang Hongxin, Huang Yuewang, Wang Jianxin, Lu Yinghua, and Zhang Jinling. Recognition of electro-magnetic leakage information from computer radiation with SVM. Computers & Security, 28(1–2):72–76, February/March 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480800093X.

Zuo:2006:IIH

[hZxZtZ06] Zhi hong Zuo, Qing xin Zhu, and Ming tian Zhou. Infection, imitation and a hierarchy of computer viruses. Computers & Security, 25(6):469–473, September 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806000307.

Iheagwara:2004:EID

[Ihe04] Charles Iheagwara. The effect of intrusion detection management methods on the return on investment. Computers & Security, 23(3):213–228, May 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804000203.

Izadinia:2006:UIS

[IKE06] Vafa D. Izadinia, D. G. Kourie, and J. H. P. Eloff. Uncovering identities: a study into VPN tunnel fingerprinting. Computers & Security, 25(2):97–105, March 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805002014.

Igure:2006:SIS

[ILW06] Vinay M. Igure, Sean A. Laughter, and Ronald D. Williams. Security issues in SCADA networks. Computers & Security, 25(7):498–506, October 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806000514.

Jayeju-akinsiku:2002:TEC

[Ja02] Babatunde Jayeju-akinsiku. Technology and Electronic Communications Act 2000. Computers & Security, 21(7):624–628, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802011100.

Jeon:2004:FXE

[JCKL04] Jae-Myeong Jeon, Yon Dohn Chung, Myoung Ho Kim, and Yoon Joon Lee. Filtering XPath expressions for XML access control. Computers & Security, 23(7):591–605, October 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804001786.

Johnston:2003:SHC

[JEL03] J. Johnston, J. H. P. Eloff, and L. Labuschagne. Security and human computer interfaces. Computers & Security, 22(8):675–684, December 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803000063.

Jiang:2005:NMB

[JHgJ05] Jun Jiang, Chen He, and Ling ge Jiang. A novel mixbased location privacy mechanism in mobile IPv6. Computers & Security, 24(8):629–641, November 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805000957.

Jeong:2007:KAKa

[JL07a] Ik Rae Jeong and Dong Hoon Lee. Key agreement for key hypergraph. Computers & Security, 26(7–8):452–458, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404807000946.

Jeong:2007:KAKb

[JL07b] Ik Rae Jeong and Dong Hoon Lee. Key agreement for key hypergraph. Computers & Security, 26(7–8):452–458, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404807000946.

Jeong:2007:KAKc

[JL07c] Ik Rae Jeong and Dong Hoon Lee. Key agreement for key hypergraph. Computers & Security, 26(7–8):452–458, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404807000946.

Lee:2007:RKD

[jLC07] Hyoung joo Lee and Sungzoon Cho. Retraining a keystroke dynamics-based authenticator with impostor patterns. Computers & Security, 26(4):300–310, June 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806002057.

Jiang:2004:FAP

[JPL04] Rui Jiang, Li Pan, and Jian-Hua Li. Further analysis of password authentication schemes based on authentication tests. Computers & Security, 23(6):469–477, September 2004. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804001038.

Jiang:2005:IEA

[JPL05] Rui Jiang, Li Pan, and Jian-Hua Li. An improvement on efficient anonymous auction protocols. Computers & Security, 24(2):169–174, March 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804002408.

Janczewski:2002:DIS

[JS02] Lech Janczewski and Frank Xinli Shi. Development of information security baselines for healthcare information systems in new Zealand. Computers & Security, 21(2):172–192, March 31, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404802002122.

Johansson:2003:DCV

[JS03] Jesper M. Johansson and E. Eugene Schultz. Dealing with contextual vulnerabilities in code: distinguishing between solutions and pseudosolutions. Computers & Security, 22(2):152–159, February 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480300213X.

Jiang:2005:RTI

[JSD05] Wenbao Jiang, Hua Song, and Yiqi Dai. Real-time intrusion detection for high-speed networks. Computers & Security, 24(4):287–294, June 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804002354.

Jian:2007:MDB

[JST^{+}07] Zhou Jian, Haruhiko Shirai, Isamu Takahashi, Jousuke Kuroiwa, Tomohiro Odaka, and Hisakazu Ogura. Masquerade detection by boosting decision stumps using UNIX commands. Computers & Security, 26(4):311–318, June 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806002069.

Juang:2004:EPA

[Jua04] Wen-Shenq Juang. Efficient password authenticated key agreement using smart cards. Computers & Security, 23 (2):167–173, March 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404804000252.

Kamel:2009:RSW

[KA09] Ibrahim Kamel and Qutaiba Albluwi. A robust software watermarking for copyright protection. Computers & Security, 28(6):395–409, September 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480900008X.

Kamel:2009:SPI

[Kam09] Ibrahim Kamel. A schema for protecting the integrity of databases. Computers & Security, 28(7):698–709, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404809000297.

Kovacich:2000:IT

[KB00] Gerald L. Kovacich and William C. Boni. Internet targets. Computers & Security, 19(2):133–140, February 1, 2000. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800878232.

Kim:2005:EES

[KC05] Han-Sung Kim and Sung-Deok Cha. Empirical evaluation of SVM-based masquerade detection using UNIX commands. Computers & Security, 24(2):160–168, March 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804001944.

Kraemer:2009:HOF

[KCC09] Sara Kraemer, Pascale Carayon, and John Clem. Human and organizational factors in computer and information security: Pathways to vulnerabilities. Computers & Security, 28(7):509–520, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000467.

Kamara:2003:AVI

[KFS^{+}03] Seny Kamara, Sonia Fahmy, Eugene Schultz, Florian Kerschbaum, and Michael Frantzen. Analysis of vulnerabilities in Internet firewalls. Computers & Security, 22(3):214–232, April 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404803003109.

Kapsalis:2006:DCA

[KHKK06] Vassilis Kapsalis, Loukas Hadellis, Dimitris Karelis, and Stavros Koubias. A dynamic context-aware access control architecture for e-services. Computers & Security, 25(7): 507–521, October 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404806000770.

King:2004:AAS

[Kin04] Stuart King. Applying application security standards — a case study. Computers & Security, 23(1):17–21, February 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804000124.

Kovacich:2001:WIP

[KJ01] Gerald L. Kovacich and Andy Jones. What InfoSec professionals should know about information warfare tactics by terrorists. Computers & Security, 21(1):35–41, First Quarter 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404802001049.

Kovacich:2002:WIP

[KJ02] Gerald L. Kovacich and Andy Jones. What InfoSec professionals should know about information warfare tactics by terrorists. Computers & Security, 21(2):113–119, March 31, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404802002031.

Kjaerland:2006:TCC

[Kja06] Maria Kjaerland. A taxonomy and comparison of computer security incidents from the commercial and government sectors. Computers & Security, 25(7):522–538, October 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001234.

Kirda:2009:CSC

[KJKV09] Engin Kirda, Nenad Jovanovic, Christopher Kruegel, and Giovanni Vigna. Client-side cross-site scripting protection. Computers & Security, 28(7):592–604, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000479.

Kokolakis:2000:AIM

[KK00] S. A. Kokolakis and E. A. Kiountouzis. Achieving interoperability in a multiple-security-policies environment. Computers & Security, 19(3):267–281, March 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800886150.

Kenny:2002:ADR

[KK02] Steve Kenny and Larry Korba. Applying digital rights management systems to privacy rights management. Computers & Security, 21(7):648–664, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802011173.

Kruger:2006:PAI

[KK06] H. A. Kruger and W. D. Kearney. A prototype for assessing information security awareness. Computers & Security, 25(4):289–296, June 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806000563.

Ksiezopolski:2007:ASM

[KK07] Bogdan Ksiezopolski and Zbigniew Kotulski. Adaptable security mechanism for dynamic environments. Computers & Security, 26(3):246–255, May 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806001878.

Kruger:2008:CRIa

[KK08a] H. A. Kruger and W. D. Kearney. Consensus ranking — an ICT security awareness case study. Computers & Security, 27(7–8):254–259, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000448.

Kruger:2008:CRIb

[KK08b] H. A. Kruger and W. D. Kearney. Consensus ranking — an ICT security awareness case study. Computers & Security, 27(7–8):254–259, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000448.

Kruger:2008:CRIc

[KK08c] H. A. Kruger and W. D. Kearney. Consensus ranking — an ICT security awareness case study. Computers & Security, 27(7–8):254–259, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000448.

Khill:2001:MPF

[KKHR01] Insoo Khill, Jiseon Kim, Ingoo Han, and Jaecheol Ryou. Multi-party fair exchange protocol using ring architecture model. Computers & Security, 20(5):422–439, July 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404801005144.

Karyda:2005:ISS

[KKK05] Maria Karyda, Evangelos Kiountouzis, and Spyros Kokolakis. Information systems security policies: a contextual perspective. Computers & Security, 24(3):246–260, May 2005. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804002378.

Kim:2007:SDC

[KL07] Sangkyun Kim and Hong Joo Lee. A study on decision consolidation methods using analytic models for security systems. Computers & Security, 26(2):145–153, March 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001404.

Kao:2009:DRA

[KLL09] Kuo-Fong Kao, I.-En Liao, and Yueh-Chia Li. Detecting rogue access points using client-side bottleneck bandwidth analysis. Computers & Security, 28(3–4):144–152, May/June 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404808001132.

Knapp:2009:ISP

[KMMB09] Kenneth J. Knapp, R. Franklin Morris, Thomas E. Marshall, and Terry Anthony Byrd. Information security policy: an organizational-level process model. Computers & Security, 28(7):493–508, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000765.

Kwak:2006:SEK

[KMWD06] DongJin Kwak, SangJae Moon, Guilin Wang, and Robert H. Deng. A secure extension of the Kwak–Moon group signcryption scheme. Computers & Security, 25(6):435–444, September 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404806000915.

Kennedy:2000:KRF

[KMZ00] John Kennedy, Stephen M. Matyas, and Nevenko Zunic. Key recovery functional model. Computers & Security, 19(1):31–36, January 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404800863609.

Kelly:2003:RCD

[KN03] Christopher Kelly and Chris Nelms. Roadmap to checking data migration. Computers & Security, 22(6):506–510, September 2003. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404803006084.

Kondakci:2009:CCA

[Kon09] Suleyman Kondakci. A concise cost analysis of Internet malware. Computers & Security, 28(7):648–659, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809000133.

Kovacich:2000:NPI

[Kov00a] Gerald L. Kovacich. Netspionage — part III: the black zone, who uses netspionage, how and why. Computers & Security, 19(6):505–519, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800060259.

Kovacich:2000:NGTa

[Kov00b] Gerald L. Kovacich. Netspionage — the global threat to information, part i: What is it and why I should care? Computers & Security, 19(4):326–336, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800040207.

Kovacich:2000:NGTb

[Kov00c] Gerald L. Kovacich. Netspionage — the global threat to information, part II: Information collection in the Gray zone. Computers & Security, 19(5):421–427, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800050288.

Kovacich:2001:CIA

[Kov01a] Gerald L. Kovacich. The corporate information assurance officer (CIAO). Computers & Security, 20(4):302–307, July 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404801004047.

Kovacich:2001:PSC

[Kov01b] Gerald L. Kovacich. Protecting 21-st Century information — It’s time for a change. Computers & Security, 20(3):207–213, May 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404801003042.

Kang:2008:IKDa

[KPsH^{+}08a] Pilsung Kang, Sunghoon Park, Seong seob Hwang, Hyoung joo Lee, and Sungzoon Cho. Improvement of keystroke data quality through artificial rhythms and cues. Computers & Security, 27(1–2):3–11, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000023.

Kang:2008:IKDb

[KPsH^{+}08b] Pilsung Kang, Sunghoon Park, Seong seob Hwang, Hyoung joo Lee, and Sungzoon Cho. Improvement of keystroke data quality through artificial rhythms and cues. Computers & Security, 27(1–2):3–11, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000023.

Kang:2008:IKDc

[KPsH^{+}08c] Pilsung Kang, Sunghoon Park, Seong seob Hwang, Hyoung joo Lee, and Sungzoon Cho. Improvement of keystroke data quality through artificial rhythms and cues. Computers & Security, 27(1–2):3–11, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000023.

Karabacak:2005:IIS

[KS05] Bilge Karabacak and Ibrahim Sogukpinar. ISRAM: information security risk analysis method. Computers & Security, 24(2):147–159, March 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804001890.

Karabacak:2006:QMI

[KS06] Bilge Karabacak and Ibrahim Sogukpinar. A quantitative method for ISO 17799 gap analysis. Computers & Security, 25(6):413–419, September 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806000757.

Kritzinger:2008:ISMa

[KS08a] E. Kritzinger and E. Smith. Information security management: an information security retrieval and awareness model for industry. Computers & Security, 27(5–6):224–231, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404808000321.

Kritzinger:2008:ISMb

[KS08b] E. Kritzinger and E. Smith. Information security management: an information security retrieval and awareness model for industry. Computers & Security, 27(5–6):224–231, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404808000321.

Kritzinger:2008:ISMc

[KS08c] E. Kritzinger and E. Smith. Information security management: an information security retrieval and awareness model for industry. Computers & Security, 27(5–6):224–231, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404808000321.

Kumar:2006:PAH

[Kum06] Sanjeev Kumar. PING attack — how bad is it? Computers & Security, 25(5):332–337, July 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805001926.

Kumar:2009:SCU

[Kum09] K. Vimal Kumar. Securing communication using function extraction technology for malicious code behavior analysis. Computers & Security, 28(1–2):77–84, February/March 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404808000904.

Komninos:2006:LSD

[KVD06] Nikos Komninos, Dimitris Vergados, and Christos Douligeris. Layered security design for mobile ad hoc networks. Computers & Security, 25(2):121–130, March 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805001562.

Komninos:2007:ALS

[KVD07] Nikos Komninos, Dimitrios D. Vergados, and Christos Douligeris. Authentication in a layered security approach for mobile ad hoc networks. Computers & Security, 26(5):373– 380, August 2007. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404807000041.

Lo:2008:SEEa

[LBE08a] Johnny Li-Chang Lo, Judith Bishop, and J. H. P. Eloff. SMSSec: an end-to-end protocol for secure SMS. Computers & Security, 27(5–6):154–167, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000151.

Lo:2008:SEEb

[LBE08b] Johnny Li-Chang Lo, Judith Bishop, and J. H. P. Eloff. SMSSec: an end-to-end protocol for secure SMS. Computers & Security, 27(5–6):154–167, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000151.

Lo:2008:SEEc

[LBE08c] Johnny Li-Chang Lo, Judith Bishop, and J. H. P. Eloff. SMSSec: an end-to-end protocol for secure SMS. Computers & Security, 27(5–6):154–167, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000151.

Lu:2007:STP

[LC07] Rongxing Lu and Zhenfu Cao. Simple three-party key exchange protocol. Computers & Security, 26(1):94–97, February 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404806001246.

Lin:2004:FSM

[LCC04] Min-Hui Lin, Chin-Chen Chang, and Yan-Ren Chen. A fair and secure mobile agent environment based on blind signature and proxy host. Computers & Security, 23(3):199–212, May 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404804000288.

Lam:2003:SME

[LCGS03] Kwok-Yan Lam, Siu-Leung Chung, Ming Gu, and Jia-Guang Sun. Security middleware for enhancing interoperability of public key infrastructure. Computers & Security, 22(6):535–546, September 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803006151.

Lee:2006:RTA

[LCK^{+}06] Soojin Lee, Byungchun Chung, Heeyoul Kim, Yunho Lee, Chanil Park, and Hyunsoo Yoon. Real-time analysis of intrusion detection alerts via correlation. Computers & Security, 25(3):169–183, May 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404805001549.

Lee:2009:CWR

[LCLL09] Junsup Lee, Sungdeok Cha, Dongkun Lee, and Hyungkyu Lee. Classification of web robots: an empirical study based on over one billion requests. Computers & Security, 28(8):795–802, November 2009. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404809000546.

Leach:2003:SES

[Lea03a] Dr John Leach. Security engineering and security RoI. Computers & Security, 22(6):482–486, September 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803006059.

Leach:2003:IUS

[Lea03b] John Leach. Improving user security behaviour. Computers & Security, 22(8):685–692, December 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803000075.

Leach:2004:TEA

[Lea04] John Leach. TBSE — an engineering approach to the design of accurate and reliable security systems. Computers & Security, 23(1):22–28, February 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804000690.

LeVine:2005:TED

[LeV05] Richard LeVine. Technology evolution drives need for greater information technology security. Computers & Security, 24(5):359–361, August 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404805000982.

Lei:2007:CSA

[LFHT07] Jun Lei, Xiaoming Fu, Dieter Hogrefe, and Jianrong Tan. Comparative studies on authentication and key exchange methods for 802.11 wireless LAN. Computers & Security, 26(5):401–409, August 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807000053.

Lekkas:2004:CNL

[LG04] Dimitris Lekkas and Dimitris Gritzalis. Cumulative notarization for long-term preservation of digital signatures. Computers & Security, 23(5):413–424, July 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804001014.

Liang:2006:VDE

[LG06] Jinqian Liang and Xiaohong Guan. A virtual disk environment for providing file system recovery. Computers & Security, 25(8):589–599, November 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001209.

Li:2007:ALBa

[LG07a] Yang Li and Li Guo. An active learning based TCM-KNN algorithm for supervised network intrusion detection. Computers & Security, 26(7–8):459–467, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807001101.

Li:2007:ALBb

[LG07b] Yang Li and Li Guo. An active learning based TCM-KNN algorithm for supervised network intrusion detection. Computers & Security, 26(7–8):459–467, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807001101.

Li:2007:ALBc

[LG07c] Yang Li and Li Guo. An active learning based TCM-KNN algorithm for supervised network intrusion detection. Computers & Security, 26(7–8):459–467, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807001101.

Levi:2009:ULM

[LG09] Albert Levi and Can Berk Güder. Understanding the limitations of S/MIME digital signatures for e-mails: a GUI based approach. Computers & Security, 28(3–4):105–120, May/June 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404808000783.

Li:2007:MAS

[LGW07] Tao Li, Zhihong Guan, and Xianyong Wu. Modeling and analyzing the spread of active worms based on P2P systems. Computers & Security, 26(3):213–218, May 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001519.

Lin:2003:PAS

[LH03] Chun-Li Lin and Tzonelih Hwang. A password authentication scheme with secure password updating. Computers & Security, 22(1):68–72, January 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803001147.

Lee:2004:ETP

[LHL04] Tian-Fu Lee, Tzonelih Hwang, and Chun-Li Lin. Enhanced three-party encrypted key exchange without server public keys. Computers & Security, 23(7):571–577, October 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804001762.

Li:2004:ARI

[Li04] Ming Li. An approach to reliably identifying signs of DDOS flood attacks based on LRD traffic pattern recognition. Computers & Security, 23(7):549–558, October 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804001245.

Li:2006:CTA

[Li06] Ming Li. Change trend of averaged Hurst parameter of traffic under DDOS flood attacks. Computers & Security, 25(3):213– 220, May 2006. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404805001963.

Liaw:2004:SEV

[Lia04] Horng-Twu Liaw. A secure electronic voting protocol for general elections. Computers & Security, 23(2):107–119, March 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404804000276.

Lin:2001:HKA

[Lin01] Chu-Hsing Lin. Hierarchical key assignment without publickey cryptography. Computers & Security, 20(7):612–619, October 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404801007118.

Li:2007:FSB

[LK07] Jun Li and Paul Knickerbocker. Functional similarities between computer worms and biological pathogens. Computers & Security, 26(4):338–347, June 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806002094.

Lou:2002:SMS

[LL02] Der-Chyuan Lou and Jiang-Lung Liu. Steganographic method for secure communications. Computers & Security, 21(5):449–460, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802005151.

Lu:2009:DHE

[LL09] Hongwei Lu and Bailing Liu. DFANS: a highly efficient strategy for automated trust negotiation. Computers & Security, 28(7):557–565, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000169.

Lee:2002:ISL

[LLL02] Younghwa Lee, Jintae Lee, and Zoonky Lee. Integrating software lifecycle process standards with security engineering. Computers & Security, 21(4):345–355, August 1, 2002.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404802004133.

Lopez:2004:AAI

[LOP04] Javier Lopez, Rolf Oppliger, and Günther Pernul. Authentication and authorization infrastructures (AAIs): a comparative survey. Computers & Security, 23(7):578–590, October 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804001828.

Lacohee:2006:RRA

[LPF06] H. Lacohee, A. D. Phippen, and S. M. Furnell. Risk and restitution: Assessing how users establish online trust. Computers & Security, 25(7):486–493, October 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001489.

Lee:2005:DAS

[LS05] Fu-Yuan Lee and Shiuhpyng Shieh. Defending against spoofed DDoS attacks with path fingerprint. Computers & Security, 24(7):571–586, October 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805000465.

Liginlal:2009:HSH

[LSK09] Divakaran Liginlal, Inkook Sim, and Lara Khansa. How significant is human error as a cause of privacy breaches? An empirical study and a framework for error management. Computers & Security, 28(3–4):215–228, May/June 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808001181.

Li:2003:GPS

[LTH03] Li-Hua Li, Shiang-Feng Tzeng, and Min-Shiang Hwang. Generalization of proxy signature-based on discrete logarithms. Computers & Security, 22(3):245–255, April 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803003122.

Luoma:2006:CFE

[Luo06] Vicki Miller Luoma. Computer forensics and electronic discovery: the new management challenge. Computers & Security, 25(2):91–96, March 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806000186.

Liao:2002:UNN

[LV02] Yihua Liao and V. Rao Vemuri. Use of K-nearest neighbor classifier for intrusion detection. Computers & Security, 21(5):439–448, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480200514X.

Lin:2009:DMG

[LWL09] Song Lin, Biao Wang, and Zhoujun Li. Digital multisignature on the generalized conic curve over Z_{n}. Computers & Security, 28(1–2):100–104, February/March 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000801.

Li:2009:BLI

[LWT^{+}09] Yang Li, Jun-Li Wang, Zhi-Hong Tian, Tian-Bo Lu, and Chen Young. Building lightweight intrusion detection system using wrapper-based feature selection mechanisms. Computers & Security, 28(6):466–475, September 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000030.

Li:2009:TSD

[LZXW09] Qi Li, Xinwen Zhang, Mingwei Xu, and Jianping Wu. Towards secure dynamic collaborations with group-based RBAC model. Computers & Security, 28(5):260–275, July 2009. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404808001375.

Machanick:2005:DSA

[Mac05] Philip Machanick. A distributed systems approach to secure Internet mail. Computers & Security, 24(6):492–499, September 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404805000489.

Matai:2003:EIW

[Mat03] D. K. Matai. The economic impact of war with Iraq — asymmetric risks. Computers & Security, 22(2):119–123, February 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404803002074.

Moulton:2003:AIS

[MC03a] Rolf Moulton and Robert S. Coles. Applying information security governance. Computers & Security, 22(7):580–584, October 2003. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404803007053.

Moulton:2003:CEI

[MC03b] Rolf Moulton and Robert S. Coles. A contest to evaluate IT security services management. Computers & Security, 22(3):204–206, April 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803003067.

McGraw:2002:BWW

[McG02] Gary McGraw. On bricks and walls: Why building secure software is hard. Computers & Security, 21(3):229–238, June 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740480200305X.

McKenna:2002:EFN

[McK02a] Brian McKenna. Enterprise in focus at NetSec 2002. Computers & Security, 21(5):421–423, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802005072.

McKenna:2002:MSS

[McK02b] Brian McKenna. Managed security services — new economy relic or wave of the future? Computers & Security, 21(7):613–616, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802011070.

McKenna:2002:NP

[McK02c] Brian McKenna. Note from the publishers. Computers & Security, 21(7):591, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802011033.

McKenna:2003:UPP

[McK03a] Brian McKenna. UK police promise charter to guard good names. Computers & Security, 22(1):38–40, January 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404803001081.

McKenna:2003:WSS

[McK03b] Brian McKenna. Web services set to provoke new sthreats: Preview of Compsec 2003, 30 Oct-1 Nov, Queen Elizabeth II Conference Centre, Westminster, London, UK. Computers & Security, 22(6):515–516, September 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803006102.

McKenna:2004:EAE

[McK04] Brian McKenna. Erratum to “Attacks on the (enhanced) Yang–Shieh authentication” [Comput Secur 22(8) (2003) 725– 727]. Computers & Security, 23(1):85, February 2004. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804000082. See [CZ03].

Meyer:2002:ARA

[Mey02] Helen Meyer. Abstracts of recent articles and literature. Computers & Security, 21(3):246–250, June 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802003097.

Magklaras:2001:ITP

[MF01] G. B. Magklaras and S. M. Furnell. Insider threat prediction tool: Evaluating the probability of IT misuse. Computers & Security, 21(1):62–73, First Quarter 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802001098.

Magklaras:2005:PME

[MF05] G. B. Magklaras and S. M. Furnell. A preliminary model of end user sophistication for insider threat prediction in IT systems. Computers & Security, 24(5):371–380, August 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804002603.

Macia-Fernandez:2008:ELRa

[MFDVGT08a] Gabriel Maciá-Fernández, Jesús E. Díaz-Verdejo, and Pedro García-Teodoro. Evaluation of a low-rate DoS attack against application servers. Computers & Security, 27(7–8):335–354, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000394.

Macia-Fernandez:2008:ELRb

[MFDVGT08b] Gabriel Maciá-Fernández, Jesús E. Díaz-Verdejo, and Pedro García-Teodoro. Evaluation of a low-rate DoS attack against application servers. Computers & Security, 27(7–8):335–354, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000394.

Macia-Fernandez:2008:ELRc

[MFDVGT08c] Gabriel Maciá-Fernández, Jesús E. Díaz-Verdejo, and Pedro García-Teodoro. Evaluation of a low-rate DoS attack against application servers. Computers & Security, 27(7–8):335–354, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000394.

Munoz-Gea:2009:PDE

[MGMSML^{+}09] Juan Pedro Muñoz-Gea, Josemaria Malgosa-Sanahuja, Pilar Manzanares-Lopez, Juan Carlos Sanchez-Aarnoutse, and Joan Garcia-Haro. Proposal, design and evaluation of a mechanism to limit the length of anonymous overlay network paths. Computers & Security, 28(8):772–784, November 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000509.

McKenna:2002:DRIa

[MH02a] Brian McKenna and Sarah Hilley. Digest of recent IT security press coverage. Computers & Security, 21(4):338–342, August 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740480200408X.

McKenna:2002:DRIb

[MH02b] Brian McKenna and Sarah Hilley. Digest of recent IT security press coverage. Computers & Security, 21(5):433–436, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802005096.

McKenna:2002:B

[MH02c] Brian McKenna and Sarah Hilley. In brief. Computers & Security, 21(5):433–436, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802005102.

Moitra:2004:EIN

[MK04] Soumyo D. Moitra and Suresh L. Konda. An empirical investigation of network attacks on computer systems. Computers & Security, 23(1):43–51, February 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804000677.

Mangipudi:2006:SIK

[MK06] Kumar Mangipudi and Rajendra Katti. A secure identification and key agreement protocol with user anonymity (SIKA).

Computers & Security, 25(6):420–425, September 2006. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404806000769.

Mourad:2008:AOA

[MLD08] Azzam Mourad, Marc-André Laverdière, and Mourad Debbabi. An aspect-oriented approach for the systematic security hardening of code. Computers & Security, 27(3–4):101–114, May/June 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000126.

Maclean:2000:OIG

[MMZ00] Abby Maclean, Stephen M. Matyas, and Nevenko Zunic. Organization implementation guidelines for recovery of encrypted information. Computers & Security, 19(1):69–81, January 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404800863658.

Mouratidis:2005:SSM

[Mou05] Haralambos Mouratidis. Safety and security in multiagent systems: Report on the 2nd SASEMAS workshop (SASE-MAS’05). Computers & Security, 24(8):614–617, November 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404805001422.

Mazhelis:2007:FBB

[MP07] Oleksiy Mazhelis and Seppo Puuronen. A framework for behavior-based detection of user substitution in a mobile context. Computers & Security, 26(2):154–176, March 2007. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404806001416.

Masri:2008:ABAa

[MP08a] Wes Masri and Andy Podgurski. Application-based anomaly intrusion detection with dynamic information flow analysis. Computers & Security, 27(5–6):176–187, October 2008. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404808000369.

Masri:2008:ABAb

[MP08b] Wes Masri and Andy Podgurski. Application-based anomaly intrusion detection with dynamic information flow analysis. Computers & Security, 27(5–6):176–187, October 2008. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404808000369.

Masri:2008:ABAc

[MP08c] Wes Masri and Andy Podgurski. Application-based anomaly intrusion detection with dynamic information flow analysis. Computers & Security, 27(5–6):176–187, October 2008. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404808000369.

Marmol:2009:STS

[MP09] Félix Gómez Mármol and Gregorio Martínez Pérez. Security threats scenarios in trust and reputation models for distributed systems. Computers & Security, 28(7):545–556, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404809000534.

Mitropoulos:2006:IHR

[MPD06] Sarandis Mitropoulos, Dimitrios Patsos, and Christos Douligeris. On incident handling and response: a state-of-the-art approach. Computers & Security, 25(5):351–370, July 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404805001574.

Moon:2004:SRM

[MPPB04] Chang-Joo Moon, Dae-Ha Park, Soung-Jin Park, and Doo-Kwon Baik. Symmetric RBAC model that takes the separation of duty and role hierarchies into consideration. Computers & Security, 23(2):126–136, March 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804000185.

Matyas:2000:BSI

[MS00] Stephen M. Matyas and Jeff Stapleton. A biometric standard for information management and security. Computers & Security, 19(5):428–441, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480005029X.

Mitropoulos:2009:SLS

[MS09] Dimitris Mitropoulos and Diomidis Spinellis. SDriver: Location-specific signatures prevent SQL injection attacks. Computers & Security, 28(3–4):121–129, May/June 2009. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404808000928.

Mourad:2009:NAO

[MSLD09] Azzam Mourad, Andrei Soeanu, Marc-André Laverdière, and Mourad Debbabi. New aspect-oriented constructs for security hardening concerns. Computers & Security, 28(6):341–358, September 2009. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740480900011X.

Marchesini:2005:KSI

[MSZ05] John Marchesini, S. W. Smith, and Meiyuan Zhao. Keyjacking: the surprising insecurity of client-side SSL. Computers & Security, 24(2):109–123, March 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480400183X.

Mannan:2009:RTF

[MvO09] Mohammad Mannan and P. C. van Oorschot. Reducing threats from flawed security APIs: the banking PIN case. Computers & Security, 28(6):410–420, September 2009. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809000078.

Markham:2000:KRH

[MW00] Tom Markham and Chuck Williams. Key recovery header for IPSEC. Computers & Security, 19(1):86–90, January 1, 2000.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800863671.

Ma:2006:ADS

[MW06] Mingchao Ma and Steve Woodhead. Authentication delegation for subscription-based remote network services. Computers & Security, 25(5):371–378, July 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806000575.

Matyas:2000:AKR

[MZ00] Stephen M. Matyas and Nevenko Zunic. Additional key recovery functions. Computers & Security, 19(1):37–40, January 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404800863610.

Nabi:2005:SBA

[Nab05] Faisal Nabi. Secure business application logic for e-commerce systems. Computers & Security, 24(3):208–217, May 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804002123.

Ng:2006:SBB

[NCS06] Wee Hock Desmond Ng, Haitham Cruickshank, and Zhili Sun. Scalable balanced batch rekeying for secure group communication. Computers & Security, 25(4):265–273, June 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404806000496.

Nichol:2000:ARAa

[Nic00a] Sandy Nichol. Abstracts of recent articles and literature. Computers & Security, 19(2):164–172, February 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800878268.

Nichol:2000:ARAb

[Nic00b] Sandy Nichol. Abstracts of recent articles and literature. Computers & Security, 19(3):257–262, March 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800886149.

Nichol:2000:ARAc

[Nic00c] Sandy Nichol. Abstracts of recent articles and literature. Computers & Security, 19(4):359–363, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800040244.

Nichol:2000:ARAd

[Nic00d] Sandy Nichol. Abstracts of recent articles and literature. Computers & Security, 19(5):443–448, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800050306.

Nikishin:2002:WLA

[Nik02a] Andy Nikishin. I-Worm.Lentin (aka Yaha). Computers & Security, 21(6):497–502, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802010040.

Nikishin:2002:KHW

[Nik02b] Andy Nikishin. The klez.h worm dissected. Computers & Security, 21(4):322–327, August 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802004054.

Nikishin:2002:MAA

[Nik02c] Andy Nikishin. Much ado about nothing: Win32.perrun. Computers & Security, 21(5):410–412, October 1, 2002. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404802005059.

Nosworthy:2000:IIS

[Nos00a] Julie D. Nosworthy. Implementing information security in the 21-st Century — do you have the balancing factors? Computers & Security, 19(4):337–347, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800040219.

Nosworthy:2000:PRA

[Nos00b] Julie D. Nosworthy. A practical risk analysis approach: Managing BCM risk. Computers & Security, 19(7):596– 614, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404800070206.

Novak:2007:IRA

[Nov07] Christopher J. Novak. Investigative response: After the breach. Computers & Security, 26(2):183–185, March 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740480600143X.

Noore:2004:EBI

[NTH04] Afzel Noore, Nikhil Tungala, and Max M. Houck. Embedding biometric identifiers in 2D barcodes for improved security. Computers & Security, 23(8):679–686, December 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804002366.

OGorman:2005:QDP

[OBB05] Lawrence O’Gorman, Amit Bagga, and Jon Bentley. Querydirected passwords. Computers & Security, 24(7):546–560, October 2005. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404805001008.

Omar:2009:RFD

[OCB09] Mawloud Omar, Yacine Challal, and Abdelmadjid Bouabdallah. Reliable and fully distributed trust model for mobile ad hoc networks. Computers & Security, 28(3–4):199–214, May/ June 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740480800117X.

Onut:2007:SNV

[OG07] Iosif-Viorel Onut and Ali A. Ghorbani. SVision: a novel visual network-anomaly identification technique. Computers & Security, 26(3):201–212, May 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001507.

Oppliger:2008:STS

[OHB08] Rolf Oppliger, Ralf Hauser, and David Basin. SSL/ TLS session-aware user authentication revisited. Computers & Security, 27(3–4):64–70, May/June 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000102.

Obiedkov:2009:BAC

[OKE09] Sergei Obiedkov, Derrick G. Kourie, and J. H. P. Eloff. Building access control models with attribute exploration. Computers & Security, 28(1–2):2–7, February/March 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000497.

Oshri:2007:ISN

[OKH07] Ilan Oshri, Julia Kotlarsky, and Corey Hirsch. Information security in networkable Windows-based operating system devices: Challenges and solutions. Computers & Security, 26(2):177–182, March 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806001490.

Oh:2003:AID

[OL03] Sang Hyun Oh and Won Suk Lee. An anomaly intrusion detection method by clustering normal user behavior. Computers & Security, 22(7):596–612, October 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803007107.

Oliveira:2007:PPC

[OZ07] Stanley R. M. Oliveira and Osmar R. Zaïane. A privacypreserving clustering approach toward secure and effective data analysis for business collaboration. Computers & Security, 26(1):81–93, February 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001222.

Palmer:2000:ARAa

[Pal00a] Chloë Palmer. Abstracts of recent articles and literature. Computers & Security, 19(6):540–546, October 1, 2000. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800060284.

Palmer:2000:ARAb

[Pal00b] Chloë Palmer. Abstracts of recent articles and literature. Computers & Security, 19(7):621–626, November 1, 2000. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740480007022X.

Palmer:2000:ARAc

[Pal00c] Chloë Palmer. Abstracts of recent articles and literature. Computers & Security, 19(8):720–725, December 1, 2000. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800080214.

Palmer:2001:ARAb

[Pal01a] Chloë Palmer. Abstracts of recent articles and literature. Computers & Security, 20(2):155–160, April 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801002061.

Palmer:2001:ARAc

[Pal01b] Chloë Palmer. Abstracts of recent articles and literature. Computers & Security, 20(3):239–243, May 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480100308X.

Palmer:2001:ARAd

[Pal01c] Chloë Palmer. Abstracts of recent articles and literature. Computers & Security, 20(4):322–326, July 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801004084.

Palmer:2001:ARAe

[Pal01d] Chloë Palmer. Abstracts of recent articles and literature. Computers & Security, 20(5):400–403, July 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801005107.

Palmer:2001:ARAf

[Pal01e] Chloë Palmer. Abstracts of recent articles and literature. Computers & Security, 20(6):509, September 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801006095.

Palmer:2008:CEAa

[Pal08a] Anthony J. Palmer. Criteria to evaluate automated personal identification mechanisms. Computers & Security, 27(7–8):260–284, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480800045X.

Palmer:2008:CEAb

[Pal08b] Anthony J. Palmer. Criteria to evaluate automated personal identification mechanisms. Computers & Security, 27(7–8):260–284, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480800045X.

Palmer:2008:CEAc

[Pal08c] Anthony J. Palmer. Criteria to evaluate automated personal identification mechanisms. Computers & Security, 27(7–8):260–284, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480800045X.

Piqueira:2008:DMCa

[PdVGA08a] Jose R. C. Piqueira, Adolfo A. de Vasconcelos, Carlos E. C. J. Gabriel, and Vanessa O. Araujo. Dynamic models for computer viruses. Computers & Security, 27(7–8):355–359, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000412.

Piqueira:2008:DMCb

[PdVGA08b] Jose R. C. Piqueira, Adolfo A. de Vasconcelos, Carlos E. C. J. Gabriel, and Vanessa O. Araujo. Dynamic models for computer viruses. Computers & Security, 27(7–8):355–359, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000412.

Piqueira:2008:DMCc

[PdVGA08c] Jose R. C. Piqueira, Adolfo A. de Vasconcelos, Carlos E. C. J. Gabriel, and Vanessa O. Araujo. Dynamic models for computer viruses. Computers & Security, 27(7–8):355–359, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000412.

Paul:2006:CJN

[PE06] Nathanael Paul and David Evans. Comparing Java and .NET security: Lessons learned and missed. Computers & Security, 25(5):338–350, July 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806000290.

Padayachee:2009:AUC

[PE09] Keshnee Padayachee and J. H. P. Eloff. Adapting usage control as a deterrent to address the inadequacies of access controls. Computers & Security, 28(7):536–544, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809000157.

Palmieri:2008:CLSa

[PF08a] Francesco Palmieri and Ugo Fiore. Containing large-scale worm spreading in the Internet by cooperative distribution of traffic filtering policies. Computers & Security, 27(1–2):48– 62, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404807001241.

Palmieri:2008:CLSb

[PF08b] Francesco Palmieri and Ugo Fiore. Containing large-scale worm spreading in the Internet by cooperative distribution of traffic filtering policies. Computers & Security, 27(1–2):48– 62, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404807001241.

Palmieri:2008:CLSc

[PF08c] Francesco Palmieri and Ugo Fiore. Containing large-scale worm spreading in the Internet by cooperative distribution of traffic filtering policies. Computers & Security, 27(1–2):48– 62, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404807001241.

Palmieri:2009:PTE

[PF09] Francesco Palmieri and Ugo Fiore. Providing true end-to-end security in converged voice over IP infrastructures. Computers & Security, 28(6):433–449, September 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404809000054.

Phan:2006:CTP

[Pha06] Raphael C.-W. Phan. Cryptanalysis of two passwordbased authentication schemes using smart cards. Computers & Security, 25(1):52–54, February 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805001938.

Philippsohn:2001:MLI

[Phi01a] Steven Philippsohn. Money laundering on the Internet. Computers & Security, 20(6):485–490, September 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480100606X.

Philippsohn:2001:TCO

[Phi01b] Steven Philippsohn. Trends in cybercrime — an overview of current financial crimes on the Internet. Computers & Security, 20(1):53–69, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801010215.

Piper:2003:RCS

[Pip03] Fred Piper. Reseacrh in cryptography and security mechanisms. Computers & Security, 22(1):22–25, January 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404803001044.

Post:2007:EIS

[PK07] Gerald V. Post and Albert Kagan. Evaluating information security tradeoffs: Restricting access can interfere with user tasks. Computers & Security, 26(3):229–237, May 2007. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404806001532.

Palmer:2001:ARAa

[PM01] Chloë Palmer and Helen Meyer. Abstracts of recent articles and literature. Computers & Security, 20(1):70–74, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404801010227.

Pudar:2009:PPM

[PML09] Srdjan Pudar, G. Manimaran, and Chen-Ching Liu. PENET: a practical method and tool for integrated modeling of security attacks and countermeasures. Computers & Security, 28(8):754–771, November 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000522.

Peyravian:2000:GRK

[PMRZ00a] Mohammad Peyravian, Stephen M. Matyas, Allen Roginsky, and Nevenko Zunic. Generation of RSA keys that are guaranteed to be unique for each user. Computers & Security, 19(3):282–288, March 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800886162.

Peyravian:2000:MBB

[PMRZ00b] Mohammad Peyravian, Stephen M. Matyas, Allen Roginsky, and Nevenko Zunic. Multiparty biometric-based authentication. Computers & Security, 19(4):369–374, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800040256.

Peyravian:2000:TCB

[PMRZ00c] Mohammad Peyravian, Stephen M. Matyas, Allen Roginsky, and Nevenko Zunic. Ticket and challenge-based protocols for timestamping. Computers & Security, 19(6):551–558, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404800060296.

Peyravian:2001:AMU

[PMRZ01a] Mohammad Peyravian, Stephen M. Matyas, Allen Roginsky, and Nevenko Zunic. Alternative method for unique RSA primes generation. Computers & Security, 20(2):173–179, April 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404801002103.

Peyravian:2001:MTE

[PMRZ01b] Mohammad Peyravian, Stephen M. Matyas, Allen Roginsky, and Nevenko Zunic. Methods for timestamping electronic documents using certificates and user-specified times. Computers & Security, 20(3):255–262, May 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801003133.

Pollard:2005:TFC

[Pol05] C. Pollard. Telecom fraud: the cost of doing nothing just went up. Computers & Security, 24(6):437–439, September 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404805001082.

Pounder:2001:CEC

[Pou01a] Chris Pounder. The council of Europe cyber-crime convention. Computers & Security, 20(5):380–383, July 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801005065.

Pounder:2001:CCB

[Pou01b] Chris Pounder. Cyber crime: the backdrop to the council of Europe convention. Computers & Security, 20(4):311–315, July 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404801004060.

Pounder:2001:EUP

[Pou01c] Chris Pounder. The European union proposal for a policy towards network and information security. Computers & Security, 20(7):573–576, October 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801007052.

Pounder:2002:ATL

[Pou02a] Chris Pounder. Anti-terrorism legislation: the impact on the processing of data. Computers & Security, 21(3):240–245, June 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802003073.

Pounder:2002:ECO

[Pou02b] Chris Pounder. The emergence of a comprehensive obligation towards computer security. Computers & Security, 21(4):328–332, August 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802004066.

Pounder:2002:SPU

[Pou02c] Chris Pounder. Security policy update. Computers & Security, 21(7):620–623, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802011094.

Pounder:2002:UNS

[Pou02d] Chris Pounder. The US’s national strategy for homeland security. Computers & Security, 21(6):503–505, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404802010052.

Pounder:2003:SUS

[Pou03a] Chris Pounder. Security with unfortunate side effects. Computers & Security, 22(2):115–118, February 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803002062.

Pounder:2003:GAI

[Pou03b] Dr Chris Pounder. Governments act to improve security. Computers & Security, 22(3):207–211, April 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803003079.

Pounder:2003:TWL

[Pou03c] Dr Chris Pounder. A tangled Web of libel lies? Computers & Security, 22(1):34–37, January 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480300107X.

Pond:2000:WAC

[PPBH00] Rachael Pond, John Podd, Julie Bunnell, and Ron Henderson. Word association computer passwords: the effect of formulation techniques on recall and guessing rates. Computers & Security, 19(7):645–656, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800070231.

Peyravian:2003:MPU

[PRZ03] Mohammad Peyravian, Allen Roginsky, and Nevenko Zunic. Methods for preventing unauthorized software distribution. Computers & Security, 22(4):316–321, May 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803004115.

Peyravian:2004:NPM

[PRZ04] Mohammad Peyravian, Allen Roginsky, and Nevenko Zunic. Non-PKI methods for public key distribution. Computers & Security, 23(2):97–103, March 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804000318.

Purser:2001:SGT

[Pur01] Steve Purser. A simple graphical tool for modelling trust. Computers & Security, 20(6):479–484, September 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404801006058.

Purser:2002:WAC

[Pur02] Steve Purser. Why access control is difficult. Computers & Security, 21(4):303–309, August 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802004030.

Purser:2004:IRS

[Pur04] Steve A. Purser. Improving the ROI of the security management process. Computers & Security, 23(7):542–546, October 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404804002329.

Posthumus:2004:FGI

[PvS04] Shaun Posthumus and Rossouw von Solms. A framework for the governance of information security. Computers & Security, 23(8):638–646, December 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804002639.

Phan:2006:SCI

[PW06] Raphael C.-W. Phan and David Wagner. Security considerations for incremental hash functions based on pair block chaining. Computers & Security, 25(2):131–136, March 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404805002063.

Peyravian:2000:MPP

[PZ00] Mohammad Peyravian and Nevenko Zunic. Methods for protecting password transmission. Computers & Security, 19(5):466–469, July 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480005032X.

Qing:2005:STI

[QW05] Sihan Qing and Weiping Wen. A survey and trends on Internet worms. Computers & Security, 24(4):334–346, June 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804002585.

Rathmell:2001:PCI

[Rat01] Andrew Rathmell. Protecting critical information infrastructures. Computers & Security, 20(1):43–52, January 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404801010203.

Reid:2001:ERA

[RF01] Randall C. Reid and Stephen A. Floyd. Extending the risk analysis model to include market-insurance. Computers & Security, 20(4):331–339, July 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801004114.

Rodwell:2007:NIBa

[RFR07a] P. M. Rodwell, S. M. Furnell, and P. L. Reynolds. A non-intrusive biometric authentication mechanism utilising physiological characteristics of the human head. Computers & Security, 26(7–8):468–478, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807001095.

Rodwell:2007:NIBb

[RFR07b] P. M. Rodwell, S. M. Furnell, and P. L. Reynolds. A non-intrusive biometric authentication mechanism utilising physiological characteristics of the human head. Computers & Security, 26(7–8):468–478, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807001095.

Rodwell:2007:NIBc

[RFR07c] P. M. Rodwell, S. M. Furnell, and P. L. Reynolds. A non-intrusive biometric authentication mechanism utilising physiological characteristics of the human head. Computers & Security, 26(7–8):468–478, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807001095.

Ray:2006:TLB

[RK06] Indrakshi Ray and Mahendra Kumar. Towards a locationbased mandatory access control model. Computers & Security, 25(1):36–44, February 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480500101X.

Rhee:2009:SEI

[RKR09] Hyeun-Suk Rhee, Cheongtag Kim, and Young U. Ryu. Self-efficacy in information security: Its influence on end users’ information security practice behavior. Computers & Security, 28(8):816–826, November 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480900056X.

Rantos:2005:MKR

[RM05] Konstantinos Rantos and Chris J. Mitchell. Matching key recovery mechanisms to business requirements. Computers & Security, 24(3):232–245, May 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804002317.

Rezgui:2008:ISAa

[RM08a] Yacine Rezgui and Adam Marks. Information security awareness in higher education: an exploratory study. Computers & Security, 27(7–8):241–253, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000485.

Rezgui:2008:ISAb

[RM08b] Yacine Rezgui and Adam Marks. Information security awareness in higher education: an exploratory study. Computers & Security, 27(7–8):241–253, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000485.

Rezgui:2008:ISAc

[RM08c] Yacine Rezgui and Adam Marks. Information security awareness in higher education: an exploratory study. Computers & Security, 27(7–8):241–253, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000485.

Ruighaver:2007:OSC

[RMC07] A. B. Ruighaver, S. B. Maynard, and S. Chang. Organisational security culture: Extending the end-user perspective. Computers & Security, 26(1):56–62, February 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480600157X.

Roberts:2006:RFI

[Rob06] C. M. Roberts. Radio frequency identification (RFID). Computers & Security, 25(1):18–26, February 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480500204X.

Roberts:2007:BAV

[Rob07] Chris Roberts. Biometric attack vectors and defences. Computers & Security, 26(1):14–25, February 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480600215X.

Rogers:2003:RCP

[Rog03] Marc Rogers. The role of criminal profiling in the computer forensics process. Computers & Security, 22(4):292–298, May 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740480300405X.

Ryan:2006:EBI

[RR06] Julie J. C. H. Ryan and Daniel J. Ryan. Expected benefits of information security investments. Computers & Security, 25(8):579–588, November 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001192.

Reinke:2003:ASC

[RS03] John Reinke and Hossein Saiedian. The availability of source code in relation to timely response to security vulnerabilities. Computers & Security, 22(8):707–724, December 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404803000117.

Rogers:2004:FCF

[RS04] Marcus K. Rogers and Kate Seigfried. The future of computer forensics: a needs analysis survey. Computers & Security, 23(1):12–16, February 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804000100.

Rezmierski:2002:USS

[RSS02] Virginia E. Rezmierski, Marshall R. Seese, and Nathaniel St. Clair. University systems security logging: who is doing it and how far can they go? Computers & Security, 21(6):557–564, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802010155.

Son:2009:FFR

[SAF09] Joon Son and Jim Alves-Foss. A formal framework for real-time information flow analysis. Computers & Security, 28(6):421–432, September 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000066.

Samet:2009:DIH

[Sam09] Refik Samet. Design and implementation of highly reliable dual-computer systems. Computers & Security, 28(7): 710–722, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404809000418.

Shaikh:2009:SAU

[SBS09] Siraj A. Shaikh, Vicky J. Bush, and Steve A. Schneider. Specifying authentication using signal events in CSP. Computers & Security, 28(5):310–324, July 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000953.

Shen:2002:NKM

[SC02] Victor R. L. Shen and Tzer-Shyong Chen. A novel key management scheme based on discrete logarithms and polynomial interpolations. Computers & Security, 21(2):164–171, March 31, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404802002110.

Shao:2006:TTS

[SC06] Jun Shao and Zhenfu Cao. A traceable threshold signature scheme with multiple signing policies. Computers & Security, 25(3):201–206, May 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480500194X.

Schneier:2001:MSM

[Sch01a] Bruce Schneier. Managed security monitoring:network security for the 21st century. Computers & Security, 20 (6):491–503, September 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801006071.

Schultz:2001:EC

[Sch01b] E. Eugene Schultz. From the editor-in-chief. Computers & Security, 21(1):2–4, First Quarter 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802001013.

Schultz:2001:SV

[Sch01c] Eugene Schultz. Security views. Computers & Security, 21(1):5–15, First Quarter 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802001025.

Schultz:2002:FUP

[Sch02a] E. Eugene Schultz. A framework for understanding and predicting insider attacks. Computers & Security, 21(6):526–531, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740480201009X.

Schultz:2002:ECa

[Sch02b] E. Eugene Schultz. From the Editor-in-chief. Computers & Security, 21(2):98–100, March 31, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802002018.

Schultz:2002:ECb

[Sch02c] E. Eugene Schultz. From the editor-in-chief. Computers & Security, 21(3):198–200, June 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802003012.

Schultz:2002:GBC

[Sch02d] E. Eugene Schultz. The gap between cryptography and information security. Computers & Security, 21(8):674–676, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404802008015.

Schultz:2002:SSL

[Sch02e] E. Eugene Schultz. The sorry state of law enforcement. Computers & Security, 21(4):290–292, August 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802004017.

Schultz:2002:TSH

[Sch02f] E. Eugene Schultz. Taking a stand on hackers. Computers & Security, 21(5):382–384, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802005011.

Schultz:2002:UGB

[Sch02g] E. Eugene Schultz. The US Government- bigger and better information security? Computers & Security, 21(7):578–580, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740480201101X.

Schultz:2002:SVa

[Sch02h] Eugene Schultz. Security views. Computers & Security, 21(2):101–112, March 31, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480200202X.

Schultz:2002:SVb

[Sch02i] Eugene Schultz. Security views. Computers & Security, 21(3):201–211, June 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802003024.

Schultz:2002:SVc

[Sch02j] Eugene Schultz. Security views. Computers & Security, 21(4):293–302, August 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802004029.

Schultz:2002:SVd

[Sch02k] Eugene Schultz. Security views. Computers & Security, 21(5):385–396, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802005023.

Schultz:2002:SVe

[Sch02l] Eugene Schultz. Security views. Computers & Security, 21(6):481–490, October 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802010027.

Schultz:2002:SVf

[Sch02m] Eugene Schultz. Security views. Computers & Security, 21(7):581–590, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802011021.

Schultz:2002:SVg

[Sch02n] Eugene Schultz. Security views. Computers & Security, 21(8):677–688, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802008027.

Schultz:2003:ISM

[Sch03a] E. Eugene Schultz. Information security and the media. Computers & Security, 22(8):652–653, December 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803000014.

Schultz:2003:ISW

[Sch03b] E. Eugene Schultz. Internet security: what’s in the future? Computers & Security, 22(2):78–79, February 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803002013.

Schultz:2003:PBS

[Sch03c] E. Eugene Schultz. Pandora’s Box: spyware, adware, autoexecution, and NGSCB. Computers & Security, 22(5):366–367, July 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404803005017.

Schultz:2003:PP

[Sch03d] E. Eugene Schultz. Patching pandemonium. Computers & Security, 22(7):556–558, October 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803007016.

Schultz:2003:WCM

[Sch03e] E. Eugene Schultz. Why can’t Microsoft stay out of the InfoSec headlines? Computers & Security, 22(4):270–272, May 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404803004012.

Schultz:2003:AHW

[Sch03f] Eugene Schultz. Attackers hit Web hosting servers. Computers & Security, 22(4):273–283, May 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803004024.

Schultz:2003:SVa

[Sch03g] Eugene Schultz. Security views. Computers & Security, 22(1):4–13, January 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803001020.

Schultz:2003:SVb

[Sch03h] Eugene Schultz. Security views. Computers & Security, 22(2):80–89, February 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803002025.

Schultz:2003:SVc

[Sch03i] Eugene Schultz. Security views. Computers & Security, 22(3):176–187, April 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480300302X.

Schultz:2003:SVd

[Sch03j] Eugene Schultz. Security views. Computers & Security, 22(5):368–377, July 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803005029.

Schultz:2003:SVe

[Sch03k] Eugene Schultz. Security views. Computers & Security, 22(8):654–663, December 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803000026.

Schultz:2003:SVO

[Sch03l] Eugene Schultz. Security views: Online piracy battle heats up in US. Computers & Security, 22(6):464–473, September 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404803006023.

Schultz:2004:IRT

[Sch04a] Dr Eugene Schultz. Incident response teams need to change. Computers & Security, 23(2):87–88, March 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480400029X.

Schultz:2004:SVb

[Sch04b] E. Schultz. Security views. Computers & Security, 23(4):267– 274, June 2004. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740480400121X.

Schultz:2004:SVc

[Sch04c] E. Schultz. Security views. Computers & Security, 23(5):355– 361, July 2004. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404804001531.

Schultz:2004:COT

[Sch04d] E. Eugene Schultz. The case for one-time credentials. Computers & Security, 23(6):441–442, September 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804001853.

Schultz:2004:GBC

[Sch04e] E. Eugene Schultz. The gap between cryptography and information security: has it narrowed? Computers & Security, 23(7):531–532, October 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804002287.

Schultz:2004:SOH

[Sch04f] E. Eugene Schultz. Sarbanes–Oxley — a huge boon to information security in the US. Computers & Security, 23(5):353– 354, July 2004. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740480400152X.

Schultz:2004:IP

[Sch04g] Eugene Schultz. Intrusion prevention. Computers & Security, 23(4):265–266, June 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804001208.

Schultz:2004:UGR

[Sch04h] Eugene Schultz. Is the U.S. Government really getting serious about information security? Computers & Security, 23(8):621–622, December 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804002664.

Schultz:2004:STA

[Sch04i] Eugene Schultz. Security training and awareness-fitting a square peg in a round hole. Computers & Security, 23(1):1–2, February 2004. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404804000094.

Schultz:2004:SVa

[Sch04j] Eugene Schultz. Security views. Computers & Security, 23(3):181–190, May 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804000999.

Schultz:2004:SVd

[Sch04k] Eugene Schultz. Security views. Computers & Security, 23(7):533–541, October 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804002305.

Schultz:2004:WVW

[Sch04l] Eugene Schultz. Worms and viruses: are we losing control? Computers & Security, 23(3):179–180, May 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804000987.

Schultz:2005:LCL

[Sch05a] Dr Eugene Schultz. Lycos crosses the line. Computers & Security, 24(2):89–90, March 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804002810.

Schultz:2005:PIC

[Sch05b] Dr Eugene Schultz. Personal information compromises: It is time for the U.S. Government to wake up. Computers & Security, 24(4):261–262, June 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805000519.

Schultz:2005:SEG

[Sch05c] Dr Eugene Schultz. Search engines: a growing contributor to security risk. Computers & Security, 24(2):87–88, March 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404805000167.

Schultz:2005:SVa

[Sch05d] E. Schultz. Security views. Computers & Security, 24(5):349– 358, August 2005. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404805000970.

Schultz:2005:SVb

[Sch05e] E. Schultz. Security views. Computers & Security, 24(6): 427–436, September 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404805001094.

Schultz:2005:SVc

[Sch05f] E. Schultz. Security views. Computers & Security, 24(8): 589–598, November 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404805001744.

Schultz:2005:ADR

[Sch05g] E. Eugene Schultz. Aligning disaster recovery and security incident response. Computers & Security, 24(7):505–506, October 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404805001513.

Schultz:2005:HFS

[Sch05h] Eugene Schultz. The human factor in security. Computers & Security, 24(6):425–426, September 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805001045.

Schultz:2005:ICW

[Sch05i] Eugene Schultz. Infosec certification: Which way do we turn from here? Computers & Security, 24(8):587–588, November 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404805001756.

Schultz:2005:NIP

[Sch05j] Eugene Schultz. Non-infosec professionals in infosec? Computers & Security, 24(5):347–348, August 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805000969.

Schultz:2005:SDM

[Sch05k] Eugene Schultz. Security dilemmas with Microsoft’s Internet explorer. Computers & Security, 24(3):175–176, May 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404805000428.

Schultz:2006:AUP

[Sch06a] E. Eugene Schultz. About “unofficial patches”. Computers & Security, 25(2):79–80, March 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806000174.

Schultz:2006:CWI

[Sch06b] E. Eugene Schultz. The changing winds of information security. Computers & Security, 25(5):315–316, July 2006. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404806000940.

Schultz:2006:DBD

[Sch06c] E. Eugene Schultz. Dilemmas and boundaries of digital rights management. Computers & Security, 25(1):1–2, February 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404805002026.

Schultz:2006:ICD

[Sch06d] E. Eugene Schultz. Issues concerning the distribution of vulnerability information. Computers & Security, 25(7):475–476, October 2006. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404806001453.

Schultz:2006:MBH

[Sch06e] E. Eugene Schultz. Microsoft is back in the hot seat. Computers & Security, 25(6):393–394, September 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001131.

Schultz:2006:PFI

[Sch06f] E. Eugene Schultz. Predicting the future of InfoSec. Computers & Security, 25(8):553–554, November 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806001830.

Schultz:2006:RIS

[Sch06g] E. Eugene Schultz. Representing information security fairly and accurately. Computers & Security, 25(4):237, June 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404806000630.

Schultz:2006:SSO

[Sch06h] E. Eugene Schultz. Special systems: Overlooked sources of security risk? Computers & Security, 25(3):155, May 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404806000538.

Schryen:2007:IPE

[Sch07a] Guido Schryen. The impact that placing email addresses on the Internet has on the receipt of spam: an empirical analysis. Computers & Security, 26(5):361–372, August 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807000028.

Schultz:2007:MCN

[Sch07b] E. Eugene Schultz. Mobile computing: the next Pandora’s Box. Computers & Security, 26(3):187, May 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807000399.

Schultz:2007:SAS

[Sch07c] E. Eugene Schultz. Struggles in the academic side of infosec. Computers & Security, 26(4):267–268, June 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807000545.

Schultz:2007:WIC

[Sch07d] E. Eugene Schultz. What infosec changes are likely to result from the recent US election? Computers & Security, 26(1):1– 2, February 2007. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404806002148.

Schultz:2007:WVM

[Sch07e] E. Eugene Schultz. Windows vista: Microsoft’s brave new world. Computers & Security, 26(2):99, March 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807000247.

Smith:2002:PAI

[SE02] E. Smith and J. H. P. Eloff. A prototype for assessing information technology risks in health care. Computers & Security, 21(3):266–284, June 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802003139.

Seba:2006:FFT

[Seb06] H. Seba. FTKM: a fault-tolerant key management protocol for multicast communications. Computers & Security, 25(6):426–434, September 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806000976.

Song:2004:RSP

[SF04] Yougang Song and Brett D. Fleisch. Rico: a security proxy for mobile code. Computers & Security, 23(4):338–351, June 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404804000173.

Sadoddin:2009:IFS

[SG09] Reza Sadoddin and Ali A. Ghorbani. An incremental frequent structure mining framework for real-time alert correlation. Computers & Security, 28(3–4):153–173, May/June 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404808001144.

Hwang:2009:KDB

[sHCP09] Seong seob Hwang, Sungzoon Cho, and Sunghoon Park. Keystroke dynamics-based authentication for mobile devices. Computers & Security, 28(1–2):85–93, February/March 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404808000965.

Sherwood:2000:OE

[She00] John Sherwood. Opening up the enterprise. Computers & Security, 19(8):710–719, December 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800080202.

Shoriak:2000:STP

[Sho00] Timothy G. Shoriak. SSL/TLS protocol enablement for key recovery. Computers & Security, 19(1):100–104, January 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404800863695.

Shahriari:2007:VTG

[SJ07] Hamid Reza Shahriari and Rasool Jalili. Vulnerability Take Grant (VTG): an efficient approach to analyze network vulnerabilities. Computers & Security, 26(5):349–360, August 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404807000375.

Sklavos:2003:DDR

[SK03] N. Sklavos and O. Koufopavlou. Data dependent rotations, a trustworthy approach for future encryption systems/ciphers: low cost and high performance. Computers & Security, 22(7):585–588, October 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803007065.

Sutherland:2006:EER

[SKBM06] Iain Sutherland, George E. Kalb, Andrew Blyth, and Gaius Mulley. An empirical examination of the reverse engineering process for binary files. Computers & Security, 25(3):221– 228, May 2006. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740480500177X.

Stathopoulos:2008:SLMa

[SKM08a] Vassilios Stathopoulos, Panayiotis Kotzanikolaou, and Emmanouil Magkos. Secure log management for privacy assurance in electronic communications. Computers & Security, 27(7–8):298–308, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000400.

Stathopoulos:2008:SLMb

[SKM08b] Vassilios Stathopoulos, Panayiotis Kotzanikolaou, and Emmanouil Magkos. Secure log management for privacy assurance in electronic communications. Computers & Security, 27(7–8):298–308, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000400.

Stathopoulos:2008:SLMc

[SKM08c] Vassilios Stathopoulos, Panayiotis Kotzanikolaou, and Emmanouil Magkos. Secure log management for privacy assurance in electronic communications. Computers & Security, 27(7–8):298–308, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000400.

Stergiou:2004:AAF

[SLG04] T. Stergiou, M. S. Leeson, and R. J. Green. An alternative architectural framework to the OSI security model. Computers & Security, 23(2):137–153, March 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480400015X.

Shen:2003:SET

[SLH03] Jau-Ji Shen, Chih-Wei Lin, and Min-Shiang Hwang. Security enhancement for the timestamp-based password authentication scheme using smart cards. Computers & Security, 22(7):591–595, October 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803007090.

Sampigethaya:2006:FTC

[SP06] Krishna Sampigethaya and Radha Poovendran. A framework and taxonomy for comparison of electronic voting schemes. Computers & Security, 25(2):137–153, March 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805001914.

Schultz:2001:USA

[SPLS01] E. Eugene Schultz, Robert W. Proctor, Mei-Ching Lien, and Gavriel Salvendy. Usability and security an appraisal of usability issues in information security methods. Computers & Security, 20(7):620–634, October 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480100712X.

Sharma:2007:IDUa

[SPP07a] Alok Sharma, Arun K. Pujari, and Kuldip K. Paliwal. Intrusion detection using text processing techniques with a kernel based similarity measure. Computers & Security, 26(7–8):488–495, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807001113.

Sharma:2007:IDUb

[SPP07b] Alok Sharma, Arun K. Pujari, and Kuldip K. Paliwal. Intrusion detection using text processing techniques with a kernel based similarity measure. Computers & Security, 26(7–8):488–495, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807001113.

Sharma:2007:IDUc

[SPP07c] Alok Sharma, Arun K. Pujari, and Kuldip K. Paliwal. Intrusion detection using text processing techniques with a kernel based similarity measure. Computers & Security, 26(7–8):488–495, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807001113.

Sharma:2007:TIS

[SS07] Sushil K. Sharma and Joshua Sefchek. Teaching information systems security courses: a hands-on approach. Computers & Security, 26(4):290–299, June 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806002045.

Sadok:2009:RRI

[SSF^{+}09] Djamel H. Sadok, Eduardo Souto, Eduardo Feitosa, Judith Kelner, and Lars Westberg. RIP — a robust IP access architecture. Computers & Security, 28(6):359–380, September 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404809000108.

Stanton:2005:AEU

[SSMJ05] Jeffrey M. Stanton, Kathryn R. Stam, Paul Mastrangelo, and Jeffrey Jolton. Analysis of end user security behaviors. Computers & Security, 24(2):124–133, March 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804001841.

Stewart:2004:RPD

[Ste04] Andrew Stewart. On risk: perception and direction. Computers & Security, 23(5):362–370, July 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804001233.

Stoupa:2007:CSC

[SV07] K. Stoupa and A. Vakali. Clustering subjects in a credential-based access control framework. Computers & Security, 26(2):120–129, March 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806001258.

Smith:2000:CIR

[SVW00] Michael Smith, Paul Van Oorschot, and Michael Willett. Cryptographic information recovery using key recovery. Computers & Security, 19(1):21–27, January 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800863580.

Shieh:2006:ERM

[SW06] Wen-Gong Shieh and Jian-Min Wang. Efficient remote mutual authentication and key agreement. Computers & Security, 25(1):72–77, February 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404805001719.

Szczepanski:2004:BRN

[SWA^{+}04] J. Szczepanski, E. Wajnryb, J. M. Amigó, Maria V. Sanchez-Vives, and M. Slater. Biometric random number generators.

Computers & Security, 23(1):77–84, February 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804000641.

Su:2009:RTN

[SYL09] Ming-Yang Su, Gwo-Jong Yu, and Chun-Yuen Lin. A realtime network intrusion detection system for large-scale attacks based on an incremental mining approach. Computers & Security, 28(5):301–309, July 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480800134X.

Tseng:2005:SIM

[TC05] H. Chris Tseng and B. Jack Culpepper. Sinkhole intrusion in mobile ad hoc networks: the problem and some detection indicators. Computers & Security, 24(7):561–570, October 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404805001021.

Trompeter:2001:FIS

[TE01] C. M. Trompeter and J. H. P. Eloff. A framework for the implementation of socio-ethical controls in information security. Computers & Security, 20(5):384–391, July 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801005077.

Tomur:2006:ATS

[TE06] Emrah Tomur and Y. M. Erten. Application of temporal and spatial role based access control in 802.11 wireless networks. Computers & Security, 25(6):452–458, September 2006. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404806000927.

Thimbleby:2003:RE

[Thi03] Harold Thimbleby. The reduced Enigma. Computers & Security, 22(7):624–642, October 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803007120.

Tzeng:2004:INT

[THY04] Shiang-Feng Tzeng, Min-Shiang Hwang, and Cheng-Ying Yang. An improvement of nonrepudiable threshold proxy signature scheme with known signers. Computers & Security, 23(2):174–178, March 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804000239.

Theoharidou:2005:ITI

[TKKK05] Marianthi Theoharidou, Spyros Kokolakis, Maria Karyda, and Evangelos Kiountouzis. The insider threat to information systems and the effectiveness of ISO17799. Computers & Security, 24(6):472–484, September 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805000684.

Torrubia:2000:ISM

[TM00] Andres Torrubia and Francisco J. Mora. Information security in multiprocessor systems based on the x86 architecture. Computers & Security, 19(6):559–563, October 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800060302.

Torkestani:2009:RRF

[TM09] Javad Akbari Torkestani and Mohammad Reza Meybodi. RAID-RMS: a fault tolerant stripped mirroring RAID architecture for distributed systems. Computers & Security, 28(1–2):40–46, February/March 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000795.

Torrubia:2001:CRC

[TMM01] Andres Torrubia, Francisco J. Mora, and Luis Marti. Cryptography regulations for e-commerce and digital rights management. Computers & Security, 20(8):724–738, December 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404801008148.

Teoh:2004:PCK

[TNG04] Andrew B. J. Teoh, David C. L. Ngo, and Alwyn Goh. Personalised cryptographic key generation based on FaceHashing. Computers & Security, 23(7):606–614, October 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804001701.

Treek:2003:IFI

[Trè03] Denis Trèek. An integral framework for information systems security management. Computers & Security, 22(4):337–360, May 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404803004139.

Tsiakis:2005:EAI

[TS05a] Theodosios Tsiakis and George Stephanides. The economic approach of information security. Computers & Security, 24(2):105–108, March 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404805000209.

Tsiakis:2005:CST

[TS05b] Theodosios Tsiakis and George Sthephanides. The concept of security and trust in electronic payments. Computers & Security, 24(1):10–15, February 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804002792.

Tsai:2008:EMS

[Tsa08] Jia-Lun Tsai. Efficient multi-server authentication scheme based on one-way hash function without verification table. Computers & Security, 27(3–4):115–121, May/June 2008. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404808000084.

Tseng:2007:RCG

[Tse07] Yuh-Min Tseng. A resource-constrained group key agreement protocol for imbalanced wireless networks. Computers & Security, 26(4):331–337, June 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404806002082.

Thomson:2005:ISO

[TvS05] Kerry-Lynn Thomson and Rossouw von Solms. Information security obedience: a definition. Computers & Security, 24(1):69–75, February 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804002627.

Tang:2009:UBA

[TXL09] Yong Tang, Bin Xiao, and Xicheng Lu. Using a bioinformatics approach to generate accurate exploit-based signatures for polymorphic worms. Computers & Security, 28(8):827–842, November 2009. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404809000583.

Ungureanu:2004:FSC

[Ung04] Victoria Ungureanu. Formal support for certificate management policies. Computers & Security, 23(4):300–311, June 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404804000719.

Vanstone:2003:NGS

[Van03] S. A. Vanstone. Next generation security for wireless: elliptic curve cryptography. Computers & Security, 22(5):412–415, July 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404803005078.

Viswanathan:2001:HKE

[VBD01] Kapali Viswanathan, Colin Boyd, and Ed Dawson. Hybrid key escrow: a new paradigm. Computers & Security, 21(1):77–92, First Quarter 2001. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802001116.

vanderHaar:2003:MDI

[vdHvS03] Helen van der Haar and Rossouw von Solms. A model for deriving information security control attribute profiles. Computers & Security, 22(3):233–244, April 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803003110.

Venter:2002:VCI

[VE02] H. S. Venter and J. H. P. Eloff. Vulnerabilities categories for intrusion detection systems. Computers & Security, 21(7):617–619, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802011082.

Venter:2003:TIS

[VE03] H. S. Venter and J. H. P. Eloff. A taxonomy for information security technologies. Computers & Security, 22(4):299–307, May 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404803004061.

Venter:2004:VFC

[VE04] H. S. Venter and J. H. P. Eloff. Vulnerability forecasting — a conceptual model. Computers & Security, 23(6):489–497, September 2004. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404804001750.

Venter:2008:SVC

[VEL08] H. S. Venter, J. H. P. Eloff, and Y. L. Li. Standardising vulnerability categories. Computers & Security, 27(3–4):71–83, May/June 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000096.

Villarroel:2005:SIS

[VFMP05] Rodolfo Villarroel, Eduardo Fernández-Medina, and Mario Piattini. Secure information systems development — a survey and comparison. Computers & Security, 24(4):308–321, June 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740480400241X.

vanKrugten:2000:BSJ

[vKH00] Petra van Krugten and Mark Hoogenboom. B2C security — be just secure enough. Computers & Security, 19(4):348–356, April 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404800040220.

vonSolms:2000:IST

[vS00] Basie von Solms. Information security — the third wave? Computers & Security, 19(7):615–620, November 1, 2000. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404800070218.

vonSolms:2001:CGI

[vS01a] Basie von Solms. Corporate governance and information security. Computers & Security, 20(3):215–218, May 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404801003054.

vonSolms:2001:ISM

[vS01b] Basie von Solms. Information security — a multidimensional discipline. Computers & Security, 20(6):504–508, September 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404801006083.

vonSolms:2005:ISGa

[vS05a] Basie von Solms. Information security governance: COBIT or ISO 17799 or both? Computers & Security, 24(2):99–104, March 2005. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404805000210.

vonSolms:2005:ISGb

[vS05b] S. H. (Basie) von Solms. Information security governance — compliance management vs operational management. Computers & Security, 24(6):443–447, September 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805001057.

vonSolms:2006:ISF

[vS06] Basie von Solms. Information security — the fourth wave. Computers & Security, 25(3):165–168, May 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480600054X.

vonSolms:2004:SWN

[vSM04] Basie von Solms and Emil Marais. From secure wired networks to secure wireless networks — what are the extra risks? Computers & Security, 23(8):633–637, December 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804002330.

Vila:2007:SMA

[VSR07] X. Vila, A. Schuster, and A. Riera. Security for a multi-agent system based on JADE. Computers & Security, 26(5):391– 400, August 2007. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404806002100.

vonSolms:2001:IIS

[vSvS01] Basie von Solms and Rossouw von Solms. Incremental information security certification. Computers & Security, 20(4):308–310, July 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801004059.

vonSolms:2004:DSI

[vSvS04a] Basie von Solms and Rossouw von Solms. The 10 deadly sins of information security management. Computers & Security, 23(5):371–376, July 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804001221.

vonSolms:2004:PC

[vSvS04b] Rossouw von Solms and Basie von Solms. From policies to culture. Computers & Security, 23(4):275–279, June 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804000331.

vonSolms:2005:ISL

[vSvS05] Basie von Solms and Rossouw von Solms. From information security to … business security? Computers & Security, 24(4):271–273, June 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404805000544.

vonSolms:2006:ISGa

[vSvS06a] Rossouw von Solms and S. H. (Basie) von Solms. Information security governance: a model based on the direct-control cycle. Computers & Security, 25(6):408–412, September 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404806001167.

vonSolms:2006:ISGb

[vSvS06b] Rossouw von Solms and S. H. (Basie) von Solms. Information security governance: Due care. Computers & Security, 25(7):494–497, October 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806001441.

Vroom:2004:TIS

[VvS04] Cheryl Vroom and Rossouw von Solms. Towards information security behavioural compliance. Computers & Security, 23(3):191–198, May 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480400032X.

Wang:2005:MLR

[Wan05] Yun Wang. A multinomial logistic regression modeling approach for anomaly intrusion detection. Computers & Security, 24(8):662–674, November 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404805000751.

Wu:2003:UFR

[WC03] Shyi-Tsong Wu and Bin-Chang Chieu. A user friendly remote authentication scheme with smart cards. Computers & Security, 22(6):547–550, September 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803006163.

Wu:2005:NDI

[WC05] Hsien-Chu Wu and Chin-Chen Chang. A novel digital image watermarking scheme based on the vector quantization technique. Computers & Security, 24(6):460–471, September 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404805000672.

Wang:2007:UHS

[WC07] Chih-Chien Wang and Sheng-Yi Chen. Using header session messages to anti-spamming. Computers & Security, 26(5):381–390, August 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404807000065.

Weir:2009:UPS

[WDCJ09] Catherine S. Weir, Gary Douglas, Martin Carruthers, and Mervyn Jack. User perceptions of security, convenience and usability for ebanking authentication tokens. Computers & Security, 28(1–2):47–62, February/March 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000941.

Webb:2000:CMH

[Web00] Steve Webb. Crimes and misdemeanours: How to protect corporate information in the Internet age. Computers & Security, 19(2):128–132, February 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800878220.

Wang:2006:PPB

[WGZY06] Wei Wang, Xiaohong Guan, Xiangliang Zhang, and Liwei Yang. Profiling program behavior for anomaly intrusion detection based on the transition and frequency property of computer audit data. Computers & Security, 25(7):539–550, October 2006. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404806000903.

Wu:2004:EUI

[WH04] Tzong-Sun Wu and Chien-Lung Hsu. Efficient user identification scheme with key distribution preserving anonymity for distributed computer networks. Computers & Security, 23(2):120–125, March 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804000197.

Wiant:2005:ISP

[Wia05] Terry L. Wiant. Information security policy’s impact on reporting security incidents. Computers & Security, 24(6):448–459, September 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404805000490.

Willett:2000:FAC

[Wil00] Michael Willett. Features, attributes, characteristics, and traits (FACTs) of key recovery schemes/products. Computers & Security, 19(1):28–30, January 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800863592.

Wang:2005:CTK

[WL05] Shyh-Yih Wang and Chi Sung Laih. Cryptanalyses of two key assignment schemes based on polynomial interpolations. Computers & Security, 24(2):134–138, March 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804001865.

Wen:2006:PSA

[WLH06] Hsiang-An Wen, Chun-Li Lin, and Tzonelih Hwang. Provably secure authenticated key exchange protocols for low power computing clients. Computers & Security, 25(2):106–113, March 2006. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404805001732.

Wang:2003:CET

[WLT03] Bin Wang, Jian-Hua Li, and Zhi-Peng Tong. Cryptanalysis of an enhanced timestamp-based password authentication scheme. Computers & Security, 22(7):643–645, October 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404803007132.

Wolfe:2003:CF

[Wol03a] Dr. Henry B. Wolfe. Computer forensics. Computers & Security, 22(1):26–28, January 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803001056.

Wolfe:2003:EE

[Wol03b] Hank Wolfe. Encountering encryption. Computers & Security, 22(5):388–391, July 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803005042.

Wolfe:2003:EAa

[Wol03c] Hank Wolfe. Evidence acquisition. Computers & Security, 22(3):193–195, April 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803003043.

Wolfe:2003:EAb

[Wol03d] Hank Wolfe. Evidence analysis. Computers & Security, 22(4):289–291, May 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803004048.

Wolfe:2003:FET

[Wol03e] Hank Wolfe. Forensic evidence testimony — some thoughts. Computers & Security, 22(7):577–579, October 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803007041.

Wolfe:2003:SEE

[Wol03f] Hank Wolfe. Setting up an electronic evidence forensics laboratory. Computers & Security, 22(8):670–672, December 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740480300004X.

Wolfe:2003:CS

[Wol03g] Henry B. Wolfe. The circumstances of seizure. Computers & Security, 22(2):96–98, February 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803002049.

Wolthusen:2008:Ea

[Wol08a] S. D. Wolthusen. Editorial. Computers & Security, 27(1–2): 1, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740480800014X.

Wolthusen:2008:Eb

[Wol08b] S. D. Wolthusen. Editorial. Computers & Security, 27(1–2): 1, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740480800014X.

Wolthusen:2008:Ec

[Wol08c] S. D. Wolthusen. Editorial. Computers & Security, 27(1–2): 1, March 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740480800014X.

Wolthusen:2008:Ed

[Wol08d] S. D. Wolthusen. Editorial. Computers & Security, 27(3–4):63, May/June 2008. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000473.

Wolthusen:2008:Ee

[Wol08e] S. D. Wolthusen. Editorial. Computers & Security, 27(7–8):233–234, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000886.

Wolthusen:2008:Ef

[Wol08f] S. D. Wolthusen. Editorial. Computers & Security, 27(7–8):233–234, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000886.

Wolthusen:2008:Eg

[Wol08g] S. D. Wolthusen. Editorial. Computers & Security, 27(7–8):233–234, December 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404808000886.

Wolthusen:2008:ECa

[Wol08h] S. D. Wolthusen. From the Editor-in-Chief. Computers & Security, 27(5–6):123, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000709.

Wolthusen:2008:ECb

[Wol08i] S. D. Wolthusen. From the Editor-in-Chief. Computers & Security, 27(5–6):123, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000709.

Wolthusen:2008:ECc

[Wol08j] S. D. Wolthusen. From the Editor-in-Chief. Computers & Security, 27(5–6):123, October 2008. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808000709.

Wolthusen:2009:E

[Wol09] S. D. Wolthusen. Editorial. Computers & Security, 28(1–2):1, February/March 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480800120X.

Wood:2001:WDR

[Woo01] Charles Cresson Wood. What do the recent terrorist attacks mean for the American information security profession? Computers & Security, 20(8):667–670, December 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801008045.

Wool:2004:UUD

[Woo04] Avishai Wool. The use and usability of direction-based filtering in firewalls. Computers & Security, 23(6):459–468, September 2004. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404804000951.

Wright:2001:CRC

[Wri01] Adrian Wright. Controlling risks of e-commerce content. Computers & Security, 20(2):147–154, April 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740480100205X.

Ward:2002:DAC

[WS02] Peter Ward and Clifton L. Smith. The development of access control policies for information technology systems. Computers & Security, 21(4):356–371, August 1, 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404802004145.

Wu:2005:HC

[Wu05] Chuan-Kun Wu. Hash channels. Computers & Security, 24(8):653–661, November 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404805000696.

Williams:2000:GIK

[WZ00] Chuck Williams and Nevenko Zunic. Global interoperability for key recovery. Computers & Security, 19(1):48–55, January 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404800863634.

Williams:2000:KRA

[WZM^{+}00] Chuck Williams, Nevenko Zunic, Stephen M. Matyas, Sarbari Gupta, and Michael Willett. Key recovery alliance (KRA) technology papers, special issue — introduction. Computers & Security, 19(1):18–20, January 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404800863579.

Wang:2009:DPW

[WZM09] Fangwei Wang, Yunkai Zhang, and Jianfeng Ma. Defending passive worms in unstructured P2P networks based on healthy file dissemination. Computers & Security, 28(7):628–636, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404809000741.

Yu:2004:KDI

[YC04] Enzhe Yu and Sungzoon Cho. Keystroke dynamics identity verification-its problems and practical solutions. Computers & Security, 23(5):428–440, July 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804000963.

Yang:2009:IBR

[YC09] Jen-Ho Yang and Chin-Chen Chang. An ID-based remote mutual authentication with key agreement scheme for mobile devices on elliptic curve cryptosystem. Computers & Security, 28(3–4):138–143, May/June 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404808001120.

Yeh:2001:HDI

[YH01] Wen-Hung Yeh and Jing-Jang Hwang. Hiding digital information using a novel system scheme. Computers & Security, 20(6):533–538, September 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404801006149.

Yang:2007:MTIa

[YH07a] Jianhua Yang and Shou-Hsuan Stephen Huang. Mining TCP/IP packets to detect stepping-stone intrusion. Computers & Security, 26(7–8):479–484, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807000934.

Yang:2007:MTIb

[YH07b] Jianhua Yang and Shou-Hsuan Stephen Huang. Mining TCP/IP packets to detect stepping-stone intrusion. Computers & Security, 26(7–8):479–484, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807000934.

Yang:2007:MTIc

[YH07c] Jianhua Yang and Shou-Hsuan Stephen Huang. Mining TCP/IP packets to detect stepping-stone intrusion. Computers & Security, 26(7–8):479–484, December 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807000934.

Yang:2007:PAA

[YH07d] Jianhua Yang and Shou-Hsuan Stephen Huang. Probabilistic analysis of an algorithm to compute TCP packet round-trip time for intrusion detection. Computers & Security, 26(2):137–144, March 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404806001398.

Yang:2004:ACH

[YL04] Cungang Yang and Celia Li. Access control in a hierarchy using one-way hash functions. Computers & Security, 23(8):659–664, December 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404804001919.

Yoon:2005:IHL

[YRY05] Eun-Jun Yoon, Eun-Kyung Ryu, and Kee-Young Yoo. An improvement of Hwang–Lee–Tang’s simple remote user authentication scheme. Computers & Security, 24(1):50–56, February 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404804001749.

Yang:2004:CUF

[YW04] Chou-Chen Yang and Ren-Chiun Wang. Cryptanalysis of a user friendly remote authentication scheme with smart cards. Computers & Security, 23(5):425–427, July 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804001002.

Yang:2004:NEU

[YWB^{+}04] Yanjiang Yang, Shuhong Wang, Feng Bao, Jie Wang, and Robert H. Deng. New efficient user identification and key distribution scheme providing enhanced security. Computers & Security, 23(8):697–704, December 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804001920.

Yang:2005:SAS

[YWL05] Chou-Chen Yang, Ren-Chiun Wang, and Wei-Ting Liu. Secure authentication scheme for session initiation protocol. Computers & Security, 24(5):381–386, August 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804002640.

Yu:2007:AMA

[YZ07] Ming Yu and Xi-Yuan Zhou. An adaptive method for anomaly detection in symmetric network traffic. Computers & Security, 26(6):427–433, September 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404807000764.

Yao:2009:STC

[YZPL09] Lihong Yao, Xiaochao Zi, Li Pan, and Jianhua Li. A study of on/off timing channel based on packet delay distribution. Computers & Security, 28(8):785–794, November 2009. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809000510.

Zhang:2005:IWP

[ZC05] Ruishan Zhang and Kefei Chen. Improvements on the WTLS protocol to avoid denial of service attacks. Computers & Security, 24(1):76–82, February 2005. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404804002597.

Zhu:2004:DSM

[ZCW04] Lie Huang Zhu, Yuan Da Cao, and Dong Wang. Digital signature of multicast streams secure against adaptive chosen message attack. Computers & Security, 23(3):229–240, May 2004. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404804000707.

Zenkin:2001:FAI

[Zen01a] Denis Zenkin. Fighting against the invisible enemy: Methods for detecting an unknown virus. Computers & Security, 20(4):316–321, July 31, 2001. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404801004072.

Zenkin:2001:GPC

[Zen01b] Denis Zenkin. Guidelines for the protecting the corporate against viruses. Computers & Security, 20(8):671–675, December 1, 2001. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404801008057.

Zhang:2009:MIE

[ZHH09] Zonghua Zhang, Pin-Han Ho, and Liwen He. Measuring IDS-estimated attack impacts for rational incident response: a decision theoretic approach. Computers & Security, 28(7):605–614, October 2009. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000273.

Zhong:2002:PKM

[Zho02] Sheng Zhong. A practical key management scheme for access control in a user hierarchy. Computers & Security, 21(8):750–759, November 2002. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404802008155.

Zhong:2003:CCC

[ZL03] Sheng Zhong and Tianwen Lin. A comment on the Chen-Chung scheme for hierarchical access control. Computers & Security, 22(5):450–452, July 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404803005133.

Zhang:2008:BNAa

[ZLCF08a] Shaojun Zhang, Jianhua Li, Xiuzhen Chen, and Lei Fan. Building network attack graph for alert causal correlation. Computers & Security, 27(5–6):188–196, October 2008. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404808000308.

Zhang:2008:BNAb

[ZLCF08b] Shaojun Zhang, Jianhua Li, Xiuzhen Chen, and Lei Fan. Building network attack graph for alert causal correlation. Computers & Security, 27(5–6):188–196, October 2008. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404808000308.

Zhang:2008:BNAc

[ZLCF08c] Shaojun Zhang, Jianhua Li, Xiuzhen Chen, and Lei Fan. Building network attack graph for alert causal correlation. Computers & Security, 27(5–6):188–196, October 2008. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404808000308.

Zuccato:2004:HSR

[Zuc04] Albin Zuccato. Holistic security requirement engineering for electronic commerce. Computers & Security, 23(1):63–76, February 2004. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404804000653.

Zuccato:2007:HSM

[Zuc07]  Albin Zuccato.  Holistic security management frame-

work applied in electronic commerce. Computers & Security, 26(3):256–265, May 2007. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480600188X.

Zunic:2000:OCR

[Zun00] Nevenko Zunic. Organization considerations for retrieval of stored data via key recovery methods. Computers & Security, 19(1):82–85, January 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480086366X.